### v2

Abstract
We provide Dutch-book arguments against two forms of misspecified Bayesian learning. An agent progressively learns about a state and is offered a bet after every new discovery. We say the agent is Dutch-booked when they are willing to accept all bets, but their payoff is negative under each state either ex-post, or in expectation given the objective conditional probabilities of the discoveries (i.e., the correct data-generating process, DGP). Respectively, an agent cannot be Dutch-booked if and only if they update their beliefs with Bayes rule either from the previous belief, even using misspecified likelihood functions, or from one lexicographic prior, using the correct data-generating process. Under a large population interpretation of the DGP, this means that a population can suffer aggregate losses under all states when different individuals update their beliefs from different (lexicographic) priors, or using misspecified likelihoods. Thus, the Dutch-book argument offers a general characterization of the perils of misspecification.
Keywords: Dutch book, Bayesian updating, lexicographic beliefs, model misspecification
1 Introduction
The classical Dutch-book argument (Ramsey 1931, De Finetti 1937) is a cornerstone of probability theory. It says that an agent (he) may accept a set of bets that return an aggregate loss under every state (i.e., a Dutch book) if his preferences cannot be rationalized by one probabilistic belief. Later (Lehman 1955, Kemeny 1955), the converse was also proven: an agent cannot be Dutch-booked if he evaluates all bets with the same probabilistic belief about the state.
In this Dutch-book argument there is no time dimension. But the agent may progressively receive information about a fundamental state, and the bookmaker (she) could offer one bet after every piece of news. In this paper, we consider the following learning environment. There are a set of decision-relevant fundamental states and a set of contingencies, partially ordered in time, in which the agent makes choices. The fundamental states represent different (probabilistic) DGPs for contingencies. Thus, the contingencies have objective probabilities conditional on the fundamental states. Then, a question arises: Should the beliefs at counterfactual contingencies be somehow coherent to guarantee sound decision-making? We answer this question with a two-fold Dutch-book argument. [^1] Epistemologists have studied this problem in a framework where the agent learns finer and finer partitions of the state space (see Rescorla, 2022 for a recent survey). Under the assumption that every state is deemed possible at the outset, Teller (1973) and Lewis (1999) have shown that an agent cannot be Dutch-booked if and only if he updates beliefs by conditioning. Rescorla (2022) extends the argument to the presence of subjectively impossible states. While in the partitional framework each state can only give rise to one learning path, and thus counterfactual contingencies can only be consistent with different states, in the framework of this paper a state can result into different paths, and thus be consistent with different counterfactual contingencies.
We call deterministic Dutch book a system of bets on the fundamental state, one for each contingency, so that the cumulative payoff of the agent is non-positive under all fundamental states and paths, and negative under some fundamental state and path. More generally, we call Dutch book a system of bets that yields a non-positive expected payoff under every fundamental state, and a negative expected payoff under some fundamental state, where the expectations are computed with the probabilities of the contingencies as determined by the fundamental state under consideration. An agent is (deterministically) Dutch-booked when she is willing to accept all the bets in a (resp. deterministic) Dutch book. [^2] We use the following tie-breaking rule: the agent does not accept bets with zero subjective expected payoff.
In other words, a deterministic Dutch book guarantees that under every possible DGP the agents surely receives a negative payoff, and possibly gets a strictly negative payoff. Instead, a Dutch book requires that the agent’s expected payoff given the probabilities over contingencies induced by every DGP is nonpositive, and strictly negative for some DGP.
The notion of deterministic Dutch book extends to this paper’s framework the notion of “diachronic Dutch book” employed by the arguments for conditioning. With this, we show that an agent cannot be deterministically Dutch-booked if and only if his beliefs are forward-consistent, i.e., the agent updates beliefs with Bayes rule whenever possible. Forward-consistency does not require to use the correct conditional probabilities of the contingencies, so the agent cannot be deterministically Dutch-booked even if they are misspecified about the relative probabilities in the likelihood function, i.e., they may update on the basis of a wrong likelihood function that links the fundamental states and the contingencies. It only rules out the more radical form of misspecification where a state is subjectively believed possible given some contingency that is objectively impossible under that state. Moreover, forward-consistency is equivalent to Bayes-plausibility (Kamenica and Gentzkow, 2011), i.e., the requirement that each current belief is a convex combination of the beliefs at the next contingencies. Thus, Bayes-plausibility is the coherence rule among beliefs at counterfactual contingencies that characterizes an agent who cannot be deterministically Dutch-booked.
The necessity of forward-consistency to avoid Dutch-booking does not follow from the existing results that establish the necessity of conditioning in the partitional framework (see Rescorla, 2022 for the most general case), because of the multiplicity of learning paths that are compatible with the same state. Thus, given a violation of forward-consistency between a contingency and its successors, the bookmaker needs to find an acceptable bet for the first contingency such that the opposite bet will be accepted in each of the next contingencies (despite the potentially different posteriors), and not just in one as in the partitional framework. [^3] Another way of seeing this difficulty is that, in this paper, the bookmaker is constrained to offer bets on one dimension of the “grand state space” that also includes the learning paths. The solution to this problem will be an application of Farkas lemma.
Leveraging the separation between subjective (over fundamental states) and objective (given the fundamental state) uncertainty, the general notion of Dutch book aggregates payoffs across counterfactual contingencies according to the objective probability over them prescribed by the fundamental state. Such weaker form of Dutch book is important for two orders of reason.
First, a population of agents making choices in different contingencies will realize the aggregate payoff under the true state. So, if the population can be Dutch-booked in objective expected terms, their combined choices will select sure aggregate losses and forfeit sure aggregate wins, suggesting an evolutionary disadvantage. Alternatively, a single agent may perish in the long run after being exposed to i.i.d. occurrences of the same dynamic decision problem.
Second, the currently accepted interpretation of Dutch-book arguments is not the literal one, where a bookmaker can truly design a bet for each contingency, but rather that the agent suffers from some form of irrationality, which the Dutch-book argument brings to light.
We illustrate how general Dutch books work with an example.
Example 1
Larry wants to go out and watch the New Year’s fireworks, but the municipality does not disclose their location anymore to avoid crowding. There are three possible locations: main square, marina, and central park. There are three bus lines on his street: the red line goes to the square, the blue line to the marina, and the green line to the park. The buses always arrive in this order, at constant time intervals. Larry is late, so his only chance to see the fireworks is taking the first bus that arrives, regardless of its color, and hoping it goes to the right location. Suppose the blue line comes. When Larry arrives at the marina, he discovers whether the fireworks are there or not. If they are not, Larry regrets not moving to the bus stop a little earlier, which meant catching the red line: “I bet the fireworks are at the square!” Overhearing this, a bookmaker approaches Larry and proposes a bet on the fireworks’ location. Larry accepts a bet that yields $9$ if the fireworks are at the square and $-10$ if the fireworks are at the park. By the same token, at the park, he would accept a bet that yields $9$ for marina and $-10$ for square, and at the square, he would accept a bet that yields $9$ for park and $-10$ for marina. Thus, no matter the fireworks’ location, Larry ends up with equal probability in each of the other two locations, and hence he suffers an expected loss from this betting behavior.
In the example, the bus schedule determines the objective conditional probabilities of the three contingencies of interest. Instead, there are no objective probabilities of the fireworks’ location, and Larry does not have a subjective prior either; after all, he cannot decide which bus to take, so he does not need one. However, to deal with the bookmaker, he needs to form a belief after discovering that the fireworks are not in the location he arrived to. What’s wrong with the beliefs he forms? The location he deems more likely is different in each of the three contingencies. This form of incoherence among beliefs can arise precisely from the lack of a prior. Thus, the Dutch-book argument provides a motivation for having a prior, even when there is no decision to make before the arrival of information. More generally, the argument provides a justification for having coherent beliefs that can be rationalized with a prior, although the agent may not truly have one, but just reason in a coherent way after different discoveries. In the population interpretation, having a prior translates into the agents having the same prior. Thus, the Dutch book arguments suggests a disadvantage for “divided” populations, in which different agents have different (prior) beliefs, in that the agents with different priors can be exploited in different situation. Such divided population naturally emerges in a repeated setting if the agents do not share their experiences, thus suggesting an important role for the aggregation of feedback.
Another possible reason for Larry’s incoherent beliefs is that he (has a prior, but) perceives an inexistent correlation between bus arrivals and location of the fireworks. This is a classic form of model misspecification about the likelihood function connecting fundamental states and contingencies. In fact, we show that any form of model misspecification about the likelihood function allows Dutch-booking. This is perhaps surprising: since there is no “bad prior” over fundamental states (because there is no objective one), it is hard to conceive in what sense updating the prior with wrong likelihoods could generate a “bad posterior” over fundamental states. The Dutch-book argument offers a precise sense in which model misspecification is bad for decision-making.
In this simple example, having one prior and updating it by conditioning suffices to avoid Dutch-booking. But this is not true in general. We determine the coherence rule among beliefs that is necessary and sufficient to avoid Dutch-booking in objective expected terms. The coherence rule is easy to interpret; it is based on the idea that differences in the subjective odds ratios of states across contingencies shall only reflect differences in the objective odds ratios of the contingencies given the states. In the example, if Larry deems square twice more likely than park when he is at the marina and marina twice more likely than square when he is at the park, to satisfy the coherence rule he should deem marina four times more likely than park when he is at the square.
In turn, we show that this coherence rule characterizes a Completely Consistent Belief System: an array of beliefs (one for each contingency) that can be derived from a Lexicographic Conditional Probability System (Blume et al., 1991), a list of theories in a priority order. Specifically, each belief is derived with Bayes rule from the first theory that is consistent with the contingency. So, the LCPS plays the role of lexicographic prior, while the CCBS specifies the posterior beliefs in different contingencies. The fact that a completely consistent belief rather than simply conditioning from a unique prior is needed will be illustrated in Section 6 with an example of an agent who is subject to correlation neglect.
Related literature
In the context of sequential games, Siniscalchi (2022) introduces the notion of structural rationality, which is based on the use of CCBSs over the opponents’ strategies. [^4] Interestingly, structural rationality is motivated by the inability of sequential rationality to capture coherent conditional preferences that can be derived by updating a prior preference relation. CCBSs guarantee that all conditional preferences can be derived from a prior lexicographic preference relation. In other words, they guarantee a stronger form of Bayesianism which might be called “Bayesian updating of preferences” compared to Bayesian updating of beliefs. In Siniscalchi (2022), as well as Battigalli et al. (2023), a CCBS is defined as an array of beliefs, one for each information set, that is consistent with a Complete Conditional Probability System: an array of probability measures, one for each subset of opponents’ strategies, disciplined by the chain rule of probability. [^5] Conditional probability systems were first introduced by Renyi (1955). CCPSs were introduced in the analysis of games by Myerson (1986). In turn, a CCPS can be derived from a LCPS by conditioning. In a sequential game, a player progressively rules out opponents’ strategies by observing their moves, that is, learning occurs in the partitional fashion. So, Bayes rule boils down to conditioning, and hence the definition of CCBS in Siniscalchi (2022) and Battigalli et al. (2023) is a special case of the definition of this paper. Thus, the Dutch-book argument for complete consistency implies that the standard notion of conditional probability system defined on the collection of observable events falls short of the level of coherence that shelters from Dutch-booking in objective expected terms. The working paper of Siniscalchi (2022) also characterizes CCBSs with a generalization of the chain rule that applies between non-nested conditional events. In the Appendix, we provide a direct proof of the equivalence between the generalized chain rule and our coherence rule in sequential games.
Pettigrew (2023) considers an agent with a credence at time $t$ and a vector of possible credences at $t^{\prime}>t$, and shows that the agent can be Dutch-booked if and only if his credence at $t$ is not in the convex hull of the credences at $t^{\prime}$ (which he calls “General Reflection Principle”). This result yields our equivalence between Bayes-plausibility and the impossibility of deterministic Dutch-booking for the special case of a tree with depth $1$ and discoveries that do not rule out any state. [^6] When some discoveries objectively rule out some states, statement II.a of Theorem 1 in Pettigrew (2023) fails. From his theorem, Pettigrew also derives an argument for conditioning in the partitional framework. In our richer framework, we can derive our more general argument for forward consistency, uncovering the limited possibility of achieving a deterministic Dutch-booking against likelihood misspecification.
Chen et al. (2015) show that the posteriors of different agents, who update a common prior based on different discoveries, are common knowledge if and only if they cannot be Dutch-booked in aggregate. [^7] See also Lehrer and Samet (2011) for an alternate characterization of “agreeing to agree” between only two agents. Their notion of Dutch book only takes the realized discoveries and sums up the payouts under each state. This amounts to assuming that the bookmaker can choose the bets based on the entire set of realized discoveries. So, when the agents have different (thus not commonly known) posteriors, the bookmaker can exploit an informational advantage over each single agent, who is unaware of the other agents’ discoveries. In other words, the bookmaker can make profit from arbitrage. This is ruled out by our notion of Dutch book, in which each possible discovery is associated with a bet, independently of what other discoveries (using the population interpretation) may have been made by other agents. This reconciles the possibility of Dutch-booking Bayesian agents in Chen et al. (2015) with the impossibility results of this paper.
There are several related contribution in the literature on misspecified Bayesian learning, who is excellently surveyed in Bohren and Hauser (2024a). Lanzani (2025) characterizes a behavior rule that prevents misspecification to damage the agent beyond his maxmin payoff preserving consistency when correctly specified. Bohren and Hauser (2024b) characterizes when a departure from Bayesian updating (e.g., underinference from signals) can be rationalized as the consequence of Bayesian updating within a misspecified model. Molavi (2024) shows that beliefs are consistent with Bayesian updating (from a possibly misspecified model) if and only if the mean posterior is absolutely continuous with respect to the prior. This finding generalizes the earlier Shmaya and Yariv (2016) to the case in which the subjective beliefs of the decision maker may not have the same support of the true data-generating process. Cerreia-Vioglio et al. (2025) show that in coordination games misspecification can be extremely damaging even for misspecification concerned agents. Compared to this literature, we provide a novel reason for why some misspecification should or should not be expected to persist based on their susceptibility to Dutch books.
Finally, our results are similar in spirit to those in the no-trade literature, the closest being in Morris (1994). Morris considers agents who trade claims over commodities that are contingent on a payoff-relevant state and on the realization of a signal. In the case of two traders, a single money commodity and risk neutrality, Morris’ model coincides with ours given a tree of depth $1$. Assuming that the traders have no incentive to exchange claims that are contingent on the state only (which in our setting boils down to having the same prior), Theorem 2.1 in Morris (1994) shows that there exists trade that gives a positive expected payoff to each trader if and only if they disagree on the likelihood function mapping states into probabilities over signals. Our main theorem further shows (in a more general learning framework) that such trade can take a very specific form, whereby one trader has a positive expected payoff conditional on each state, and the other has a positive expected payoff conditional on each signal. If the first trader (the bookmaker) is using the correct likelihood function, then the second trader (our agent) gets an objectively negative expected payoff conditional on each state, and thus is Dutch-booked. Conversely, we show that the existence of such trade guarantees that the disagreement between the two traders must be about the likelihood function, not (just) the prior.
The paper is organized as follows. Section 1 introduces the framework. Section 2 illustrates belief systems. Section 3 introduces the novel notions of Dutch book. Section 4 states and proves the Dutch-book theorem for forward-consistency. Section 5 is devoted to the Dutch-book theorem for complete consistency. In the Appendix, we formalize the special case of sequential games and the connection with Siniscalchi (2022).
2 Framework
There is a finite set $\mathcal{S}$ of possible (fundamental) states. We describe the possible learning processes through an arborescence (i.e., a set of trees) of contingencies $\mathcal{H}$. With this, we are not assuming the existence of just one initial contingency in which the agent forms beliefs before learning anything about the state. A learning path is a path of contingencies from an “initial” contingency that is not preceded by any other contingency (i.e., the root of a tree), to a “terminal” contingency that is not followed by any other contingency (i.e., a leaf of that tree). Let $\mathcal{L}$ denote the set of possible learning paths and $L(h)$ the set of paths that go through a contingency $h$, and $H(l)$ the set of contingencies through which learning path $l$ goes. Each state $s\in\mathcal{S}$ induces an objective distribution $\eta^{s}\in\Delta(\mathcal{L})$ over learning paths, and thus states can be interpreted as different possible data-generating processes. Let $L(s)=\mathrm{supp}\eta^{s}$. For each $h\in\mathcal{H}$, let $p(h|s)=\sum_{l\in L(h)}\eta^{s}(l)$ denote the probability of reaching contingency $h$ under state $s$. Let $S(h)$ denote the set of states $s$ that are consistent with contingency $h$, that is, $p(h|s)>0$. Without loss of generality, we assume that every contingency is consistent with some state. An important special case is the one in which the probability of a contingency $h$ is the same across all the consistent states $s$, that is, $p(h|s)=p(h|s^{\prime})=:p^{h}$ for all $h\in\mathcal{H}$ and $s,s^{\prime}\in S(h)$. We will refer to this as the “independence case”.
To illustrate the framework, we formalize the example of the Introduction.
Example 2
There are three possible states, corresponding to the location of the fireworks: $\mathcal{S=}\left\{sq,ma,pa\right\}$. There are six possible contingencies, each with a different set of consistent states, so we label each contingency $h$ with the only state or with the initials of the two states in $S(h)$:
=$\mathcal{H=}\left\{sq,ma,pa,sm,mp,ps\right\}.$
For instance, if Larry takes the bus that goes to the square, he ends up in contingency $sq$ if the fireworks are there, and in contingency $mp$ otherwise. Note that all the contingencies are unordered, therefore contingencies and learning paths coincide. The arrival process of the buses is independent of the state. Thus, we are in the independence case: for each $h\in\left\{sm,mp,ps\right\}$ and $s,s^{\prime}\in S(h)$, $p(h|s)$ and $p(h|s^{\prime})$ are equal, because they coincide with the probability of the bus that leads to contingency $h$. In particular, since the three lines arrive with equal frequencies, each state $s\in\mathcal{S}$ induces a uniform distribution $\eta^{s}$ over the three learning paths/contingencies $h$ such that $s\in S(h)$. So, $p(h|s)=p(h|s^{\prime})=1/3$. $\triangle$
3 Belief systems
We assume that, in each contingency $h$, the agent has a probabilistic belief $\mu(\cdot|h)\in\Delta(\mathcal{S})$ such that $\mu(S(h)|h)=1$ — we call $(\mu(\cdot|h))_{h\in\mathcal{H}}$ belief system. This maintained assumption, besides requiring probabilistic sophistication at any point in time, rules out the most severe forms of model misspecification in which an agent assigns positive probability to states that have become objectively impossible.
We say that a belief system is forward-consistent if the agent’s beliefs are consistent with Bayesian updating from his previous belief using some (conditional) likelihood function, whenever possible. Observe that such conditional likelihood function is not required to be the objectively correct one.
Definition 1
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is forward-consistent if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist conditional probability distributions
$\left(\widetilde{p}(\cdot|h,s)\right)_{s\in S(h)}\in\left(\Delta(\left\{h^{1},...,h^{n}\right\})\right)^{\left|S(h)\right|}$
such that, for each $m\in\left\{1,...,n\right\}$, if $\widetilde{p}(h^{m}|h,s)>0$ for some $s\in\mathrm{supp}\mu(\cdot|h)\subseteq S(h)$, then $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$.
It is well-known that beliefs are consistent with this form of updating if and only if the “prior belief” (i.e., the belief at $h$) is a convex combination of the “posterior beliefs” (i.e., the beliefs at $h^{1},...,h^{n}$). (The Appendix contains a formal proof of this fact.)
Definition 2
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is Bayes-plausible if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist $\alpha^{1},...,\alpha^{n}\in\left[0,1\right]$ such that $\alpha^{1}+...+\alpha^{n}=1$ and
$\sum_{m=1}^{n}\alpha^{m}\mu(\cdot|h^{m})=\mu(\cdot|h).$
As we will show, Bayes-plausibility is the minimal requirement of coherence among beliefs that shelters from deterministic Dutch-booking.
We will define a Completely Consistent Belief System as a belief system that can be derived from a Lexicographic Conditional Probability System.
Definition 3 (Blume et al., 1991)
A list of probability measures $\bar{\mu}=(\mu^{1},...,\mu^{n})$ over $\mathcal{S}$ is a Lexicographic Conditional Probability System if for every $s\in\mathcal{S}$, there exists exactly one $m\in\left\{1,...,n\right\}$ such that $\mu^{m}(s)>0$.
Definition 3 embodies a “full-support” requirement: every state is given positive probability at some level of the LCPS. In this way, from every LCPS, one can always derive a belief system as follows.
Definition 4
A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is completely consistent if there exists a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$ such that, for each $h\in\mathcal{H}$, $\mu(\cdot|h)$ can be derived with Bayes rule using $\left(\eta^{s}\right)_{s\in\mathcal{S}}$ from the earliest $\mu^{m}$ such that $\mu^{m}(S(h))>0$, i.e., $m=\min\left\{k:\mu^{k}(S(h))>0\right\}$ and
$\mu(s|h)=\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in S}p(h|s^{\prime})\mu^{m}(s^{\prime})}\qquad\forall s\in S(h)\text{.}$
[…]
We will characterize CCBSs in terms of coherence among the subjective relative probabilities of states in different contingencies. However, given that the objective probabilities of contingencies may differ across states, these comparisons are meaningful only after discounting the probability attributed to a state in a contingency by the probability that the state induces the contingency.
Definition 5
Given a belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$, a contingency $h$, and a pair of states $s,s^{\prime}\in S(h)$ such that $\mu(s|h)+\mu(s^{\prime}|h)>0$, the (possibly infinite) discounted odds ratio between $s$ and $s^{\prime}$ at $h$ is
$o(s,s^{\prime}|h)=\frac{\mu(s|h)}{p(h|s)}\frac{p(h|s^{\prime})}{\mu(s^{\prime}|h)}.$
A generalized odds ratio between $s$ and $s^{\prime}$ is the product of a finite concatenation of discounted odds ratios
$o(s,s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{\prime}|h^{n})$
such that, if one discounted odds ratio is 00, no discounted odds ratio is infinite.
A discounted odds ratio compares the probabilities attributed to two states in the same contingency. A generalized odds ratio compares the probabilities of two states, $s$ and $s^{\prime}$, in two different contingencies, $h^{1}$ and $h^{n}$, via a chain of comparisons with other states ($s$ with $s^{1}$, $s^{1}$ with $s^{2}$, …, until $s^{n-1}$ with $s^{\prime}$) in different contingencies ($h^{1}$ for $s,s^{1}$, $h^{2}$ for $s^{1},s^{2}$, and so on). [^8] The formal definition allows contingencies and states to repeat, because this will be convenient for the proofs. In the independence case, discounted odds ratios boil down to (standard) odds ratios.
The strong rule of coherence that shelters from Dutch-booking in objective expected terms is the equivalence of all generalized odds ratios between any two states.
Definition 6
We say that a belief system is completely coherent when, for every $s,s^{\prime}\in\mathcal{S}$, the generalized odds ratios between $s$ and $s^{\prime}$ are all identical.
We illustrate the notions of belief system and generalized odds ratio through the example from the Introduction.
Example 3
For each state $s$, a belief system $\mu=\left(\mu(\cdot|h)\right)_{h\in\mathcal{H}}$ must assign probability $1$ to $s$ in the contingency labelled with $s$. Also, in this example, every belief system is trivially forward-consistent as every path consists of just one (terminal) contingency.
To obtain a CCBS, let us fix a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$. Suppose that $\mu^{1}$ assigns positive probability to at least $2$ states. Then, $\mu^{1}(S(h))>0$ for each $h\in\left\{sm,mp,ps\right\}$. Because each contingency has the same probability under the two consistent states, we are in the independent case: Bayes rule boils down to the chain rule, and discounted odds ratios boil down to odds ratios. So, for each $h\in\left\{sm,mp,ps\right\}$, writing $S(h)=\left\{s,s^{\prime}\right\}$, we have
$\mu(s|h)=\frac{\mu^{1}(s)}{\mu^{1}(s)+\mu^{1}(s^{\prime})}.$
If instead $\mu^{1}$ assigns probability $1$ to a state $\bar{s}$, we obtain $\mu(\bar{s}|h)=1$ whenever $\bar{s}\in S(h)$, and the belief in the contingency $h$ such that $S(h)=\mathcal{S}\backslash\left\{\bar{s}\right\}=\left\{s,s^{\prime}\right\}$ is given by
$\mu(s|h)=\frac{\mu^{2}(s)}{\mu^{2}(s)+\mu^{2}(s^{\prime})}.$
Given any two states $s,s^{\prime}$, one can compute an odds ratio between $s$ and $s^{\prime}$ in the only contingency $h$ such that $S(h)=\left\{s,s^{\prime}\right\}$; call it $o(s,s^{\prime})$. Then, one can also compute a generalized odds ratio between $s$ and $s^{\prime}$ as $o(s,s^{\prime\prime})\cdot o(s^{\prime\prime},s^{\prime})$ (s′′{$s^{\prime\prime}\notin\left\{s,s^{\prime}\right\}$), unless one is zero and the other infinite. For instance, when $\mu^{1}(ma)\in(0,1)$, we can compute
$o(sq,ma)\cdot o(ma,pa)$ $=$ $\frac{\mu(sq|sm)}{\mu(ma|sm)}\cdot\frac{\mu(ma|mp)}{\mu(pa|mp)}=\frac{\mu^{1}(sq)}{\mu^{1}(ma)}\cdot\frac{\mu^{1}(ma)}{\mu^{1}(pa)}$
$=$ $\frac{\mu^{1}(sq)}{\mu^{1}(pa)}=\frac{\mu(sq|ps)}{\mu(pa|ps)}=o(sq,pa),$
which can be infinite if $\mu^{1}(pa)=0$ and thus $\mu^{1}(sq)>0$, or zero, vice versa. Note the equality between the odds ratio and the generalized odds ratio.
An alternative belief system $\bar{\mu}$, which rationalizes Larry’s betting behavior, is the following:
$\bar{\mu}(sq|ps)=\bar{\mu}(ma|sm)=\bar{\mu}(pa|mp)=3/4\text{.}$
In this case, the odds ratio and the generalized odds ratio between any two states differ. For instance, in contingency $ps$ we have $o(sq,pa)=3$, but in contingency $sm$ we have $o(sq,ma)=1/3$ and in contingency $mp$ we have $o(ma,pa)=1/3$, therefore we can compute a generalized odds ratio between $sq$ and $pa$ of $1/9$ instead of $3$triangle
4 Dutch books and Dutch-booking
Traditionally, a Dutch book is a set of bets such that, no matter the realization of the state, the cumulative payoff from the bets is never positive and sometimes negative for the gambler (thus never negative and sometimes positive for the bookmaker). An agent is Dutch-booked if he accepts all the bets in a Dutch book. Our agent, however, would never accept a classical Dutch book if all the bets are proposed at the same time, as he has a well-defined probabilistic belief at any point in time. We now investigate whether, instead, a system of gambles proposed in different contingencies can induce the same kind of state-wise losses. [^9] Our agent risk neutrality and probabilistic sophistication at every contingency imply that, when he is willing to accept a bet, he is also willing to accept the sum of the bets, and thus we can assume without loss of generality that our agent is proposed just one bet in each contingency.
Definition 7
A gamble is a map $\gamma:\mathcal{S}\rightarrow\mathbb{R}$ that specifies, for each $s\in\mathcal{S}$, the gain or loss $\gamma(s)$ for the agent in case of $s$.
We say that $\gamma$ is a gamble on $\widetilde{S}\subseteq\mathcal{S}$ if $\gamma(s)=0$ for every $s\in\mathcal{S}\backslash\widetilde{S}$.
A system of gambles is an array $g=(g(\cdot|h))_{h\in\mathcal{H}}$ where, for each $h\in\mathcal{H}$, $g(\cdot|h)$ is a gamble on $S(h)$. [^10] The results would not change if the bets in a Dutch book were allowed to yield non-zero payoffs also under objectively impossible states.
What does it mean, for a system of gambles, to be a “Dutch book”? An intuitive notion is that the bookmaker obtains a never negative and sometimes positive profit under every state and path. We call this notion “deterministic Dutch book”. But the existence of objective conditional probabilities of paths allows to introduce also a more permissive notion of Dutch book, whereby the bookmaker obtains a non-negative/positive profit conditional on each funamental state only in objective expected terms.
Definition 8
A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a Dutch book if for every $s\in\mathcal{S}$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)\leq 0,$ (1)
and for some $s\in\mathcal{S}$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)<0.$ (2)
A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a deterministic Dutch book if for every $s\in\mathcal{S}$ and $l\in L\left(s\right)$
$\sum_{h\in H(l)}g(s|h)\leq 0,$ (3)
and for some $s\in\mathcal{S}$ and $l\in L\left(s\right)$,
$\sum_{h\in H(l)}g(s|h)<0.$ (4)
Note that a deterministic Dutch book does not depend on the objective probabilities of paths, hence it is a Dutch book no matter such probabilities. Moreover, Dutch books, deterministic or not, do not depend on the objective probabilities of the states.
Since each gamble is proposed to the agent in a specific contingency, the acceptance of the gamble will depend on the contingent belief.
Definition 9
Fix a probability measure $\nu\in\Delta(\mathcal{S})$ and a gamble $\gamma$. We say that the agent is willing to accept $\gamma$ given $\nu$ if
$\sum_{s\in\mathcal{S}}\nu(s)\gamma(s)>0\text{,}$
or $\gamma(s)=0$ for every $s\in\mathcal{S}$.
We say that an agent with belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is willing to accept a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ if, for every $h\in\mathcal{H}$, he is willing to accept $g(\cdot|h)$ given $\mu(\cdot|h)$.
We say that the agent is (deterministically) Dutch-booked if he is willing to accept a (deterministic) Dutch book.
Definition 9 captures the following tie-breaking rule: our agent declines gambles that yield zero subjective expected payoff (except for the trivial “null” gamble, as the bookmaker is free to offer no gamble). This rules out in a simple way the uninteresting case in which the agent can be Dutch-booked by just charging negative payoffs for the states he subjectively deems impossible.
To conclude this section, we illustrate Dutch-booking in the running example.
Example 4
Suppose that Larry has the belief system $\bar{\mu}$ constructed before. The bookmaker constructs the following system of bets:
$g(sq|ps)$ $=$ $g(ma|sm)=g(pa|mp)=9;$
$g(pa|ps)$ $=$ $g(sq|sm)=g(ma|mp)=-10;$
$g(s|h)$ $=$ 0 $0\text{ \ \ \qquad\qquad\qquad\qquad\qquad\qquad\qquad}\forall h\in\left\{sq,ma,pa\right\}.$
Under every state $s$, we have
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)=\frac{1}{3}(9-10+0)=-\frac{1}{3}<0,$
i.e., Larry has an objective expected payoff of $-1/3$. Therefore, the system is a Dutch book. There remains to show that Larry is willing to accept it. We have
$\sum_{s\in\mathcal{S}}\bar{\mu}(s|h)g(s|h)=\frac{3}{4}\cdot 9+\frac{1}{4}(-10)+0\cdot 0>0\quad\forall h\in\left\{sm,mp,ps\right\}.$
Therefore, Larry is Dutch-booked. Note that deterministic Dutch-booking is impossible in this example, as every path consists of just one contingency, in which the agent has a well-defined probabilistic belief. $\triangle$
5 The Dutch-book theorem for forward-consistency
[…]
Theorem 1
Consider an agent with belief system $\mu=(\mu(\cdot|h))_{h\in\mathcal{H}}$. The following are equivalent:
  1. 1.
    $\mu$ is Bayes-plausible;
  2. 2.
    $\mu$ is forward-consistent;
  3. 3.
    the agent cannot be deterministically Dutch-booked.
Note that forward consistency does not depend on the models the agent uses for updating. In this sense, misspecification about the relative probabilities of the likelihood function is immune to deterministic Dutch-booking. However, that is not the case for the stronger form of misspecification where at some contingency $h$ some objectively impossible state $s\notin S\left(h\right)$ is deemed possible by the decision maker. This is because, by assigning an arbitrarily high positive payout $g(s|h)>>0$ to the agent in $s$ given $h$, the bookmaker can obtain a profit under some state $s^{\prime}\in S\left(h\right)$ that is objectively possible.
Proof of Theorem 1.
$\mathbf{1\Rightarrow 2)}$
This is a known fact, however we provide formal proof in the Appendix.
$\mathbf{2\Rightarrow 3)}$
[…]
$\mathbf{3\Rightarrow 1)}$
We prove the counterpositive. Suppose that $\mu$ is not Bayes-plausible. Thus, there exists a non-terminal contingency $h$ such that $\mu(\cdot|h)$ is not a convex combination of the beliefs at the next contingencies $h^{1},...,h^{n}$. Let $A\in\mathbb{R}^{\left|\mathcal{S}\right|\times n}$ be the matrix of subjective probabilities $\mu(s|h^{j})$, and let $b\in\mathbb{R}^{\left|\mathcal{S}\right|}$ be the vector of subjective probabilities $\mu(s|h)$. So, for every transposed vector $x=(x^{1},...,x^{n})\in\left[0,1\right]^{n}$ such that $\sum_{m=1}^{n}x^{m}=1$, $Ax\not=b$. Moreover, for every $x\in\left[0,\infty\right)^{n}$ such that $\sum_{m=1}^{n}x^{m}\not=1$, $Ax\not=b$, because $Ax=b$ implies
$x^{1}+...+x^{n}$ $=$
$\sum_{j=1}^{n}\left(\sum_{s}\mu(s|h^{j})\right)x^{j}$ $=$
$\sum_{s}\sum_{j=1}^{n}x^{j}\mu(s|h^{j})$ $=$
$\sum_{s}\mu(s|h)$ $=$ $1$
Thus, there is no $x\geq 0$ such that $Ax=b$. But then, by Farkas lemma, there exists $y\in\mathbb{R}^{\left|\mathcal{S}\right|}$ such that
$A^{T}y\geq 0\text{ and }b^{T}y<0.$
For some $\varepsilon>0$, by $-b^{T}y>0$ the gamble with vector of prizes $-y-2\overrightarrow{\varepsilon}$ is acceptable at $h$, and by $A^{T}y\geq 0$ the gamble with vector of prizes $y+\overrightarrow{\varepsilon}$ is acceptable at each $h^{j}$. Complete the system with gambles with zero prizes for all states at every other contingency. We obtain a deterministic Dutch book: if one of the paths that go through $h$ realizes, and those includes at least one $l\in L(s)$ for some $s\in S(h)$, the profit of the bookmaker is $\varepsilon>0$ (satisfying 4), otherwise it is 00 (satisfying 3). $\blacksquare$
6 The Dutch-book theorem for complete consistency
The Dutch-book theorem for complete consistency generalizes the insights from the example: incongruities among generalized odds ratios can be exploited for Dutch-booking and can only be displayed by a belief system that is not completely consistent. The theorem also shows the opposite implications: when there is no incongruity between generalized odds ratios, the agent’s belief system is completely consistent and he cannot be Dutch-booked. Thus, coherence among generalized odds ratios fully characterizes complete consistency and impossibility of Dutch-booking.
Theorem 2
Fix a belief system $\mu$. The following are equivalent.
  1. 1.
    $\mu$ is completely coherent;
  2. 2.
    $\mu$ is completely consistent;
  3. 3.
    the agent cannot be Dutch-booked.
Proof. The acronyms DOR and GOR will stand for discounted/generalized odds ratio.
$\mathbf{1\Rightarrow 2)}$
Note preliminarly that, by Condition 1, if a state $s$ has a DOR with another state $s^{\prime}$, all the DORs between $s$ and $s^{\prime}$ take the same value; [^12] Recall that an “indeterminate DOR” betwen $s$ and $s^{\prime}$ is just not considered a DOR, and likewise for GOR. We say that $s$ has a DOR with $s^{\prime}$ when there exists $h\in\mathcal{H}$ such that $s,s^{\prime}\in S(h)$ and $\mu(s|h)+\mu(s^{\prime}|h)>0$, the DOR between $s$ and $s^{\prime}$ being $\mu(s|h)/\mu(s^{\prime}|h)$. Similarly, we say that $s$ has a GOR with $s^{\prime}$ when it is possible to compute at least one GOR between $s$ and $s^{\prime}$. call it $o(s,s^{\prime})$.
Let $P^{1}$ be the set of all $s\in\mathcal{S}$ that have no zero GOR with any $s^{\prime}$.
Let $P^{2}$ be the set of all $s\in\mathcal{S}\backslash P^{1}$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash P^{1}$.
Let $P^{3}$ be the set of all $s\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$.
And so on.
Claim 1: there exists $n\geq 1$ such that $(P^{1},...,P^{n})$ is a partition of $\mathcal{S}$, with non-empty elements.
Proof: Let $n$ be the smallest non-negative integer such that $P^{n+1}=\emptyset$; it exists because $\mathcal{S}$ is finite and the sets are disjoint by construction. There remains to show that $\cap_{m=1}^{n}P^{m}=\mathcal{S}$ (which also implies that $n\geq 1$). Let $\widetilde{S}=\mathcal{S}\backslash\left(\cap_{m=1}^{n}P^{m}\right)$ and suppose by contraposition that $\widetilde{S}\not=\emptyset$. The fact that $P^{n+1}=\emptyset$ implies that every $s\in\widetilde{S}$ has a zero GOR with some $s^{\prime}\in\widetilde{S}$. Consider a directed graph with nodes $\widetilde{S}$ and arrows that represent a zero GOR between the start and the end node. Since every node is the start of an arrow, the graph has a cycle. So, the product of all the (zero) GORs along the cycle is a zero GOR between $s^{1}$ and itself. Inverting all the DORs in this GOR, we obtain an infinite GOR between $s^{1}$ and itself. This violates the assumption that all the GORs between $s^{1}$ and itself are identical by Condition 1. $\square$
Claim 2: there exists a LCPS $\bar{\mu}=(\nu^{1},...,\nu^{n})$ such that, for each $m=1,...,n$, (a) $\mathrm{supp}\nu^{m}=P^{m}$ and (b), for every pair $\left(s,s^{\prime}\right)\in P^{m}\times P^{m}$ which has a DOR $o(s^{\prime},s)$, $\nu^{m}(s^{\prime})/\nu^{m}(s)=o(s^{\prime},s)$.
Proof: For each $m=1,...,n$, we will construct a probability measure $\nu^{m}$ that satisfies (a) and (b); then, $(\nu^{1},...,\nu^{n})$ is an LCPS because, by Claim 1, $(P^{1},...,P^{n})$ partitions $\mathcal{S}$.
Fix $m=1,...,n$ and construct a graph with nodes $P^{m}$ and connections between nodes that have a DOR. For every connected component $\widetilde{P}\subseteq P^{m}$, pick a state $s$ arbitrarily, let $\widetilde{\nu}(s)=1$, and progressively let $\widetilde{\nu}(s^{\prime\prime})=\widetilde{\nu}(s^{\prime})\cdot o(s^{\prime\prime},s^{\prime})$ for every $s^{\prime\prime}\in\widetilde{P}$, where $o(s^{\prime\prime},s^{\prime})$ is a DOR between $s^{\prime\prime}$ and some $s^{\prime}\in\widetilde{P}$ for which $\widetilde{\nu}(s^{\prime})$ was already defined; by definition of $P^{m}$, $o(s^{\prime\prime},s^{\prime})$ is not zero and not infinite (otherwise $o(s^{\prime},s^{\prime\prime})$ would be zero). So, we can obtain a probability measure $\nu^{m}$ with support $P^{m}$ by rescaling $\widetilde{\nu}$.
To show that $\nu^{m}$ satisfies the desiderata, fix $s,s^{\prime}\in P^{m}$ that have a DOR $o(s^{\prime},s)$, thus $s$ and $s^{\prime}$ are in the same connected component. Then, the graph has a path $(\bar{s}^{1},...,\bar{s}^{T})$ with $\bar{s}^{1}=s$ and $\bar{s}^{T}=s^{\prime\prime}$ such that, for each $t=2,...,T$, $\widetilde{\nu}(\bar{s}^{t})=\widetilde{\nu}(\bar{s}^{t-1})\cdot o(\bar{s}^{t},\bar{s}^{t-1})$, where $o(\bar{s}^{t},\bar{s}^{t-1})\in\mathbb{R}^{++}$ is a DOR between $\bar{s}^{t}$ and $\bar{s}^{t-1}$. (Note that, if $\widetilde{\nu}(\bar{s}^{t-1})$ was defined as $\widetilde{\nu}(\bar{s}^{t})\cdot o(\bar{s}^{t-1},\bar{s}^{t})$ and not vice versa, there is also a DOR $o(\bar{s}^{t},\bar{s}^{t-1})=1/o(\bar{s}^{t-1},\bar{s}^{t})$ between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.) Therefore
$\frac{\nu^{m}(s^{\prime})}{\nu^{m}(s)}=\frac{\widetilde{\nu}(s^{\prime})}{\widetilde{\nu}(s)}\equiv\frac{\widetilde{\nu}(\bar{s}^{T})}{\widetilde{\nu}(\bar{s}^{1})}=o(\bar{s}^{T},\bar{s}^{T-1})\cdot\frac{\widetilde{\nu}(\bar{s}^{T-1})}{\widetilde{\nu}(\bar{s}^{1})}=...=\prod\limits_{t=2}^{T}o(\bar{s}^{t},\bar{s}^{t-1}).$
This is a GOR between $s^{\prime}$ and $s$, and it is equal to $o(s^{\prime},s)$ by Condition 1. $\square$
Now fix $h\in\mathcal{H}$. Let $m$ be the smallest $k$ such that $\nu^{k}(S(h))>0$. We show that $\mu(\cdot|h)$ can be derived from $\nu^{m}$ with Bayes rule. Let $\hat{\nu}^{m}$ be the probability measure derived from $\nu^{m}$ with Bayes rule given $h$ (and $\left(\eta^{s}\right)_{s\in\mathcal{S}}$). By property (a) of $\bar{\mu}$, $\mathrm{supp}\nu^{m}=P^{m}$, thus $\mathrm{supp}\hat{\nu}^{m}=P^{m}\cap S(h)=:\hat{S}$. Moreover, for all $s,s^{\prime}\in\hat{S}$,
$\frac{\hat{\nu}^{m}(s)}{\hat{\nu}^{m}(s^{\prime})}=\frac{p(h|s)\nu^{m}(s)}{p(h|s^{\prime})\nu^{m}(s^{\prime})}.$
To show that $\mu(\cdot|h)$ and $\hat{\nu}^{m}$ coincide, it is enough to show that $\mathrm{supp}\mu(\cdot|h)=\hat{S}$ and $\mu(s|h)/\mu(s^{\prime}|h)=\hat{\nu}^{m}(s)/\hat{\nu}^{m}(s^{\prime})$ for all $s,s^{\prime}\in\hat{S}$.
First we show that $\mathrm{supp}\mu(\cdot|h)=P^{m}\cap S(h)$.
Fix $s\in\mathrm{supp}\mu(\cdot|h)$. Thus, $s\in S(h)$. Hence, for every $k<m$, $s\not\in P^{k}$, as $\mathrm{supp}\nu^{k}=P^{k}$ but $\nu^{k}(S(h))=0$. Now we show that, for every $k>m$, $s\not\in P^{k}$ either, so that $s\in P^{m}$. Suppose by contradiction that $s\in P^{k}$. As $\mathrm{supp}\nu^{m}=P^{m}$ and $\nu^{m}(S(h))>0$, there exists $\bar{s}\in S(h)\cap P^{m}$ such that $\nu^{m}(\bar{s})>0$. Since $s\in\mathrm{supp}\mu(\cdot|h)$, $s$ has a non-zero DOR with every $s^{\prime}\in S(h)$, so in particular with $\bar{s}$. Since $s\not\in P^{1}\cup...\cup P^{m}$, $s$ has a zero GOR with some $s^{\prime}\in P^{m}\cup...\cup P^{n}$. The two things combined impy that $\bar{s}$ has a zero GOR with $s^{\prime}$ too. But this contradicts $\bar{s}\in P^{m}$.
Now fix $s\not\in\mathrm{supp}\mu(\cdot|h)$. If $s\not\in S(h)$, we are done. If $s\in S(h)$, $s\not\in P^{m}$, because $s$ has a zero DOR with every $s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and we have just shown that $\mathrm{supp}\mu(\cdot|h)\subseteq P^{m}$.
To conclude the proof, fix a pair $s,s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and write
$\frac{\mu(s|h)}{\mu(s^{\prime}|h)}=\frac{p(h|s)}{p(h|s^{\prime})}\frac{\mu(s|h)}{p(h|s)}\frac{p(h|s^{\prime})}{\mu(s^{\prime}|h)}=\frac{p(h|s)}{p(h|s^{\prime})}o(s,s^{\prime})=\frac{p(h|s)\nu^{m}(s)}{p(h|s^{\prime})\nu^{m}(s^{\prime})},$
where $o(s,s^{\prime})$ is the DOR between $s$ and $s^{\prime}$ at $h$, and the last equality follows from property (b) of $\nu^{m}$.
$\mathbf{2\Rightarrow 3)}$
Let $\bar{\mu}=(\mu^{1},...,\mu^{n})$ be a LCPS from which the belief system is derived. For each $m=1,...,n$, let $S^{m}$ denote the support of $\mu^{m}$, and let $H^{m}$denote the set of all $h\in\mathcal{H}$ such that $\mu(\cdot|h)$ is derived from $\mu^{m}$.
Fix a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$. Suppose that $i$ is willing to accept $g$. We show that $g$ is not a Dutch book. Suppose that $g$satisfies condition (1) of a Dutch book, so that,
$\forall s\in\mathcal{S},\text{ \ \ }\sum_{h\in\mathcal{H}}p(h|s)g(s|h)\leq 0.$ (5)
We are going to show that $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in\mathcal{H}$, so that $g$ does not satisfy condition (2) of a Dutch book.
The proof is by induction on $m$. Fix $m=1,...,n$ and suppose by way of induction that for every $k<m$, $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in H^{k}$. (This is vacuously true for $m=1$.) Fix $s\in S^{m}$. For every $k>m$ and $h\in H^{k}$, $p(h|s)=0$, otherwise $\mu(\cdot|h)$ could be derived from $\mu^{m}$. By this and the induction hypothesis, inequality (5) yields
$\sum_{h\in H^{m}}p(h|s)g(s|h)\leq 0.$
So, considering that $\mu^{m}(s^{\prime})=0$ for every $s^{\prime}\not\in S^{m}$, we can write
$\sum_{s\in S}\mu^{m}(s)\sum_{h\in H^{m}}p(h|s)g(s|h)\leq 0\text{.}$ (6)
Since $\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})>0$ for every $h\in H^{m}$, we can rewrite (6) as
$\sum_{h\in H^{m}}\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\leq 0.$ (7)
For each $h\in H^{m}$, since the agent is willing to accept $g(\cdot|h)$, we have
$\sum_{s\in\mathcal{S}}\mu(s|h)g(s|h)\geq 0.$ (8)
Since $\mu(\cdot|h)$ can be derived from $\mu^{m}$ with Bayes rule, we can rewrite (8) as
$\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\geq 0.$ (9)
Equations (9) and (7) imply
$\forall h\in H^{m},\text{ \ \ }\sum_{s\in S}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)=0,$
that is,
$\sum_{s\in S}\mu(s|h)g(s|h)=0.$
This means that the agent has no strict incentive to accept $g(\cdot|h)$, therefore acceptance implies $g(s|h)=0$ for all $s\in\mathcal{S}$.
$\mathbf{3\Rightarrow 1)}$
The proof that an agent with two different GORs between a state $s$ and a state $s^{\prime}$ can be Dutch-booked resembles the classical Dutch-book arguments of De Finetti and Ramsey. Therefore, the proof is deferred to the Appendix. $\blacksquare$
The next example illustrates the theorem.
Example 5
This is a variation of the famous Monty Hall problem, in which the initially selected door is determined at random and not revealed to the decision maker. There are three doors. A car is placed behind one of the doors, and a goat is placed behind each of the other two doors. The state represents the door with the car: $\mathcal{S}=\left\{A,B,C\right\}$. Then, one door, call it $d$, is drawn from a uniform distribution (in the Monty Hall problem, this is the door assigned to the agent). Finally, another door behind which there is a goat, call it $h\not=d$, is opened; it is the first non-car door preceding $d$ in an $a$-$b$-$c$ circle. So, for instance, conditional on $s=B$, door $a$ is opened with probability $2/3$ (the probability that $d\in\left\{b,c\right\}$), and door $c$ is opened with probability $1/3$ (the probability that $d=a$). The agent only observes which door is opened, so the possible contingencies are: =$\mathcal{H}=\left\{a,b,c\right\}$. Overall, conditional on each $s\in\mathcal{S}$, the probabilities of each $h\in\mathcal{H}$ are:
$p(h|s)$ $h=a$ $h=b$ $h=c$
$s=A$ 00 $\frac{1}{3}$ $\frac{2}{3}$
$s=B$ $\frac{2}{3}$ 00 $\frac{1}{3}$
$s=C$ $\frac{1}{3}$ $\frac{2}{3}$ 00
We next consider two possible belief systems for the agent.
First, suppose that, at every $h\in\mathcal{H}$, the agent assigns equal probability to each of the two states in $S(h)$. Then, the agent can be Dutch-booked (in objective expected terms). To see this, compute the three discounted odds ratios:
$o(B,C|a)$ $=$ $\frac{\Pr(a|C)}{\Pr(a|B)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(C,A|b)$ $=$ $\frac{\Pr(b|A)}{\Pr(b|C)}=\frac{1/3}{2/3}=\frac{1}{2};$
$o(B,A|c)$ $=$ $\frac{\Pr(c|A)}{\Pr(c|B)}=\frac{2/3}{1/3}=2.$
Since
$o(B,C|a)\cdot o(C,A|b)\not=o(B,A|c),$
we have two different generalized odds ratios between $B$ and $A$, and thus by Theorem 2 the agent can be Dutch-booked. For example, suppose that, at each $h\in\mathcal{H}$, a bookmaker offers to the agent the following bets:
$g(C|a)$ $=$ 3, $3,\text{ \ }g(B|a)=-2;$
$g(A|b)$ $=$ 3, $3,\text{ \ }g(C|b)=-2;$
$g(B|c)$ $=$ 3, $2\displaystyle 3,\text{ \ }g(A|c)=-$ 2 .
Clearly, the agent accepts all bets, and the objective expected payoff of the bookmaker conditional on each $s\in\mathcal{S}$ is:
$-g(A|b)p(b|A)-g(A|c)p(c|A)$ $=$ $-g(B|c)p(c|B)-g(B|a)p(a|B)=-g(C|a)p(a|C)-g(C|b)p(b|C)$
[…]
So, the agent is harmed by the “correlation neglect” he experiences.
Second, suppose that the agent derives his beliefs as follows. He has a uniform prior, and, thinking of the correct updating in the original Monty Hall problem, believes that the car is more likely to be behind the unopened door that has not been assigned to him (i.e., not door $d$). For instance, suppose that $h=a$ and the agent correctly computes
$\Pr\left(d=b|h=a\right)$
$=$ $\frac{\Pr(h=a|d=b)\Pr(d=b)}{\Pr(h=a|d=b)\Pr(d=b)+\Pr(h=a|d=c)\Pr(d=c)}$
$=$ $\frac{\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)}{\left(\begin{array}[]{c}\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)\\ +\Pr(h=a|d=c,s=B)+\Pr(h=a|d=c,s=C)\end{array}\right)}$
$=$ $\frac{2}{2+1}=\frac{2}{3}.$
So, he has $\mu(s=B|h=a)<1/2$. However, given $h=a$, the car is more likely to be behind door $b$:
$\Pr\left(s=B|h=a\right)$
$=$ $\frac{\Pr(h=a|s=B)\Pr(s=B)}{\Pr(h=a|s=B)\Pr(s=B)+\Pr(h=a|s=C)\Pr(s=C)}$
$=$ $\frac{\Pr(s=B)}{\Pr(s=B)+\Pr(d=b)\Pr(s=C)}$
$=$ $\frac{1}{1+\frac{1}{3}}=\frac{3}{4}.$
In other words, the agent perceives the opposite correlation between the fundamental state and the observed contingency. This increases the discrepancy between generalized odds ratios, making it easier for the bookmaker to Dutch-book the agent. [^13] To be precise, without limited liability any discrepancy allows the bookmaker to extract an infinite amount of money (in expectation) from the agent. But under limited liability, the higher discrepancy allows to strictly increase the bookmaker’s expected profit.
7 Conclusion
Dutch-book arguments have already been used to motivate probabilistic sophistication and conditioning. We introduce novel Dutch-book arguments that motivate two forms of belief consistency in a learning framework a la Anscombe-Aumann, with separation between subjective uncertainty (the state) and objective uncertainty (lotteries over learning paths conditional on the state). Bayes-plausibility is necessary and sufficient to avoid being Dutch-booked in a deterministic sense, i.e., under each realized state and learning path. A stronger form of coherence, guaranteed by deriving all beliefs from a subjective lexicographic prior with the correct models, is necessary and sufficient to avoid being Dutch-booked “in objective expected terms”, i.e., under each realized state, given the corresponding lottery on the learning paths. The latter form of Dutch-booking entails sure aggregate losses across a population of agents, thus showing the perils of heterogenous priors across different individuals, and of model misspecification even in absence of an objective prior.
Appendix A Appendix
A.1 Sequential games: connection with Battigalli et al. (2023) and Siniscalchi (2020, 2022)
In the context of a sequential game with complete information and perfect recall, Battigalli et al. (2023) introduce a classification of the belief systems of a player over the opponents’ strategies. We study their classification of belief systems as a special case of the analysis of the main body of the paper. In a game, the moves of a player are independent of the moves of the co-players that are yet to be observed, and the two jointly determine the path of the game, thus the learning path. Therefore, the assumption that every contingency is reached with the same probability under every consistent state is satisfied in that context. For concreteness, we stick to the sequential games application, but the analysis extends seamlessly to situations where this assumption is satisfied, for instance in situations where the an agent does not choose what to observe but the factors that determine the observations are independent of the state itself, as in the running example of the paper.
Let $S_{-i}$ be the set of stategy profiles of the opponents of player $i$. Let $H_{i}$ be the arborescence of information sets of player $i$, endowed with the precedence relation $\prec$. Each information set $h_{i}$ corresponds to the observation of some past moves of the opponents, summarized by the event $S_{-i}(h_{i})\subseteq S_{-i}$. A belief system is an array $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ of probability measures over $S_{-i}$ such that $\mu_{i}(S_{-i}(h_{i})|h_{i})=1$ for every $h_{i}\in H_{i}$.
Taking as primitive a collection of conditioning events $\mathcal{C\subseteq}2^{S_{-i}}\backslash\left\{\emptyset\right\}$, a conditional probability system over $S_{-i}$ is an array of probability measures $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in\mathcal{C}}$ such that $\bar{\mu}_{i}(C|C)=1$ for every $C\in\mathcal{C}$, and for every $D\in\mathcal{C}$ with $D\subset C$, for each $E\subseteq D$,
$\bar{\mu}_{i}(E|C)=\bar{\mu}_{i}(E|D)\bar{\mu}_{i}(D|C).$
A conditional probability system is complete when it is defined with respect to the collection $\mathcal{C=}2^{S_{-i}}\backslash\left\{\emptyset\right\}$.
Definition 10 (Battigalli et al., 2023)
A belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ over the opponents’ strategies is forward consistent if, for all $h_{i},h_{i}^{\prime}\in H_{i}$, if $h_{i}\prec h_{i}^{\prime}$, then
$\mu_{i}(E|h_{i})=\mu_{i}(S_{-i}(h_{i}^{\prime})|h_{i})\mu(E|h_{i}^{\prime})$
for every $E\subseteq S_{-i}(h_{i}^{\prime})$.
Analogously to the general notion of Definition 1, forward consistency corresponds to Bayesian updating of conjectures along the path of play (with the inconsequential difference that consistency is not only required between a contingency and its immediate successors). However, in this particular setting Bayesian updating boils down to the chain rule.
Definition 11
[…]
$\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i})).$
The definition of complete consistency of Battigalli et al. (2023) uses CCPS’s to discipline beliefs at counterfactual contingencies. An information set $h$ does not carry any additional information about the strategies of the opponents besides ruling out all the strategies that do not belong to $S_{-i}(h)$; for this reason, it is appropriate to derive each belief by mere conditioning in place of Bayes rule, from the most general theory that explains the contingency.
Proposition 1
For every information set $h_{i}$, for every pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h_{i})$, and for every probability distribution over player $i$’s strategies, the probability of reaching the information set is identical under $s_{-i}$ and $s_{-i}^{\prime}$: $p(h_{i}|s_{-i})=p(h_{i}|s_{-i}^{\prime})$.
Proof. Under each $s_{-i}$, the probability of reaching $h_{i}$ is the probability of the strategies of $i$ that, with $s_{-i}$, induce a path that goes through $h_{i}$. So all we need to show is that such set of strategies of $i$ is the same under all $s_{-i}\in S_{-i}(h_{i})$. By perfect recall, all these strategies must prescribe the same moves before reaching $h_{i}$. Suppose by contradiction that there exists a strategy of $i$ that induces a path that goes through $h_{i}$ with some $s_{-i}\in S_{-i}(h_{i})$ but not with some other $s_{-i}^{\prime}\in S_{-i}(h_{i})$. But then, all other strategies of $i$ that prescribe the same moves before $h_{i}$, will have the same effect: $h_{i}$ is not reached under $s_{-i}^{\prime}$. This contradicts that $s_{-i}^{\prime}$ is consistent with $h_{i}$. $\blacksquare$
Although the use of CCPSs is appropriate in this context, one might as well model the theories from which a player derives just as a LCPS, and derive each belief by conditioning the first theory in the list that is consistent with the contingency; the two definitions of complete consistency are equivalent. [^14] The equivalence between LCPSs and CCPSs is well-known. We formalize and prove this result for the analysis to be self-contained.
Proposition 2
Fix a belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$. The following are equivalent:
  1. 1.
[…]
    $\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}));$
  2. 2.
    there exists a LCPS $\hat{\mu}_{i}=(\mu_{i}^{1},...,\mu_{i}^{n})$ such that for each $h_{i}\in H_{i}$ and $s_{-i}\in S_{-i}(h_{i})$,
    $\mu_{i}(s_{-i}|h_{i})=\frac{\mu_{i}^{k}(s_{-i})}{\mu_{i}^{k}(S_{-i}(h_{i}))},$
    where $k$ is the smallest $m$ such that $\mu_{i}^{k}(S_{-i}(h_{i}))>0$.
Proof.
[…]
1$\Rightarrow$2) Create a decreasing chain of conditioning events $(C^{1},...,C^{n})$ as follows. Let $C^{1}=S_{-i}$. For each $k>1$, let $C^{k}$ be the union of all the $s_{-i}\in S_{-i}$ such that $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$. Let $n$ be the smallest $k$ such that $C^{k+1}=\emptyset$. Consider the list of measures $\hat{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C^{1}),...,\bar{\mu}_{i}(\cdot|C^{n}))$. For each $h_{i}\in H_{i}$, let $k$ be the smallest $j\leq n$ such that $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})>0$ — it exists by $C^{n+1}=\emptyset$. By $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})=0$ for every $j<k$, it follows that $S_{-i}(h_{i})\subseteq C^{j}$. Hence, $\mu_{i}(\cdot|h_{i})$ can be derived from $\bar{\mu}_{i}(\cdot|C^{k})$ by updating, and $\bar{\mu}_{i}(\cdot|C^{k})$ is also the first measure in $\hat{\mu}_{i}$ that gives positive probability to $S_{-i}(h_{i})$. There remains to show that $\hat{\mu}_{i}$ is a LCPS. For each $s_{-i}\in S_{-i}$, by $C^{n+1}=\emptyset$, $\bar{\mu}_{i}(s_{-i}|C^{k})>0$ for some $k\leq n$, so $\hat{\mu}_{i}$ has full joint support. Moreover, by $s_{-i}\in C^{k}$, $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$, and by $\bar{\mu}_{i}(s_{-i}|C^{k})>0$, $s_{-i}\not\in C^{j}$ for every $j>k$, which implies $\bar{\mu}_{i}(s_{-i}|C^{j})=0$. Hence, $\hat{\mu}_{i}$ has disjoint supports and is therefore a LCPS. $\blacksquare$
Proposition 2 implies that the main result of this paper also applies to belief consistency as it was defined for sequential games. With this, the characterization of complete consistency with the coherence rule among discounted odds ratios can be rewritten using traditional odds ratios in place of discounted odds ratios, because in this context the two coincide by Proposition 1.
Definition 12
Given a belief system $(\mu_{i}(\cdot|h))_{h\in H_{i}}$, an information set $h$, and a pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h)$, let
$o^{\ast}(s_{-i},s_{-i}^{\prime}|h)=\frac{\mu_{i}(s_{-i}|h)}{\mu_{i}(s_{-i}^{\prime}|h)}$
denote the odds ratio between $s_{-i}$ and $s_{-i}^{\prime}$ at $h$, provided that it is not indeterminate. A simple generalized odds ratio between $s_{-i}$ and $s_{-i}^{\prime}$ is the product of a finite concatenation of odds ratios
$o^{\ast}(s_{-i},s_{-i}^{1}|h^{1})\cdot o^{\ast}(s_{-i}^{1},s_{-i}^{2}|h^{2})\cdot....\cdot o^{\ast}(s_{-i}^{n-1},s_{-i}^{\prime}|h^{n})$
that is not indeterminate.
The following is a corollary of the Dutch-book theorem for complete consistency and of the last two propositions.
Corollary 1
Fix a belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$. The following are equivalent:
the belief system is completely consistent;
for every $s_{-i},s_{-i}^{\prime}\in S_{-i}$, all the simple generalized odds ratios between $s_{-i}$ and $s_{-i}^{\prime}$ are identical.
Siniscalchi (2020) (the working paper of Siniscalchi 2022) comes to a similar characterization of complete consistency starting from a different idea. Consider a belief system that is isomorphic to a conditional probability system defined over the collection of observable events. Such belief system thus satisfies the chain rule between contingencies that correspond to nested observations, but the chain rule may have little bite when the observable events are not nested, and the corresponding beliefs are not disciplined by the preliminary observation of a larger event. Therefore, Siniscalchi (2020) generalizes the chain rule to apply also between overlapping, but non-nested observable events. We conclude this section by showing the convergence between Siniscalchi’s approach and mine.
Call “generalized self-odds ratio” a generalized odds ratio between $s_{-i}$ and itself.
Remark 1
All the generalized odds ratios of each pair are identical if and only if all generalized self-odds ratios are $1$.
Proof. Only if. Fix a generalized self-odds ratios. Break it into two and invert one of the two halves. We have two concatenations of odds ratios that start and end with the same element. Thus, the two corresponding generalized odds ratios are identical by assumption. Therefore, the generalized self-odds ratio is $1$.
If. Fix two concatenations of odds ratios where either no odds ratio is infinite or there is no zero. If both contain a zero, the corresponding generalized odds ratios are both zero. If they both contain an infinite, they are both infinite. If one contains an infinite and the other contains a zero, the generalized self-odds ratio obtained by inverting one of the two is zero or infinite, which violates the assumption. Otherwise, the generalized self-odds ratio (in one direction) is finite. By assumption, it is $1$. Therefore the two generalized odds ratios are identical. $\blacksquare$
Definition 13 (Siniscalchi 2020)
A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if for every $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$,
$\mu^{i}(E|h_{i}^{1})\prod_{m=1}^{n-1}\mu^{i}(S_{-i}(h_{i}^{m})\cap S_{-i}(h_{i}^{m+1})|h_{i}^{m+1})=\mu^{i}(E|h_{i}^{n})\prod_{m=1}^{n-1}\mu^{i}(S_{-i}(h_{i}^{m})\cap S_{-i}(h_{i}^{m+1})|h_{i}^{m}).$ (11)
Theorem 3
A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if and only if it is completely consistent.
Proof. We will use the characterization of complete consistency of Corollary 1.
If part. Fix $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$. If both sides of equation 11 are zero, consistency holds. Otherwise, suppose without loss of generality that the right-hand side is not zero, so that equation 11 can be rewritten as
$\frac{\mu^{i}(E|h_{i}^{1})}{\mu^{i}(S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{2})|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(S_{-i}(h_{i}^{n-1})\cap S_{-i}(h_{i}^{n})|h_{i}^{n})}{\mu^{i}(E|h_{i}^{n})}=1.$ (12)
Note that equation 12 holds if for every $s_{-i}\in E$ such that the ratio $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ is not indeterminate,
$\frac{\mu^{i}(s_{-i}|h_{i}^{1})}{\mu^{i}(S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{2})|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(S_{-i}(h_{i}^{n-1})\cap S_{-i}(h_{i}^{n})|h_{i}^{n})}{\mu^{i}(s_{-i}|h_{i}^{n})}=1,$ (13)
because this means that all the non-indeterminate ratios $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ are identical, which implies that the ratio $\mu^{i}(E|h_{i}^{1})/\mu^{i}(E|h_{i}^{n})$ takes the same value too. We are going to show that for every $m=2,...,n$, there is $s_{-i}^{m}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that
$\frac{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m-1})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})};$
then, we can rewrite equation 13 as
$\frac{\mu^{i}(s_{-i}|h_{i}^{1})}{\mu^{i}(s_{-i}^{2}|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(s_{-i}^{n}|h_{i}^{n})}{\mu^{i}(s_{-i}|h_{i}^{n})}=1,$ (14)
which is true by complete consistency (cf. Remark 1). Let $s_{-i}^{m}$ be any $s_{-i}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that $\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})>0$; one exists by the fact that the right-hand side of 11 is not zero. By complete consistency, for each $s_{-i}^{\prime}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$, we have
$\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{\prime}|h_{i}^{m})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}{\mu^{i}(s_{-i}^{\prime}|h_{i}^{m-1})}.$
With some algebra, it is easy to show that then
$\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m-1})|h_{i}^{m-1})},$
as we wanted to show.
Only if part. By Remark 1 we can show that for each $s_{-i},\in S_{-i}$, for every generalized self-odds ratio, equation 14 holds. Without loss of generality, we can assume that the generalized self-odds ratio is not infinite (because an infinite one can be inverted and becomes zero), so that no denominator in equation 14 is zero. For each $m=2,...,n$, by consistency, equation 11 for $E=\left\{s_{-i}^{m}\right\}$ and $\left(h_{i}^{1},...,h_{i}^{n}\right)=(h_{i}^{m-1},h_{i}^{m})$ yields
$\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}=\frac{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m-1})|h_{i}^{m-1})}.$
Then, substituing in 14, we obtain equation 13, which coincides with equation 12 for $E=\left\{s_{-i}\right\}$, and thus holds by consistency.
A.2 Completion of the proofs of the theorems
Proof of Theorem 1: $1\Rightarrow 2$
Fix a non-terminal contingency $h$ and let $h^{1},...,h^{n}$ be the next contingencies. Let $\alpha^{1},...,\alpha^{n}$ denote the weights that verify the definition of Bayes-plausibility. We construct probability distributions $(p(h^{m}|h,s))_{m=1}^{n}$ of reaching each $h^{m}$ given $h$ and each $s\in S(h)$. Fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, let $p(h^{m}|h,s)=0$ for all $s\in\mathcal{S}$. Suppose now that $\alpha^{m}>0$. For each $s\in\mathcal{S}$, let
$p(h^{m}|h,s)$ $=$ $\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\text{ \ \ if }s\in\mathrm{supp}\mu(\cdot|h),$
$p(h^{m}|h,s)$ $=$ $\alpha^{m}\text{ \ \ \ \ \ \ \ \ \ if }s\not\in\mathrm{supp}\mu(\cdot|h).$
For each $s\not\in\mathrm{supp}\mu(\cdot|h)$, we have
$\sum_{m=1}^{n}p(h^{m}|h,s)=\sum_{m=1}^{n}\alpha^{m}=1.$
For each $s\in\mathrm{supp}\mu(\cdot|h)$, we have
$\sum_{m=1}^{n}p(h^{m}|h,s)=\frac{\sum_{m=1}^{n}\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}=1,$
where the last equality follows from Bayes-plausibility.
Now fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, then
$\sum_{s\in S(h)}p(h^{m}|h,s)=0.$
If $\alpha^{m}>0$, by Bayes-plausibility
$\mathrm{supp}\mu(\cdot|h^{m})\subseteq\mathrm{supp}\mu(\cdot|h).$ (15)
So we have
$\frac{p(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}p(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=\frac{\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\mu(s|h)}{\alpha^{m}\sum_{s^{\prime}\in\mathrm{supp}\mu(\cdot|h)}\mu(s|h^{m})}=\mu(s|h^{m}),\qquad\forall s\in\mathcal{S}$
where the last equality follows from (15). Thus, $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(p(h^{m}|h,s))_{s\in S(h)}$ whenever possible.
Proof of Theorem 2: $3\Rightarrow 1$
We prove the counterpositive. Suppose that there exist two different GORs between some $s$ and $s^{\prime}$. Then, inverting all the DORs in the second GOR and multiplying by the first GOR, [^15] Since the two GORs are different, they cannot be both infinite or both zero, therefore after inverting one of the two their product is not indeterminate. we obtain a new GOR between $s$ and itself, which is not $1$; write it as
$r:=o(s^{0},s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{n}|h^{n}),$
where $n\geq 2$, $s^{0}=s^{n}=s$, and for each $k=1,...,n$,
$o(s^{k-1},s^{k}|h^{k})=\frac{p(h^{k}|s^{k-1})}{p(h^{k}|s^{k})}\frac{\mu(s^{k-1}|h^{k})}{\mu(s^{k}|h^{k})}.$ (16)
Wwithout loss of generality, we can suppose that:
$s^{0},...,s^{n-1}$ are all distinct, because if $s^{j}=s^{k}$, either $r^{\prime}=o(s^{j},s^{j+1}|h^{j})\cdot...\cdot o(s^{k-1},s|h^{k})$, or $r^{\prime\prime}=r/r^{\prime}$ is a new GOR between $s$ and itself that is not $1$.
$r<1$, because if a GOR is larger than $1$, inverting all the DORs we obtain a GOR smaller than $1$ (still between $s$ and itself).
With this, we construct a Dutch Book that the agent is willing to accept. For every $h\in\mathcal{H}\backslash\left\{h^{1},...,h^{n}\right\}$, let $g(\widetilde{s}|h)=0$ for every $\widetilde{s}\in\mathcal{S}$, so that $i$ is willing to accept $g(\cdot|h)$. Fix $\varepsilon>0$. Let $g(\cdot|h^{1})$ be the following gamble on $S(h^{1})$:
$g(s^{0}|h^{1})$ $=$ $-1,$
$g(s^{1}|h^{1})$ $=$ $-g(s^{0}|h^{1})\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon=\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon,$ (17)
and $g(s^{\prime}|h^{1})=0$ for every $s^{\prime}\not=s^{0},s^{1}$. Recursively, for each $m=2,...,n$, let $g(\cdot|h^{m})$ be the following gamble on $S(h^{m})$:
$g(s^{m-1}|h^{m})$ $=$ $-g(s^{m-1}|h^{m-1})\frac{p(h^{m-1}|s^{m-1})}{p(h^{m}|s^{m-1})}-\varepsilon,$ (18)
$g(s^{m}|h^{m})$ $=$ $-g(s^{m-1}|h^{m})\frac{\mu(s^{m-1}|h^{m})}{\mu(s^{m}|h^{m})}+\varepsilon,$ (19)
and $g(s^{\prime}|h^{m})=0$ for every $s^{\prime}\not=s^{m},s^{m-1}$. For each $m=1,...,n$, the agent is willing to accept $g(\cdot|h^{m})$ because
$\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})+\mu(s^{m}|h^{m})\cdot g(s^{m}|h^{m})$
$=$ $\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})-g(s^{m-1}|h^{m})\cdot\mu(s^{m-1}|h^{m})+\varepsilon\cdot\mu(s^{m}|h^{m})$
$=$ $\varepsilon\cdot\mu(s^{m}|h^{m})>0\text{,}$
where $\mu(s^{m}|h^{m})>0$ because $r\not=\infty$. Thus, $i$ is willing to accept $g=(g(\cdot|h))_{h\in\mathcal{H}}$. If $m\not=n$, since $s^{m}\not=s^{m^{\prime}}$ for every $m^{\prime}\not=m$, we also have
$\sum_{h\in\mathcal{H}}p(h|s^{m})g(s^{m}|h)$ $=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})+p(h^{m+1}|s^{m})g(s^{m}|h^{m+1})$
$=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})-p(h^{m+1}|s^{m})g(s^{m}|h^{m})\frac{p(h^{m}|s^{m})}{p(h^{m+1}|s^{m})}-p(h^{m+1}|s^{m})\varepsilon$
$=$ $-p(h^{m+1}|s^{m})\varepsilon<0,$
where $p(h^{m+1}|s^{m})>0$ because $s^{m}\in S(h^{m+1})$ (DORs at a contingency are only defined between consistent states). Thus, $s^{m}$ verifies the strict inequality condition of a Dutch book. Every $\widetilde{s}\in\mathcal{S}\backslash\left\{s^{1},...,s^{n}\right\}$ trivially satisfies the weak inequality condition. Finally, we need to show that, for some $\varepsilon>0$, $s$ satisfies the weak inequality condition as well, so that $g$ is a Dutch book. Note that, for $\varepsilon=0$,
$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)$
$=$ $p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{n}|s)g(s^{n-1}|h^{n})\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
$=$ $-p(h^{1}|s)+p(h^{n}|s)g(s^{n-1}|h^{n-1})\frac{p(h^{n-1}|s^{n-1})}{p(h^{n}|s^{n-1})}\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-1}|h^{n-1})o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-1})\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-2})\frac{p(h^{n-2}|s^{n-2})}{p(h^{n-1}|s^{n-2})}\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)+p(h^{n-2}|s^{n-2})g(s^{n-2}|h^{n-2})o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h^{n})$
$=$ $-p(h^{1}|s)-p(h^{1}|s^{1})g(s^{0}|h^{1})\frac{\mu(s|h^{1})}{\mu(s^{1}|h^{1})}\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h)$
$=$ $-p(h^{1}|s)+p(h^{1}|s)r<0\text{,}$
where the second equality uses (19), the third (18), the fourth (16), and so on. By continuity, the strict inequality is preserved for sufficiently small $\varepsilon>0$. $\blacksquare$

### v3

@@
 Abstract
-We provide Dutch-book arguments against two forms of misspecified Bayesian learning.
-An agent progressively learns about a state and is offered a bet after every new discovery.
-We say the agent is Dutch-booked when they are willing to accept all bets, but their payoff is negative under each state either ex-post, or in expectation given the objective conditional probabilities of the discoveries (i.e., the correct data-generating process, DGP).
-Respectively, an agent cannot be Dutch-booked if and only if they update their beliefs with Bayes rule either from the previous belief, even using misspecified likelihood functions, or from one lexicographic prior, using the correct data-generating process.
-Under a large population interpretation of the DGP, this means that a population can suffer aggregate losses under all states when different individuals update their beliefs from different (lexicographic) priors, or using misspecified likelihoods.
-Thus, the Dutch-book argument offers a general characterization of the perils of misspecification.
+We provide Dutch-book arguments against misspecified Bayesian learning.
+An agent progressively learns about a state and is offered a bet after every discovery.
+We say the agent is deterministically Dutch-booked when they would accept all bets, but their payoff is ex-post negative under each state.
+More generally, we say that the agent is Dutch-booked when they would accept all bets, but their expected payoff under each fundamental state is negative.
+With this, the agent cannot be deterministically Dutch-booked if and only if they update their beliefs using Bayes’ rule, even with misspecified likelihoods.
+The agent cannot be Dutch booked if and only if they update their beliefs using Bayes’ rule with a lexicographic prior and using the correct data-generating process.
+We show that offers of financial instruments and behavior in Monty Hall problems can be viewed as Dutch books that extract a sure expected gain from a misspecified population.
 Keywords: Dutch book, Bayesian updating, lexicographic beliefs, model misspecification
 1 Introduction
+A growing literature (surveyed in Bohren and Hauser, 2025) documents the consequences of updating beliefs using a misspecified model of the world.
+Model misspecification leads entirely rational (i.e., Bayesian) agents astray by inducing them to misinterpret the data and compute incorrect posteriors.
+Is there a general sense in which model misspecification is harmful for decision-making?
+We provide an answer to this question through a two-fold Dutch-book argument.
 The classical Dutch-book argument (Ramsey 1931, De Finetti 1937) is a cornerstone of probability theory.
-It says that an agent (he) may accept a set of bets that return an aggregate loss under every state (i.e., a Dutch book) if his preferences cannot be rationalized by one probabilistic belief.
+It says that an agent (he) may accept a set of bets that returns an aggregate loss under every state (i.e., a Dutch book) if his preferences cannot be rationalized by one probabilistic belief.
 Later (Lehman 1955, Kemeny 1955), the converse was also proven: an agent cannot be Dutch-booked if he evaluates all bets with the same probabilistic belief about the state.
-In this Dutch-book argument there is no time dimension.
-But the agent may progressively receive information about a fundamental state, and the bookmaker (she) could offer one bet after every piece of news.
-In this paper, we consider the following learning environment.
-There are a set of decision-relevant fundamental states and a set of contingencies, partially ordered in time, in which the agent makes choices.
+In the classical Dutch-book argument, there is no time dimension.
+However, the agent may progressively receive information about a fundamental state, and the bookmaker (she) could offer one bet conditional on every piece of news.
+Epistemologists have studied this problem in a framework where the agent learns finer and finer partitions of the state space on which he bets.
+[^1] An exception is Pettigrew (2023): see the literature review.
+Under the assumption that every state is deemed possible at the outset, Teller (1973) and Lewis (1999) have shown that an agent cannot be Dutch-booked if and only if he updates beliefs by conditioning.
+[^2] Rescorla (2022) surveys Dutch-book arguments and extends the argument for conditioning to the presence of subjectively impossible states.
+However, in many problems of dynamic choice, the decision-relevant state does not include the observed events; hence, learning is best represented as the acquisition of noisy signals.
+That is the approach that is followed in the entire economics literature on misspecified learning.
+In particular, this literature employs the following learning environment, which we also adopt here.
+There is a set of decision-relevant fundamental states and a set of contingencies, partially ordered in time, in which the agent makes choices.
 The fundamental states represent different (probabilistic) DGPs for contingencies.
 Thus, the contingencies have objective probabilities conditional on the fundamental states.
-Then, a question arises: Should the beliefs at counterfactual contingencies be somehow coherent to guarantee sound decision-making?
-We answer this question with a two-fold Dutch-book argument.
-[^1] Epistemologists have studied this problem in a framework where the agent learns finer and finer partitions of the state space (see Rescorla, 2022 for a recent survey).
-Under the assumption that every state is deemed possible at the outset, Teller (1973) and Lewis (1999) have shown that an agent cannot be Dutch-booked if and only if he updates beliefs by conditioning.
-Rescorla (2022) extends the argument to the presence of subjectively impossible states.
-While in the partitional framework each state can only give rise to one learning path, and thus counterfactual contingencies can only be consistent with different states, in the framework of this paper a state can result into different paths, and thus be consistent with different counterfactual contingencies.
-We call deterministic Dutch book a system of bets on the fundamental state, one for each contingency, so that the cumulative payoff of the agent is non-positive under all fundamental states and paths, and negative under some fundamental state and path.
-More generally, we call Dutch book a system of bets that yields a non-positive expected payoff under every fundamental state, and a negative expected payoff under some fundamental state, where the expectations are computed with the probabilities of the contingencies as determined by the fundamental state under consideration.
-An agent is (deterministically) Dutch-booked when she is willing to accept all the bets in a (resp.
-deterministic) Dutch book.
+Then, the question arises: what if the agent is (Bayesian but) misspecified about these probabilities?
+We call a deterministic Dutch book a system of bets on the fundamental state, one for each contingency, so that the cumulative payoff of the agent is nonpositive under all fundamental states and paths, and negative under some fundamental state and path.
+[^3] Thus, the notion of deterministic Dutch book extends to this paper’s framework the notion of “diachronic Dutch book” employed by the arguments for conditioning.
+More generally, we call a system of bets a Dutch book if it yields a nonpositive expected payoff under every fundamental state and a negative expected payoff under some fundamental state, where the expectations are computed with the probabilities of the contingencies as determined by the fundamental state under consideration.
+An agent is (deterministically) Dutch-booked when she is willing to accept all the bets in a (deterministic) Dutch book.
 [^4] We use the following tie-breaking rule: the agent does not accept bets with zero subjective expected payoff.
-In other words, a deterministic Dutch book guarantees that under every possible DGP the agents surely receives a negative payoff, and possibly gets a strictly negative payoff.
-Instead, a Dutch book requires that the agent’s expected payoff given the probabilities over contingencies induced by every DGP is nonpositive, and strictly negative for some DGP.
-The notion of deterministic Dutch book extends to this paper’s framework the notion of “diachronic Dutch book” employed by the arguments for conditioning.
-With this, we show that an agent cannot be deterministically Dutch-booked if and only if his beliefs are forward-consistent, i.e., the agent updates beliefs with Bayes rule whenever possible.
-Forward-consistency does not require to use the correct conditional probabilities of the contingencies, so the agent cannot be deterministically Dutch-booked even if they are misspecified about the relative probabilities in the likelihood function, i.e., they may update on the basis of a wrong likelihood function that links the fundamental states and the contingencies.
-It only rules out the more radical form of misspecification where a state is subjectively believed possible given some contingency that is objectively impossible under that state.
-Moreover, forward-consistency is equivalent to Bayes-plausibility (Kamenica and Gentzkow, 2011), i.e., the requirement that each current belief is a convex combination of the beliefs at the next contingencies.
-Thus, Bayes-plausibility is the coherence rule among beliefs at counterfactual contingencies that characterizes an agent who cannot be deterministically Dutch-booked.
-The necessity of forward-consistency to avoid Dutch-booking does not follow from the existing results that establish the necessity of conditioning in the partitional framework (see Rescorla, 2022 for the most general case), because of the multiplicity of learning paths that are compatible with the same state.
-Thus, given a violation of forward-consistency between a contingency and its successors, the bookmaker needs to find an acceptable bet for the first contingency such that the opposite bet will be accepted in each of the next contingencies (despite the potentially different posteriors), and not just in one as in the partitional framework.
-[^3] Another way of seeing this difficulty is that, in this paper, the bookmaker is constrained to offer bets on one dimension of the “grand state space” that also includes the learning paths.
+In other words, a deterministic Dutch book guarantees that, under every possible DGP, the agent surely receives a nonpositive payoff and possibly a strictly negative payoff.
+Instead, a Dutch book requires that the agent’s expected payoff, given the probabilities over contingencies induced by every DGP, is nonpositive and strictly negative for some DGP.
+First, we show that an agent cannot be deterministically Dutch-booked if and only if his beliefs are forward-consistent, i.e., the agent updates his beliefs using Bayes rule whenever possible.
+Since forward-consistency does not require the use of the correct conditional probabilities of the contingencies, agents that are Bayesian but misspecified about the relative probabilities in the likelihood function (i.e., they may apply Bayes rule to the wrong likelihood function linking states and the contingencies) are immune to deterministic Dutch books.
+However, an agent with the more radical form of misspecification, in which a state is subjectively believed possible given a contingency objectively impossible under that state, can be deterministically Dutch-booked.
+Moreover, forward consistency is equivalent to Bayes-plausibility (Kamenica and Gentzkow, 2011), i.e., the requirement that each belief is a convex combination of the beliefs that immediately follow.
+Thus, Bayes-plausibility is the coherence rule among beliefs in counterfactual contingencies that characterizes an agent who cannot be deterministically Dutch-booked.
+[^5] Note that the necessity of forward-consistency to avoid Dutch-booking does not follow from the existing results that establish the necessity of conditioning in the partitional framework, because of the multiplicity of learning paths that are consistent with the same state.
+Thus, given a violation of forward-consistency between a contingency and its successors, the bookmaker needs to find an acceptable bet for the first contingency such that the opposite bet will be accepted in each of the following contingencies (despite the potentially different posteriors), and not just in one as in the partitional framework.
+Another way of seeing this difficulty is that, in this paper, the bookmaker is constrained to offer bets on only one dimension of the “grand state space,” which also includes the learning paths.
 The solution to this problem will be an application of Farkas lemma.
-Leveraging the separation between subjective (over fundamental states) and objective (given the fundamental state) uncertainty, the general notion of Dutch book aggregates payoffs across counterfactual contingencies according to the objective probability over them prescribed by the fundamental state.
-Such weaker form of Dutch book is important for two orders of reason.
-First, a population of agents making choices in different contingencies will realize the aggregate payoff under the true state.
-So, if the population can be Dutch-booked in objective expected terms, their combined choices will select sure aggregate losses and forfeit sure aggregate wins, suggesting an evolutionary disadvantage.
-Alternatively, a single agent may perish in the long run after being exposed to i.i.d.
+Second, and most importantly, we show that a Bayesian agent cannot be Dutch-booked (“in objective expected terms”) if and only if he updates beliefs from a lexicographic prior with the correct likelihood function.
+This is perhaps surprising: since there is no “bad prior” over fundamental states (because there is no objective one), it is hard to conceive in what sense updating the prior with wrong likelihoods could generate a “bad posterior” over fundamental states.
+Our Dutch-book argument offers a precise sense in which model misspecification is harmful for decision-making: a misspecified agent could accept sure expected losses and forfeit sure expected wins.
+The argument also offers a motivation for having a prior, even when there is no decision to take before the arrival of information.
+The use of a lexicographic prior (in place of free belief revision after each surprise) is crucial to guarantee the coherence of beliefs across counterfactual contingencies, thereby shielding the agent from a Dutch book.
+We call a belief system completely consistent (henceforth CCBS) when it is derived from a subjective lexicographic prior with the correct likelihood function.
+More generally, the argument justifies having coherent beliefs that can be rationalized with a prior, although the agent may not truly have one; rather, the agent may reason coherently after different discoveries.
+We call complete coherence the rule of coherence between beliefs in counterfactual contingencies that shelters the agent from Dutch-booking.
+The rule is easy to interpret; it is based on the idea that differences in the subjective odds ratios of states across contingencies shall only reflect differences in the objective odds ratios of the contingencies given the states.
+Importantly, we also show that the possibility of Dutch-booking a misspecified agent does not rely on the bookmaker knowing the true data generating process perfectly.
+As long as a well defined measure of discrepancy between the agent and bookmaker beliefs is sufficiently large compared to the bookmaker misspecification, a Dutch book can be constructed.
+Leveraging the separation between subjective (over fundamental states) and objective (given the fundamental state) uncertainty, our general notion of Dutch book aggregates payoffs across counterfactual contingencies according to the objective probability over them prescribed by the fundamental state.
+Besides being a useful theoretical tool to uncover a form of “irrationality”, this weaker form of Dutch book could also be of applied relevance, even though a single agent will only find himself in one of many counterfactual contingencies.
+First, a large population of agents making choices in different contingencies will realize the aggregate payoff under the actual state.
+So, our Dutch-book argument suggests a disadvantage for “divided” populations, in which different agents hold different (prior) beliefs and can thus be exploited in different situations.
+Such a divided population naturally emerges in a repeated setting when agents do not share their experiences, suggesting an important role for aggregating feedback.
+Second, a single agent may perish in the long run after being exposed to i.i.d.
 occurrences of the same dynamic decision problem.
-Second, the currently accepted interpretation of Dutch-book arguments is not the literal one, where a bookmaker can truly design a bet for each contingency, but rather that the agent suffers from some form of irrationality, which the Dutch-book argument brings to light.
-We illustrate how general Dutch books work with an example.
-Example 1
-Larry wants to go out and watch the New Year’s fireworks, but the municipality does not disclose their location anymore to avoid crowding.
-There are three possible locations: main square, marina, and central park.
-There are three bus lines on his street: the red line goes to the square, the blue line to the marina, and the green line to the park.
-The buses always arrive in this order, at constant time intervals.
-Larry is late, so his only chance to see the fireworks is taking the first bus that arrives, regardless of its color, and hoping it goes to the right location.
-Suppose the blue line comes.
-When Larry arrives at the marina, he discovers whether the fireworks are there or not.
-If they are not, Larry regrets not moving to the bus stop a little earlier, which meant catching the red line: “I bet the fireworks are at the square!” Overhearing this, a bookmaker approaches Larry and proposes a bet on the fireworks’ location.
-Larry accepts a bet that yields $9$ if the fireworks are at the square and $-10$ if the fireworks are at the park.
-By the same token, at the park, he would accept a bet that yields $9$ for marina and $-10$ for square, and at the square, he would accept a bet that yields $9$ for park and $-10$ for marina.
-Thus, no matter the fireworks’ location, Larry ends up with equal probability in each of the other two locations, and hence he suffers an expected loss from this betting behavior.
-In the example, the bus schedule determines the objective conditional probabilities of the three contingencies of interest.
-Instead, there are no objective probabilities of the fireworks’ location, and Larry does not have a subjective prior either; after all, he cannot decide which bus to take, so he does not need one.
-However, to deal with the bookmaker, he needs to form a belief after discovering that the fireworks are not in the location he arrived to.
-What’s wrong with the beliefs he forms?
-The location he deems more likely is different in each of the three contingencies.
-This form of incoherence among beliefs can arise precisely from the lack of a prior.
-Thus, the Dutch-book argument provides a motivation for having a prior, even when there is no decision to make before the arrival of information.
-More generally, the argument provides a justification for having coherent beliefs that can be rationalized with a prior, although the agent may not truly have one, but just reason in a coherent way after different discoveries.
-In the population interpretation, having a prior translates into the agents having the same prior.
-Thus, the Dutch book arguments suggests a disadvantage for “divided” populations, in which different agents have different (prior) beliefs, in that the agents with different priors can be exploited in different situation.
-Such divided population naturally emerges in a repeated setting if the agents do not share their experiences, thus suggesting an important role for the aggregation of feedback.
-Another possible reason for Larry’s incoherent beliefs is that he (has a prior, but) perceives an inexistent correlation between bus arrivals and location of the fireworks.
-This is a classic form of model misspecification about the likelihood function connecting fundamental states and contingencies.
-In fact, we show that any form of model misspecification about the likelihood function allows Dutch-booking.
-This is perhaps surprising: since there is no “bad prior” over fundamental states (because there is no objective one), it is hard to conceive in what sense updating the prior with wrong likelihoods could generate a “bad posterior” over fundamental states.
-The Dutch-book argument offers a precise sense in which model misspecification is bad for decision-making.
-In this simple example, having one prior and updating it by conditioning suffices to avoid Dutch-booking.
-But this is not true in general.
-We determine the coherence rule among beliefs that is necessary and sufficient to avoid Dutch-booking in objective expected terms.
-The coherence rule is easy to interpret; it is based on the idea that differences in the subjective odds ratios of states across contingencies shall only reflect differences in the objective odds ratios of the contingencies given the states.
-In the example, if Larry deems square twice more likely than park when he is at the marina and marina twice more likely than square when he is at the park, to satisfy the coherence rule he should deem marina four times more likely than park when he is at the square.
-In turn, we show that this coherence rule characterizes a Completely Consistent Belief System: an array of beliefs (one for each contingency) that can be derived from a Lexicographic Conditional Probability System (Blume et al., 1991), a list of theories in a priority order.
-Specifically, each belief is derived with Bayes rule from the first theory that is consistent with the contingency.
-So, the LCPS plays the role of lexicographic prior, while the CCBS specifies the posterior beliefs in different contingencies.
-The fact that a completely consistent belief rather than simply conditioning from a unique prior is needed will be illustrated in Section 6 with an example of an agent who is subject to correlation neglect.
+Third, an agent may be offered the whole set of conditional bets at the ex-ante stage, as in a speculative-trade context (see the related literature below).
+Traditional Dutch-book arguments are often criticized for the possibility that the agent would realize being Dutch-booked when all bets are offered simultaneously.
+However, in our case, a Dutch-booked agent will still compute a positive expected payoff under some state by using a misspecified likelihood function.
+[^6] Dutch-book arguments have also been criticized for the need of a bookmaker who knows the agent’s beliefs.
+We disagree with this critique, as a bookmaker may find a Dutch book that the agent wants to accept by trial and error.
+In a financial market, the agent may also self-pick a set of assets that jointly constitute a Dutch book, without the presence of a strategic bookmaker.
+We then apply our theorems to one commonly studied form of misspecification, correlation neglect.
+Since agents with this bias update beliefs using Bayes rule, our first result implies that they are shielded from deterministic Dutch books.
+However, they use an incorrect likelihood that neglects the correlation between contingencies, thereby exposing them to general Dutch books, consistent with our second result.
+In particular, we illustrate the nature of those Dutch books in a venture capital setting and in a variation of the Monty Hall game, with secretly assigned doors.
+In the former application, the Dutch book consists of the venture capitalist offering different combinations of convertible debt and ownership purchases depending on whether initial milestones are achieved.
+In the second case, the Dutch book offers terms that seem inviting if the remaining doors were equally likely, but that induce sure expected loss under the mechanism used to determine what door is open initially.
 Related literature
@@
 CCBSs guarantee that all conditional preferences can be derived from a prior lexicographic preference relation.
-In other words, they guarantee a stronger form of Bayesianism which might be called “Bayesian updating of preferences” compared to Bayesian updating of beliefs.
-In Siniscalchi (2022), as well as Battigalli et al.
-(2023), a CCBS is defined as an array of beliefs, one for each information set, that is consistent with a Complete Conditional Probability System: an array of probability measures, one for each subset of opponents’ strategies, disciplined by the chain rule of probability.
+In other words, they guarantee a stronger form of Bayesianism, which might be called “Bayesian updating of preferences” rather than “Bayesian updating of beliefs.” In Siniscalchi (2022), as well as Battigalli et al.
+(2023), a CCBS is defined as an array of beliefs, one for each information set, that is consistent with a Complete Conditional Probability System (CCPS): an array of probability measures, one for each subset of opponents’ strategies, disciplined by the chain rule of probability.
 [^8] Conditional probability systems were first introduced by Renyi (1955).
 CCPSs were introduced in the analysis of games by Myerson (1986).
-In turn, a CCPS can be derived from a LCPS by conditioning.
-In a sequential game, a player progressively rules out opponents’ strategies by observing their moves, that is, learning occurs in the partitional fashion.
+In turn, a CCPS can be derived from a lexicographic probability system by conditioning.
+In a sequential game, a player progressively rules out opponents’ strategies by observing their moves; that is, learning occurs in a partitional fashion.
 So, Bayes rule boils down to conditioning, and hence the definition of CCBS in Siniscalchi (2022) and Battigalli et al.
 (2023) is a special case of the definition of this paper.
-Thus, the Dutch-book argument for complete consistency implies that the standard notion of conditional probability system defined on the collection of observable events falls short of the level of coherence that shelters from Dutch-booking in objective expected terms.
-The working paper of Siniscalchi (2022) also characterizes CCBSs with a generalization of the chain rule that applies between non-nested conditional events.
+Thus, the Dutch-book argument for complete consistency implies that the standard notion of a conditional probability system defined on the collection of observable events falls short of the level of coherence that shelters from Dutch-booking in objective expected terms.
+The working paper of Siniscalchi (2022) also characterizes CCBSs by generalizing the chain rule to apply to non-nested conditional events.
 In the Appendix, we provide a direct proof of the equivalence between the generalized chain rule and our coherence rule in sequential games.
 Pettigrew (2023) considers an agent with a credence at time $t$ and a vector of possible credences at $t^{\prime}>t$, and shows that the agent can be Dutch-booked if and only if his credence at $t$ is not in the convex hull of the credences at $t^{\prime}$ (which he calls “General Reflection Principle”).
-This result yields our equivalence between Bayes-plausibility and the impossibility of deterministic Dutch-booking for the special case of a tree with depth $1$ and discoveries that do not rule out any state.
+This result yields our equivalence between Bayes plausibility and the impossibility of deterministic Dutch-booking for the special case of a tree with depth $1$ and discoveries that do not rule out any state.
 [^9] When some discoveries objectively rule out some states, statement II.a of Theorem 1 in Pettigrew (2023) fails.
 From his theorem, Pettigrew also derives an argument for conditioning in the partitional framework.
-In our richer framework, we can derive our more general argument for forward consistency, uncovering the limited possibility of achieving a deterministic Dutch-booking against likelihood misspecification.
+In our richer framework, we can derive our more general argument for forward consistency, establishing the impossibility of achieving a deterministic Dutch-booking against likelihood misspecification.
 Chen et al.
@@
 [^10] See also Lehrer and Samet (2011) for an alternate characterization of “agreeing to agree” between only two agents.
-Their notion of Dutch book only takes the realized discoveries and sums up the payouts under each state.
+Their notion of a Dutch book considers only realized discoveries and sums up the payouts under each state.
 This amounts to assuming that the bookmaker can choose the bets based on the entire set of realized discoveries.
-So, when the agents have different (thus not commonly known) posteriors, the bookmaker can exploit an informational advantage over each single agent, who is unaware of the other agents’ discoveries.
-In other words, the bookmaker can make profit from arbitrage.
-This is ruled out by our notion of Dutch book, in which each possible discovery is associated with a bet, independently of what other discoveries (using the population interpretation) may have been made by other agents.
+So, when the agents have different (and thus not commonly known) posteriors, the bookmaker can exploit an informational advantage over each agent, who is unaware of the other agents’ discoveries.
+In other words, the bookmaker can make a profit from arbitrage.
+This is ruled out by our notion of a Dutch book, in which each possible discovery is associated with a bet, independently of what other discoveries (using the population interpretation) may have been made by other agents.
 This reconciles the possibility of Dutch-booking Bayesian agents in Chen et al.
 (2015) with the impossibility results of this paper.
-There are several related contribution in the literature on misspecified Bayesian learning, who is excellently surveyed in Bohren and Hauser (2024a).
-Lanzani (2025) characterizes a behavior rule that prevents misspecification to damage the agent beyond his maxmin payoff preserving consistency when correctly specified.
-Bohren and Hauser (2024b) characterizes when a departure from Bayesian updating (e.g., underinference from signals) can be rationalized as the consequence of Bayesian updating within a misspecified model.
-Molavi (2024) shows that beliefs are consistent with Bayesian updating (from a possibly misspecified model) if and only if the mean posterior is absolutely continuous with respect to the prior.
-This finding generalizes the earlier Shmaya and Yariv (2016) to the case in which the subjective beliefs of the decision maker may not have the same support of the true data-generating process.
+Molavi (2025) shows that beliefs are consistent with Bayesian updating if and only if the mean posterior is absolutely continuous with respect to the prior.
+This finding generalizes the earlier work of Shmaya and Yariv (2016) to the case in which the decision maker’s subjective beliefs may not have the same support as the true data-generating process.
+Fudenberg and Lanzani (2026) characterize when beliefs are consistent with Bayesian updating from a conditional i.i.d.
+model.
+There are several related contributions in the literature on misspecified Bayesian learning, which is excellently surveyed in Bohren and Hauser (2025) and Esponda and Pouzo (2026).
+Here we discuss the part of this literature concerned with measuring the loss induced by misspecification and how to avoid these losses.
+Lanzani (2025) characterizes a behavior rule that prevents misspecification from damaging the agent beyond his maxmin payoff, preserving consistency when correctly specified.
+Frick et al.
+(2024) provide a rank of forms of misspecification on the basis of the damages they induce, and Gossner and Steiner (2018) provides a measure of cost of misperceiving the payoff function.
+Bohren and Hauser (2024) characterize the conditions under which a departure from Bayesian updating (e.g., underinference from signals) can be rationalized as a consequence of Bayesian updating within a misspecified model.
 Cerreia-Vioglio et al.
-(2025) show that in coordination games misspecification can be extremely damaging even for misspecification concerned agents.
+(2025) show that misspecification in coordination games can be highly damaging, even for agents concerned about it.
 Compared to this literature, we provide a novel reason for why some misspecification should or should not be expected to persist based on their susceptibility to Dutch books.
+Moreover, as our application to correlation neglect clearly shows, the form of misspecification that are susceptible to Dutch book significantly differ from the one that have been deemed unstable due to light-bulb realizations or evolutionary pressure (see, Gagnon-Bartsch et al.
+2023, Fudenberg and Lanzani 2023, Ba 2026, He and Libgober 2025).
 Finally, our results are similar in spirit to those in the no-trade literature, the closest being in Morris (1994).
 Morris considers agents who trade claims over commodities that are contingent on a payoff-relevant state and on the realization of a signal.
-In the case of two traders, a single money commodity and risk neutrality, Morris’ model coincides with ours given a tree of depth $1$.
-Assuming that the traders have no incentive to exchange claims that are contingent on the state only (which in our setting boils down to having the same prior), Theorem 2.1 in Morris (1994) shows that there exists trade that gives a positive expected payoff to each trader if and only if they disagree on the likelihood function mapping states into probabilities over signals.
+In the case of two traders, a single money commodity, and risk neutrality, Morris’ model coincides with ours with a tree of depth $1$.
+Assuming that the traders have no ex-ante incentive to exchange claims that are contingent on the state only (which in our setting boils down to having the same prior), Theorem 2.1 in Morris (1994) shows that there exists trade that gives a positive expected payoff to each trader if and only if they disagree on the likelihood function mapping states into probabilities over signals.
 Our main theorem further shows (in a more general learning framework) that such trade can take a very specific form, whereby one trader has a positive expected payoff conditional on each state, and the other has a positive expected payoff conditional on each signal.
-If the first trader (the bookmaker) is using the correct likelihood function, then the second trader (our agent) gets an objectively negative expected payoff conditional on each state, and thus is Dutch-booked.
+Suppose the first trader (the bookmaker) is using the correct likelihood function.
+In that case, the second trader (our agent) gets an objectively negative expected payoff conditional on each state, and thus is Dutch-booked.
 Conversely, we show that the existence of such trade guarantees that the disagreement between the two traders must be about the likelihood function, not (just) the prior.
-The paper is organized as follows.
-Section 1 introduces the framework.
-Section 2 illustrates belief systems.
-Section 3 introduces the novel notions of Dutch book.
-Section 4 states and proves the Dutch-book theorem for forward-consistency.
-Section 5 is devoted to the Dutch-book theorem for complete consistency.
-In the Appendix, we formalize the special case of sequential games and the connection with Siniscalchi (2022).
 2 Framework
@@
 We will refer to this as the “independence case”.
-To illustrate the framework, we formalize the example of the Introduction.
+We now introduce an example that we will repeatedly use to illustrate our contribution.
 Example 1
-There are three possible states, corresponding to the location of the fireworks: $\mathcal{S=}\left\{sq,ma,pa\right\}$.
-There are six possible contingencies, each with a different set of consistent states, so we label each contingency $h$ with the only state or with the initials of the two states in $S(h)$:
-=$\mathcal{H=}\left\{sq,ma,pa,sm,mp,ps\right\}.$
-For instance, if Larry takes the bus that goes to the square, he ends up in contingency $sq$ if the fireworks are there, and in contingency $mp$ otherwise.
-Note that all the contingencies are unordered, therefore contingencies and learning paths coincide.
-The arrival process of the buses is independent of the state.
-Thus, we are in the independence case: for each $h\in\left\{sm,mp,ps\right\}$ and $s,s^{\prime}\in S(h)$, $p(h|s)$ and $p(h|s^{\prime})$ are equal, because they coincide with the probability of the bus that leads to contingency $h$.
-In particular, since the three lines arrive with equal frequencies, each state $s\in\mathcal{S}$ induces a uniform distribution $\eta^{s}$ over the three learning paths/contingencies $h$ such that $s\in S(h)$.
-So, $p(h|s)=p(h|s^{\prime})=1/3$.
+Larry wants to go out and watch the New Year’s fireworks, but the municipality does not disclose their location anymore to avoid crowding.
+There are three possible locations: the main square, the marina, and the central park, $\mathcal{S=}\left\{sq,ma,pa\right\}$.
+Three bus lines run on his street: the red line goes to the square, the blue to the marina, and the green to the park.
+Then, after taking a bus (say the red bus), Larry may find himself in one of two contingencies: the one in which he arrives at the correct location ($sq$) and the one in which the fireworks are elsewhere ($mp$).
+So, without introducing an initial contingency, we have $\mathcal{H=}\left\{sq,ma,pa,sm,mp,ps\right\}$, where each contingency $h$ is labeled with the only state or with the initials of the two states $S(h)$ that are consistent with it.
+Since the contingencies are all unordered, the learning paths have length $1$ and coincide with them.
+Larry is late, so his only chance to see the fireworks is to take the first bus that arrives, regardless of the line, and hope it goes to the correct location.
+Suppose that the three lines arrive in the given order at constant time intervals, independently of the state.
+Then, each state $s\in\mathcal{S}$ induces a uniform distribution $\eta^{s}$ over the three learning paths/contingencies $h$ such that $s\in S(h)$, and we are in the independence case: for each $h\in\left\{sm,mp,ps\right\}$ and $s,s^{\prime}\in S(h)$, $p(h|s)=p(h|s^{\prime})=1/3$.
 $\triangle$
 3 Belief systems
-We assume that, in each contingency $h$, the agent has a probabilistic belief $\mu(\cdot|h)\in\Delta(\mathcal{S})$ such that $\mu(S(h)|h)=1$ — we call $(\mu(\cdot|h))_{h\in\mathcal{H}}$ belief system.
-This maintained assumption, besides requiring probabilistic sophistication at any point in time, rules out the most severe forms of model misspecification in which an agent assigns positive probability to states that have become objectively impossible.
+We assume that, in each contingency $h$, the agent has a probabilistic belief $\mu(\cdot|h)\in\Delta(\mathcal{S})$ such that $\mu(S(h)|h)=1$— we call $(\mu(\cdot|h))_{h\in\mathcal{H}}$ a belief system.
+This maintained assumption, besides requiring probabilistic beliefs at any point in time, rules out the most severe forms of model misspecification in which an agent assigns positive probability to states that have become objectively impossible.
 We say that a belief system is forward-consistent if the agent’s beliefs are consistent with Bayesian updating from his previous belief using some (conditional) likelihood function, whenever possible.
-Observe that such conditional likelihood function is not required to be the objectively correct one.
+Observe that such a conditional likelihood function is not required to be the objective one.
 Definition 1
 A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is forward-consistent if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist conditional probability distributions
-$\left(\widetilde{p}(\cdot|h,s)\right)_{s\in S(h)}\in\left(\Delta(\left\{h^{1},...,h^{n}\right\})\right)^{\left|S(h)\right|}$
-such that, for each $m\in\left\{1,...,n\right\}$, if $\widetilde{p}(h^{m}|h,s)>0$ for some $s\in\mathrm{supp}\mu(\cdot|h)\subseteq S(h)$, then $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$.
-It is well-known that beliefs are consistent with this form of updating if and only if the “prior belief” (i.e., the belief at $h$) is a convex combination of the “posterior beliefs” (i.e., the beliefs at $h^{1},...,h^{n}$).
+$\left(\widetilde{p}(\cdot|h,s)\right)_{s\in S(h)}\in\left(\Delta(\left\{h^{1},...,h^{n}\right\})\right)^{S(h)}$
+such that, for each $m\in\left\{1,...,n\right\}$, if $\widetilde{p}(h^{m}|h,s)>0$ for some $s\in\mathrm{supp}\mu(\cdot|h)$, then $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$.
+It is well known that beliefs are consistent with this form of updating if and only if the “prior belief” (i.e., the belief at $h$) is a convex combination of the “posterior beliefs” (i.e., the beliefs at $h^{1},...,h^{n}$).
 (The Appendix contains a formal proof of this fact.)
 Definition 2
-A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is Bayes-plausible if, for every non-terminal $h\in\mathcal{H}$, calling $h^{1},...,h^{n}$ its immediate successors, there exist $\alpha^{1},...,\alpha^{n}\in\left[0,1\right]$ such that $\alpha^{1}+...+\alpha^{n}=1$ and
-$\sum_{m=1}^{n}\alpha^{m}\mu(\cdot|h^{m})=\mu(\cdot|h).$
-As we will show, Bayes-plausibility is the minimal requirement of coherence among beliefs that shelters from deterministic Dutch-booking.
+A belief system $(\mu(\cdot|h))_{h\in\mathcal{H}}$ is Bayes-plausible if, for every non-terminal $h\in\mathcal{H}$ with immediate successors $h^{1},...,h^{n}$, there exist $\alpha_{1},...,\alpha_{n}\in\left[0,1\right]$ such that
+$\sum_{m=1}^{n}\alpha_{m}=1\quad\text{and\qquad}\sum_{m=1}^{n}\alpha_{m}\mu(\cdot|h^{m})=\mu(\cdot|h).$
+As we will show, Bayes-plausibility is the minimal requirement of coherence among beliefs that shields against deterministic Dutch-booking.
 We will define a Completely Consistent Belief System as a belief system that can be derived from a Lexicographic Conditional Probability System.
@@
 $\mu(s|h)=\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in S}p(h|s^{\prime})\mu^{m}(s^{\prime})}\qquad\forall s\in S(h)\text{.}$
-[…]
+In a CCBS, given each contingency $h$, the belief $\mu(\cdot|h)$ is derived with Bayes rule from the first theory in the LCPS that can explain $h$, i.e., that assigns positive probability to some $s\in S(h)$.
+[^11] Ortoleva (2012) proposes and axiomatizes a generalization of LCPS that allows the agent to change prior after small but positive probability event in a way that better matches the empirical evidence about belief revisions.
+Moreover, Bayes rule must be applied using the objective likelihood $p$.
+Note that, for any two contingencies $h,h^{\prime}$ such thatS(h)=S(h)\ $S(h^{\prime})$, the beliefs derived from the LCPS with Bayes rule may differ.
+However, they are derived from the same measure in the LCPS.
+This is because a contingency may convey more information than the set of consistent states, as it may be reached with different probabilities under different states.
+When this is not the case, i.e., in the independence case, Bayes rule boils down to conditioning and Definition 4 coincides with the definition of CCBS in Siniscalchi (2022) and Battigalli et al.
+(2023) — see the Appendix for details.
 We will characterize CCBSs in terms of coherence among the subjective relative probabilities of states in different contingencies.
-However, given that the objective probabilities of contingencies may differ across states, these comparisons are meaningful only after discounting the probability attributed to a state in a contingency by the probability that the state induces the contingency.
+However, because the objective probabilities of contingencies may differ across states, these comparisons are meaningful only after discounting the probability attributed to a state in a contingency by the probability that the state induces the contingency.
 Definition 5
@@
 $o(s,s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{\prime}|h^{n})$
-such that, if one discounted odds ratio is 00, no discounted odds ratio is infinite.
+such that, if one discounted odds ratio is 0, no discounted odds ratio is infinite.
 A discounted odds ratio compares the probabilities attributed to two states in the same contingency.
@@
 We say that a belief system is completely coherent when, for every $s,s^{\prime}\in\mathcal{S}$, the generalized odds ratios between $s$ and $s^{\prime}$ are all identical.
-We illustrate the notions of belief system and generalized odds ratio through the example from the Introduction.
+We illustrate the notions of belief system and generalized odds ratio through the running example.
 Example 2
-For each state $s$, a belief system $\mu=\left(\mu(\cdot|h)\right)_{h\in\mathcal{H}}$ must assign probability $1$ to $s$ in the contingency labelled with $s$.
-Also, in this example, every belief system is trivially forward-consistent as every path consists of just one (terminal) contingency.
 To obtain a CCBS, let us fix a LCPS $\bar{\mu}=(\mu^{1},...,\mu^{n})$.
-Suppose that $\mu^{1}$ assigns positive probability to at least $2$ states.
-Then, $\mu^{1}(S(h))>0$ for each $h\in\left\{sm,mp,ps\right\}$.
-Because each contingency has the same probability under the two consistent states, we are in the independent case: Bayes rule boils down to the chain rule, and discounted odds ratios boil down to odds ratios.
-So, for each $h\in\left\{sm,mp,ps\right\}$, writing $S(h)=\left\{s,s^{\prime}\right\}$, we have
-$\mu(s|h)=\frac{\mu^{1}(s)}{\mu^{1}(s)+\mu^{1}(s^{\prime})}.$
-If instead $\mu^{1}$ assigns probability $1$ to a state $\bar{s}$, we obtain $\mu(\bar{s}|h)=1$ whenever $\bar{s}\in S(h)$, and the belief in the contingency $h$ such that $S(h)=\mathcal{S}\backslash\left\{\bar{s}\right\}=\left\{s,s^{\prime}\right\}$ is given by
-$\mu(s|h)=\frac{\mu^{2}(s)}{\mu^{2}(s)+\mu^{2}(s^{\prime})}.$
+Because each contingency has the same probability under the two consistent states (i.e., we are in the independence case), Bayes rule reduces to the chain rule, and discounted odds ratios reduce to odds ratios.
+So, from $\bar{\mu}$, we can derive a belief system $\mu=\left(\mu(\cdot|h)\right)_{h\in\mathcal{H}}$ as follows.
+For each $h\in\left\{sq,ma,pa\right\}$, $\mu(s|h)=1$ for $s=h$.
+Next, suppose first that $\mu^{1}$ assigns positive probability to at least $2$ states.
+Then, for each $h\in\left\{sm,mp,ps\right\}$, $\mu^{1}(S(h))>0$, and we get
+$\mu(s|h)=\frac{\mu^{1}(s)}{S(h)}\qquad\forall s\in S\left(h\right).$
+If instead $\mu^{1}$ assigns probability $1$ to a state $\bar{s}$, we obtain $\mu(\bar{s}|h)=1$ whenever $\bar{s}\in S(h)$, and the belief in any contingency $h$ such that $\bar{s}\notin S(h)$ is given by
+$\mu(s|h)=\frac{\mu^{2}(s)}{\mu^{2}(S(h))}\qquad\forall s\in S\left(h\right).$
 Given any two states $s,s^{\prime}$, one can compute an odds ratio between $s$ and $s^{\prime}$ in the only contingency $h$ such that $S(h)=\left\{s,s^{\prime}\right\}$; call it $o(s,s^{\prime})$.
-Then, one can also compute a generalized odds ratio between $s$ and $s^{\prime}$ as $o(s,s^{\prime\prime})\cdot o(s^{\prime\prime},s^{\prime})$ (s′′{$s^{\prime\prime}\notin\left\{s,s^{\prime}\right\}$), unless one is zero and the other infinite.
+Then, one can also compute a generalized odds ratio between $s$ and $s^{\prime}$ as $o(s,s^{\prime\prime})\cdot o(s^{\prime\prime},s^{\prime})$ ($s^{\prime\prime}\notin\left\{s,s^{\prime}\right\}$), unless one is zero and the other infinite.
 For instance, when $\mu^{1}(ma)\in(0,1)$, we can compute
@@
 which can be infinite if $\mu^{1}(pa)=0$ and thus $\mu^{1}(sq)>0$, or zero, vice versa. Note the equality between the odds ratio and the generalized odds ratio.
-An alternative belief system $\bar{\mu}$, which rationalizes Larry’s betting behavior, is the following:
-$\bar{\mu}(sq|ps)=\bar{\mu}(ma|sm)=\bar{\mu}(pa|mp)=3/4\text{.}$
-In this case, the odds ratio and the generalized odds ratio between any two states differ.
-For instance, in contingency $ps$ we have $o(sq,pa)=3$, but in contingency $sm$ we have $o(sq,ma)=1/3$ and in contingency $mp$ we have $o(ma,pa)=1/3$, therefore we can compute a generalized odds ratio between $sq$ and $pa$ of $1/9$ instead of $3$triangle
+Suppose now that, after not finding the fireworks, Larry regrets not going down to the bus stop a little earlier and believes that, most likely, the previous bus would have brought him to the right location.
+In this case, he could have the following belief system $\hat{\mu}$:
+$\hat{\mu}(sq|sq)$ $=$ $\hat{\mu}(ma|ma)=\hat{\mu}(pa|pa)=1,$
+$\hat{\mu}(sq|ps)$ $=$ $\hat{\mu}(ma|sm)=\hat{\mu}(pa|mp)=3/4\text{.}$
+Under this belief system, the odds ratio and the generalized odds ratio between any two states differ.
+For instance, in contingency $ps$ we have $o(sq,pa)=3$, but in contingency $sm$ we have $o(sq,ma)=1/3$ and in contingency $mp$ we have $o(ma,pa)=1/3$, therefore we can compute a generalized odds ratio between $sq$ and $pa$ of $1/9$ instead of $3$.
+Thus, $\hat{\mu}$ is not completely coherent.
+Every belief system is trivially forward-consistent, as every path consists of just one (terminal) contingency.
+If we introduce a belief of Larry before taking the bus, i.e., an initial contingency $h^{0}$, then Bayes-plausibility is guaranteed by the condition that for each $h\in\left\{sq,ma,pa\right\}$, $\mu(h|h)=1$, as every initial belief would trivially be a convex combination of these three vertices of $\Delta\left(\mathcal{S}\right)$.
+Thus, every belief system would still be forward-consistent.
+$\triangle$
 4 Dutch books and Dutch-booking
@@
 An agent is Dutch-booked if he accepts all the bets in a Dutch book.
-Our agent, however, would never accept a classical Dutch book if all the bets are proposed at the same time, as he has a well-defined probabilistic belief at any point in time.
+Our agent, however, would never accept a classical Dutch book if all the bets are proposed simultaneously, as he has a well-defined probabilistic belief at any point in time.
 We now investigate whether, instead, a system of gambles proposed in different contingencies can induce the same kind of state-wise losses.
-[^9] Our agent risk neutrality and probabilistic sophistication at every contingency imply that, when he is willing to accept a bet, he is also willing to accept the sum of the bets, and thus we can assume without loss of generality that our agent is proposed just one bet in each contingency.
+[^13] Our agent’s risk neutrality and probabilistic sophistication at every contingency imply that, when he is willing to accept a bet, he is also willing to accept the sum of the bets.
+Thus, we can assume, without loss of generality, that our agent is offered only one bet in each contingency.
 Definition 7
-A gamble is a map $\gamma:\mathcal{S}\rightarrow\mathbb{R}$ that specifies, for each $s\in\mathcal{S}$, the gain or loss $\gamma(s)$ for the agent in case of $s$.
-We say that $\gamma$ is a gamble on $\widetilde{S}\subseteq\mathcal{S}$ if $\gamma(s)=0$ for every $s\in\mathcal{S}\backslash\widetilde{S}$.
-A system of gambles is an array $g=(g(\cdot|h))_{h\in\mathcal{H}}$ where, for each $h\in\mathcal{H}$, $g(\cdot|h)$ is a gamble on $S(h)$.
-[^10] The results would not change if the bets in a Dutch book were allowed to yield non-zero payoffs also under objectively impossible states.
+A gamble is a map $\gamma:\mathcal{S}\rightarrow\mathbb{R}$ that specifies, for each $s\in\mathcal{S}$, the gain or loss $\gamma(s)$ for the agent if $s$ realizes.
+A system of gambles is an array $g=(g(\cdot|h))_{h\in\mathcal{H}}$ where, for each $h\in\mathcal{H}$, $g(\cdot|h)$ is a gamble.
 What does it mean, for a system of gambles, to be a “Dutch book”?
@@
 We call this notion “deterministic Dutch book”.
-But the existence of objective conditional probabilities of paths allows to introduce also a more permissive notion of Dutch book, whereby the bookmaker obtains a non-negative/positive profit conditional on each funamental state only in objective expected terms.
+But the existence of objective conditional probabilities of paths allows us to introduce also a more permissive notion of Dutch book, whereby the bookmaker obtains a non-negative/positive profit conditional on each fundamental state only in objective expected terms.
 Definition 8
@@
 $\sum_{h\in\mathcal{H}}p(h|s)g(s|h)<0.$ (2)
-A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a deterministic Dutch book if for every $s\in\mathcal{S}$ and $l\in L\left(s\right)$
-$\sum_{h\in H(l)}g(s|h)\leq 0,$ (3)
+A system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a deterministic Dutch book if
+$\sum_{h\in H(l)}g(s|h)\leq 0,\qquad\forall s\in S,\forall l\in L\left(s\right)$ (3)
 and for some $s\in\mathcal{S}$ and $l\in L\left(s\right)$,
 $\sum_{h\in H(l)}g(s|h)<0.$ (4)
-Note that a deterministic Dutch book does not depend on the objective probabilities of paths, hence it is a Dutch book no matter such probabilities.
+Note that a deterministic Dutch book does not depend on the objective probabilities of paths, hence it is a Dutch book no matter what such probabilities are.
 Moreover, Dutch books, deterministic or not, do not depend on the objective probabilities of the states.
@@
 Example 3
-Suppose that Larry has the belief system $\bar{\mu}$ constructed before. The bookmaker constructs the following system of bets:
-$g(sq|ps)$ $=$ $g(ma|sm)=g(pa|mp)=9;$
-$g(pa|ps)$ $=$ $g(sq|sm)=g(ma|mp)=-10;$
-$g(s|h)$ $=$ 0 $0\text{ \ \ \qquad\qquad\qquad\qquad\qquad\qquad\qquad}\forall h\in\left\{sq,ma,pa\right\}.$
+A bookmaker follows Larry and, in case the bus does not arrive at the correct location, she is ready to propose a bet on the fireworks’ location.
+If the bus goes to the marina, she is thinking of offering a bet that yields $9$ if the fireworks are at the square and 10-10 if they are at the park; likewise at the square and at the park.
+Altogether, we have the following system of bets:
+$\hat{g}(sq|ps)$ $=$ $\hat{g}(ma|sm)=\hat{g}(pa|mp)=9;$
+$\hat{g}(pa|ps)$ $=$ $\hat{g}(sq|sm)=\hat{g}(ma|mp)=-10;$
+$\hat{g}(s|h)$ $=$ $0\qquad\forall h\in\left\{sq,ma,pa\right\}.$
 Under every state $s$, we have
-$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)=\frac{1}{3}(9-10+0)=-\frac{1}{3}<0,$
-i.e., Larry has an objective expected payoff of $-1/3$. Therefore, the system is a Dutch book. There remains to show that Larry is willing to accept it. We have
-$\sum_{s\in\mathcal{S}}\bar{\mu}(s|h)g(s|h)=\frac{3}{4}\cdot 9+\frac{1}{4}(-10)+0\cdot 0>0\quad\forall h\in\left\{sm,mp,ps\right\}.$
-Therefore, Larry is Dutch-booked.
-Note that deterministic Dutch-booking is impossible in this example, as every path consists of just one contingency, in which the agent has a well-defined probabilistic belief.
+$\sum_{h\in\mathcal{H}}p(h|s)\hat{g}(s|h)=\frac{1}{3}(9-10+0)=-\frac{1}{3}<0,$
+so, this is a Dutch book.
+If the bookmaker can also propose a bet to Larry before the bus arrives, she may try to deterministically Dutch-book him.
+However, for him to accept this bet in contingency $h^{0}$, she would need to offer a positive prize under at least one state $s$.
+But then, to obtain a deterministic Dutch book, she would need to offer a negative prize for $s$ in contingency $h=s$.
+Under the imposed condition that Larry “believes what he knows” (i.e., $\mu(s|h)=1$), Larry would reject the bet.
+Thus, Larry cannot be deterministically Dutch-booked.
+As already observed in Example 2, the same condition also guarantees Bayes plausibility/forward consistency.
+This is not a coincidence, and the argument is general: If, in a tree of depth 2, the agent has the possibility of learning the state (whichever it is), his beliefs are always forward-consistent, and he cannot be deterministically Dutch-booked.
+In the next section, we will show that the tight connection between deterministic Dutch-booking and forward-consistency extends beyond this special case.
 $\triangle$
@@
 In this sense, misspecification about the relative probabilities of the likelihood function is immune to deterministic Dutch-booking.
-However, that is not the case for the stronger form of misspecification where at some contingency $h$ some objectively impossible state $s\notin S\left(h\right)$ is deemed possible by the decision maker.
+However, that is not the case for the stronger form of misspecification where in some contingency $h$ some objectively impossible state $s\notin S\left(h\right)$ is deemed possible by the decision maker.
 This is because, by assigning an arbitrarily high positive payout $g(s|h)>>0$ to the agent in $s$ given $h$, the bookmaker can obtain a profit under some state $s^{\prime}\in S\left(h\right)$ that is objectively possible.
+Before proving the theorem, we draw the conclusions of the running example concerning deterministic Dutch booking.
+Example 4
+As anticipated, even if we introduced an initial belief for Larry, his beliefs would certainly be forward-consistent/Bayes-plausible, and he could not be deterministically Dutch-booked.
+So, to introduce the possibility of deterministic Dutch-booking, assume the following version of “Murphy’s law”: the arriving bus never goes to the fireworks location.
+Then, only the three contingencies $ps,sm,mp$ are possible after $h^{0}$.
+Suppose that, at $h^{0}$, Larry assigns probability $2/3$ to marina and $1/6$ to each of the other two states.
+However, he has uniform beliefs over the two possible states in each of the three following contingencies.
+Then, his beliefs do not satisfy Bayes-plausibility, as he assigns to marina a larger probability at $h^{0}$ than in any other future contingency.
+Indeed, Larry can be deterministically Dutch-booked.
+At $h^{0}$, the bookmaker can offer him a bet that pays $x+\varepsilon$ for marina and 2x-2x for park and square.
+Then, at each $h\in\left\{sm,mp\right\}$, she can offer $x-2\varepsilon$ for marina and $2x$ for the other possible state.
+(Complete the system of bets with the null bet at $ps$.) For sufficiently small $\varepsilon$, Larry will accept.
+It is easy to check that the system of bets is a deterministic Dutch book.
+[^14] Note that, if we reintroduce the three contingencies where Larry learns the state and associate them with null bets, the system of bets would no longer be a deterministic Dutch book (consistently with what we observed in Example 3).
+The reason is that the possibility to learn state marina introduces the learning path $(h^{0},ma)$ along which Larry makes a positive payoff.
+$\triangle$
 Proof of Theorem 1.
-$\mathbf{1\Rightarrow 2)}$
-This is a known fact, however we provide formal proof in the Appendix.
-$\mathbf{2\Rightarrow 3)}$
-[…]
-$\mathbf{3\Rightarrow 1)}$
-We prove the counterpositive.
+$\mathbf{1\Rightarrow 2)}$ This is a known fact; however, we provide formal proof in the Appendix.
+$\mathbf{2\Rightarrow 3)}$ This can be easily derived as a consequence of Dutch-book arguments for conditionalization, and therefore a proof is deferred to the Appendix.
+$\mathbf{3\Rightarrow 1)}$ We prove the counterpositive.
 Suppose that $\mu$ is not Bayes-plausible.
-Thus, there exists a non-terminal contingency $h$ such that $\mu(\cdot|h)$ is not a convex combination of the beliefs at the next contingencies $h^{1},...,h^{n}$.
+Thus, there exists a non-terminal contingency $h$ such that $\mu(\cdot|h)$ is not a convex combination of the beliefs in the next contingencies $h^{1},...,h^{n}$.
 Let $A\in\mathbb{R}^{\left|\mathcal{S}\right|\times n}$ be the matrix of subjective probabilities $\mu(s|h^{j})$, and let $b\in\mathbb{R}^{\left|\mathcal{S}\right|}$ be the vector of subjective probabilities $\mu(s|h)$.
-So, for every transposed vector $x=(x^{1},...,x^{n})\in\left[0,1\right]^{n}$ such that $\sum_{m=1}^{n}x^{m}=1$, $Ax\not=b$.
-Moreover, for every $x\in\left[0,\infty\right)^{n}$ such that $\sum_{m=1}^{n}x^{m}\not=1$, $Ax\not=b$, because $Ax=b$ implies
-$x^{1}+...+x^{n}$ $=$
-$\sum_{j=1}^{n}\left(\sum_{s}\mu(s|h^{j})\right)x^{j}$ $=$
-$\sum_{s}\sum_{j=1}^{n}x^{j}\mu(s|h^{j})$ $=$
-$\sum_{s}\mu(s|h)$ $=$ $1$
+So, for every transposed vector $x=(x_{1},...,x_{n})\in\left[0,1\right]^{n}$ such that $\sum_{m=1}^{n}x_{m}=1$, $Ax\not=b$.
+Moreover, for every $x\in\left[0,\infty\right)^{n}$ such that $\sum_{m=1}^{n}x_{m}\not=1$, $Ax\not=b$, because $Ax=b$ implies
+$\sum_{m=1}^{n}x_{m}$ $=$ $\sum_{j=1}^{n}\left(\sum_{s\in\mathcal{S}}\mu(s|h^{j})\right)x_{j}$
+$=$ $\sum_{s\in\mathcal{S}}\sum_{j=1}^{n}x_{j}\mu(s|h^{j})=\sum_{s\in\mathcal{S}}\mu(s|h)=1.$
 Thus, there is no $x\geq 0$ such that $Ax=b$. But then, by Farkas lemma, there exists $y\in\mathbb{R}^{\left|\mathcal{S}\right|}$ such that
 $A^{T}y\geq 0\text{ and }b^{T}y<0.$
-For some $\varepsilon>0$, by $-b^{T}y>0$ the gamble with vector of prizes $-y-2\overrightarrow{\varepsilon}$ is acceptable at $h$, and by $A^{T}y\geq 0$ the gamble with vector of prizes $y+\overrightarrow{\varepsilon}$ is acceptable at each $h^{j}$.
-Complete the system with gambles with zero prizes for all states at every other contingency.
-We obtain a deterministic Dutch book: if one of the paths that go through $h$ realizes, and those includes at least one $l\in L(s)$ for some $s\in S(h)$, the profit of the bookmaker is $\varepsilon>0$ (satisfying 4), otherwise it is 00 (satisfying 3).
+For some $\varepsilon>0$, by $b^{T}y>0$ the gamble with vector of prizes $y-2\overrightarrow{\varepsilon}$ is acceptable at $h$, and by $A^{T}y\geq 0$ the gamble with vector of prizes $y+\overrightarrow{\varepsilon}$ is acceptable at each $h^{j}$.
+Complete the system with gambles with zero prizes for all states in every other contingency.
+We obtain a deterministic Dutch book: if one of the paths in $L\left(h\right)$ realizes (note that $L\left(h\right)$ contains at least some $l\in L(s)$ for some $s\in S\left(h\right)\neq\emptyset$) the profit of the bookmaker is $\varepsilon>0$ (satisfying (4)), otherwise it is 0 (satisfying (3)).
 $\blacksquare$
@@
 The Dutch-book theorem for complete consistency generalizes the insights from the example: incongruities among generalized odds ratios can be exploited for Dutch-booking and can only be displayed by a belief system that is not completely consistent.
-The theorem also shows the opposite implications: when there is no incongruity between generalized odds ratios, the agent’s belief system is completely consistent and he cannot be Dutch-booked.
+The theorem also shows the opposite implications: when there is no incongruity between generalized odds ratios, the agent’s belief system is completely consistent, and he cannot be Dutch-booked.
 Thus, coherence among generalized odds ratios fully characterizes complete consistency and impossibility of Dutch-booking.
 Theorem 2
-Fix a belief system $\mu$. The following are equivalent.
+Fix a belief system $\mu$. The following are equivalent:
   1. 1.
@@
     the agent cannot be Dutch-booked.
+Before proving the theorem, we conclude the running example.
+Example 5
+Suppose that Larry has the belief system $\hat{\mu}$ constructed before.
+As we observed, $\hat{\mu}$ is not completely coherent.
+Thus, by Theorem 2, $\hat{\mu}$ is not completely consistent, and Larry can be Dutch-booked.
+In particular, he is willing to accept Dutch book $\hat{g}$:
+$\sum_{s\in\mathcal{S}}\hat{\mu}(s|h)\hat{g}(s|h)=\frac{3}{4}\cdot 9+\frac{1}{4}(-10)+0\cdot 0>0\qquad\forall h\in\left\{sm,mp,ps\right\}.$
+Thus, no matter the fireworks’ location, Larry ends up with equal probability in each of the other two locations, and hence he suffers an expected loss from this betting behavior.
+The lack of complete consistency means that Larry’s beliefs cannot be derived from a (lexicographic) prior under the true DGP.
+This could be because Larry does not have a prior; after all, he does not need one, as he cannot choose which bus to take.
+Or, it could be that Larry updates a prior with a misspecified DGP, perceiving a correlation between bus arrivals and the state when, in fact, there is none.
+In Section 7, we will instead provide opposite examples of correlation neglect.
+$\triangle$
 Proof. The acronyms DOR and GOR will stand for discounted/generalized odds ratio.
-$\mathbf{1\Rightarrow 2)}$
-Note preliminarly that, by Condition 1, if a state $s$ has a DOR with another state $s^{\prime}$, all the DORs between $s$ and $s^{\prime}$ take the same value; [^12] Recall that an “indeterminate DOR” betwen $s$ and $s^{\prime}$ is just not considered a DOR, and likewise for GOR.
-We say that $s$ has a DOR with $s^{\prime}$ when there exists $h\in\mathcal{H}$ such that $s,s^{\prime}\in S(h)$ and $\mu(s|h)+\mu(s^{\prime}|h)>0$, the DOR between $s$ and $s^{\prime}$ being $\mu(s|h)/\mu(s^{\prime}|h)$.
+$\mathbf{1\Rightarrow 2)}$ Note preliminarily that, by Condition 1, if a state $s$ has a DOR with another state $s^{\prime}$, all the DORs between $s$ and $s^{\prime}$ take the same value; [^15] Recall that by definition an “indeterminate DOR” between $s$ and $s^{\prime}$ is just not considered a DOR, and likewise for GOR.
+We say that $s$ has a DOR with $s^{\prime}$ when there exists $h\in\mathcal{H}$ such that $s,s^{\prime}\in S(h)$ and $\mu(s|h)+\mu(s^{\prime}|h)>0$, the DOR between $s$ and $s^{\prime}$ being $\left(\mu(s|h)p(h|s^{\prime})\right)/\left(\mu(s^{\prime}|h)p(h|s)\right)$.
 Similarly, we say that $s$ has a GOR with $s^{\prime}$ when it is possible to compute at least one GOR between $s$ and $s^{\prime}$.
 call it $o(s,s^{\prime})$.
-Let $P^{1}$ be the set of all $s\in\mathcal{S}$ that have no zero GOR with any $s^{\prime}$.
-Let $P^{2}$ be the set of all $s\in\mathcal{S}\backslash P^{1}$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash P^{1}$.
-Let $P^{3}$ be the set of all $s\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$ that have no zero GOR with any $s^{\prime}\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$.
+Let $P^{1}$ be the set of all $s\in\mathcal{S}$ such that for all $s^{\prime}\in\mathcal{S}$, all the GORs between $s$ and $s^{\prime}$ are different from 0.
+Let $P^{2}$ be the set of all $s\in\mathcal{S}\backslash P^{1}$ such that for all $s^{\prime}\in\mathcal{S}\backslash P^{1}$, all the GORs between $s$ and $s^{\prime}$ are different from 0.
+Let $P^{3}$ be the set of all $s\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$ such that for all $s^{\prime}\in\mathcal{S}\backslash\left(P^{1}\cup P^{2}\right)$, all the GORs between $s$ and $s^{\prime}$ are different from 0.
 And so on.
-Claim 1: there exists $n\geq 1$ such that $(P^{1},...,P^{n})$ is a partition of $\mathcal{S}$, with non-empty elements.
-Proof: Let $n$ be the smallest non-negative integer such that $P^{n+1}=\emptyset$; it exists because $\mathcal{S}$ is finite and the sets are disjoint by construction.
-There remains to show that $\cap_{m=1}^{n}P^{m}=\mathcal{S}$ (which also implies that $n\geq 1$).
-Let $\widetilde{S}=\mathcal{S}\backslash\left(\cap_{m=1}^{n}P^{m}\right)$ and suppose by contraposition that $\widetilde{S}\not=\emptyset$.
+Claim 1
+There exists $n\geq 1$ such that $(P^{1},...,P^{n})$ is a partition of $\mathcal{S}$, with non-empty elements.
+Proof of the Claim.
+Let $n$ be the smallest non-negative integer such that $P^{n+1}=\emptyset$; it exists because $\mathcal{S}$ is finite and the sets are disjoint by construction.
+There remains to show that $\cup_{m=1}^{n}P^{m}=\mathcal{S}$ (which also implies that $n\geq 1$).
+Let $\widetilde{S}=\mathcal{S}\backslash\left(\cup_{m=1}^{n}P^{m}\right)$ and suppose by contraposition that $\widetilde{S}\not$$\emptyset$.
 The fact that $P^{n+1}=\emptyset$ implies that every $s\in\widetilde{S}$ has a zero GOR with some $s^{\prime}\in\widetilde{S}$.
@@
 Since every node is the start of an arrow, the graph has a cycle.
-So, the product of all the (zero) GORs along the cycle is a zero GOR between $s^{1}$ and itself.
+So, the product of all the (zero) GORs along the cycle is a zero GOR between some $s^{1}\in\widetilde{S}$ and itself.
 Inverting all the DORs in this GOR, we obtain an infinite GOR between $s^{1}$ and itself.
@@
 $\square$
-Claim 2: there exists a LCPS $\bar{\mu}=(\nu^{1},...,\nu^{n})$ such that, for each $m=1,...,n$, (a) $\mathrm{supp}\nu^{m}=P^{m}$ and (b), for every pair $\left(s,s^{\prime}\right)\in P^{m}\times P^{m}$ which has a DOR $o(s^{\prime},s)$, $\nu^{m}(s^{\prime})/\nu^{m}(s)=o(s^{\prime},s)$.
-Proof: For each $m=1,...,n$, we will construct a probability measure $\nu^{m}$ that satisfies (a) and (b); then, $(\nu^{1},...,\nu^{n})$ is an LCPS because, by Claim 1, $(P^{1},...,P^{n})$ partitions $\mathcal{S}$.
-Fix $m=1,...,n$ and construct a graph with nodes $P^{m}$ and connections between nodes that have a DOR.
+Claim 2
+There exists a LCPS $\bar{\mu}=(\nu^{1},...,\nu^{n})$ such that, for each $m\in\left\{1,...,n\right\}$, (a) $\mathrm{supp}\nu^{m}=P^{m}$ and (b), for every pair $\left(s,s^{\prime}\right)\in P^{m}\times P^{m}$ which has a DOR $o(s^{\prime},s)$, $\nu^{m}(s^{\prime})/\nu^{m}(s)=o(s^{\prime},s)$.
+Proof of the Claim.
+For each $m\in\left\{1,...,n\right\}$, we will construct a probability measure $\nu^{m}$ that satisfies (a) and (b); then, $(\nu^{1},...,\nu^{n})$ is an LCPS because, by Claim 1, $(P^{1},...,P^{n})$ partitions $\mathcal{S}$.
+Fix $m\in\left\{1,...,n\right\}$ and construct a graph with nodes $P^{m}$ and connections between nodes that have a DOR.
 For every connected component $\widetilde{P}\subseteq P^{m}$, pick a state $s$ arbitrarily, let $\widetilde{\nu}(s)=1$, and progressively let $\widetilde{\nu}(s^{\prime\prime})=\widetilde{\nu}(s^{\prime})\cdot o(s^{\prime\prime},s^{\prime})$ for every $s^{\prime\prime}\in\widetilde{P}$, where $o(s^{\prime\prime},s^{\prime})$ is a DOR between $s^{\prime\prime}$ and some $s^{\prime}\in\widetilde{P}$ for which $\widetilde{\nu}(s^{\prime})$ was already defined; by definition of $P^{m}$, $o(s^{\prime\prime},s^{\prime})$ is not zero and not infinite (otherwise $o(s^{\prime},s^{\prime\prime})$ would be zero).
@@
 To show that $\nu^{m}$ satisfies the desiderata, fix $s,s^{\prime}\in P^{m}$ that have a DOR $o(s^{\prime},s)$, thus $s$ and $s^{\prime}$ are in the same connected component.
-Then, the graph has a path $(\bar{s}^{1},...,\bar{s}^{T})$ with $\bar{s}^{1}=s$ and $\bar{s}^{T}=s^{\prime\prime}$ such that, for each $t=2,...,T$, $\widetilde{\nu}(\bar{s}^{t})=\widetilde{\nu}(\bar{s}^{t-1})\cdot o(\bar{s}^{t},\bar{s}^{t-1})$, where $o(\bar{s}^{t},\bar{s}^{t-1})\in\mathbb{R}^{++}$ is a DOR between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.
+Then, the graph has a path $(\bar{s}^{1},...,\bar{s}^{T})$ with $\bar{s}^{1}=s$ and $\bar{s}^{T}=s^{\prime}$ such that, for each $t\in\left\{2,...,T\right\}$, $\widetilde{\nu}(\bar{s}^{t})=\widetilde{\nu}(\bar{s}^{t-1})\cdot o(\bar{s}^{t},\bar{s}^{t-1})$, where $o(\bar{s}^{t},\bar{s}^{t-1})\in\mathbb{R}^{++}$ is a DOR between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.
+That is, there is a path in the graph such that for every two consecutive states, the value of $\tilde{\nu}$ for one state has been defined by the other.
 (Note that, if $\widetilde{\nu}(\bar{s}^{t-1})$ was defined as $\widetilde{\nu}(\bar{s}^{t})\cdot o(\bar{s}^{t-1},\bar{s}^{t})$ and not vice versa, there is also a DOR $o(\bar{s}^{t},\bar{s}^{t-1})=1/o(\bar{s}^{t-1},\bar{s}^{t})$ between $\bar{s}^{t}$ and $\bar{s}^{t-1}$.) Therefore
@@
 To show that $\mu(\cdot|h)$ and $\hat{\nu}^{m}$ coincide, it is enough to show that $\mathrm{supp}\mu(\cdot|h)=\hat{S}$ and $\mu(s|h)/\mu(s^{\prime}|h)=\hat{\nu}^{m}(s)/\hat{\nu}^{m}(s^{\prime})$ for all $s,s^{\prime}\in\hat{S}$.
-First we show that $\mathrm{supp}\mu(\cdot|h)=P^{m}\cap S(h)$.
+First, we show that $\mathrm{supp}\mu(\cdot|h)=P^{m}\cap S(h)$.
 Fix $s\in\mathrm{supp}\mu(\cdot|h)$.
@@
 As $\mathrm{supp}\nu^{m}=P^{m}$ and $\nu^{m}(S(h))>0$, there exists $\bar{s}\in S(h)\cap P^{m}$ such that $\nu^{m}(\bar{s})>0$.
-Since $s\in\mathrm{supp}\mu(\cdot|h)$, $s$ has a non-zero DOR with every $s^{\prime}\in S(h)$, so in particular with $\bar{s}$.
-Since $s\not\in P^{1}\cup...\cup P^{m}$, $s$ has a zero GOR with some $s^{\prime}\in P^{m}\cup...\cup P^{n}$.
-The two things combined impy that $\bar{s}$ has a zero GOR with $s^{\prime}$ too.
+Since $\in\mathrm{supp}\mu(\cdot|h), ss$ has a non-zero DOR (which can be finite or infinite) with every $s^{\prime}\in S(h)$, so in particular with $\bar{s}$.
+Since $\not\in\cup_{i=1}^{m}P^{i}, ss$ has a zero GOR with some $s^{\prime}\in\cup_{i=m}^{n}P^{i}$.
+The two things combined imply that $\bar{s}$ has a zero GOR with $s^{\prime}$ as well.
 But this contradicts $\bar{s}\in P^{m}$.
@@
 If $s\not\in S(h)$, we are done.
-If $s\in S(h)$, $s\not\in P^{m}$, because $s$ has a zero DOR with every $s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and we have just shown that $\mathrm{supp}\mu(\cdot|h)\subseteq P^{m}$.
+If $s\in S(h)$, $s\not\in P^{m}$, because $s$ has a zero DOR with every $s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$, and we have just shown that $\mathrm{supp}\mu(\cdot|h)\subseteq P^{m}$.
 To conclude the proof, fix a pair $s,s^{\prime}\in\mathrm{supp}\mu(\cdot|h)$ and write
@@
 where $o(s,s^{\prime})$ is the DOR between $s$ and $s^{\prime}$ at $h$, and the last equality follows from property (b) of $\nu^{m}$.
-$\mathbf{2\Rightarrow 3)}$
-Let $\bar{\mu}=(\mu^{1},...,\mu^{n})$ be a LCPS from which the belief system is derived.
-For each $m=1,...,n$, let $S^{m}$ denote the support of $\mu^{m}$, and let $H^{m}$denote the set of all $h\in\mathcal{H}$ such that $\mu(\cdot|h)$ is derived from $\mu^{m}$.
+$\mathbf{2\Rightarrow 3)}$ Let $\bar{\mu}=(\mu^{1},...,\mu^{n})$ be an LCPS from which the belief system is derived.
+For each $m\in\left\{1,...,n\right\}$, let $S^{m}$ denote the support of $\mu^{m}$, and let $H^{m}\$denote the set of all $h\in\mathcal{H}$ such that $\mu(\cdot|h)$ is derived from $\mu^{m}$.
 Fix a system of gambles $g=(g(\cdot|h))_{h\in\mathcal{H}}$.
@@
 We show that $g$ is not a Dutch book.
-Suppose that $g$satisfies condition (1) of a Dutch book, so that,
+Suppose that gg\ satisfies condition (1) of a Dutch book, so that,
 $\forall s\in\mathcal{S},\text{ \ \ }\sum_{h\in\mathcal{H}}p(h|s)g(s|h)\leq 0.$ (5)
@@
 The proof is by induction on $m$.
-Fix $m=1,...,n$ and suppose by way of induction that for every $k<m$, $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in H^{k}$.
+Fix $m\in\left\{1,...,n\right\}$ and suppose by induction that for every $k<m$, $g(s|h)=0$ for all $s\in\mathcal{S}$ and $h\in H^{k}$.
 (This is vacuously true for $m=1$.) Fix $s\in S^{m}$.
@@
 $\sum_{s\in S}\mu^{m}(s)\sum_{h\in H^{m}}p(h|s)g(s|h)\leq 0\text{.}$ (6)
-Since $\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})>0$ for every $h\in H^{m}$, we can rewrite (6) as
-$\sum_{h\in H^{m}}\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\leq 0.$ (7)
+Since $\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})>0$ for every $h\in H^{m}$, we can rewrite (6) as
+$\sum_{h\in H^{m}}\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\leq 0.$ (7)
 For each $h\in H^{m}$, since the agent is willing to accept $g(\cdot|h)$, we have
 $\sum_{s\in\mathcal{S}}\mu(s|h)g(s|h)\geq 0.$ (8)
-Since $\mu(\cdot|h)$ can be derived from $\mu^{m}$ with Bayes rule, we can rewrite (8) as
-$\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\geq 0.$ (9)
+Since $\mu(\cdot|h)$ can be derived from $\mu^{m}$ by Bayes rule, we can rewrite (8) as
+$\forall h\in H^{m},\text{ \ \ }\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)\geq 0.$ (9)
 Equations (9) and (7) imply
-$\forall h\in H^{m},\text{ \ \ }\sum_{s\in S}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)=0,$
+$\forall h\in H^{m},\text{ \ \ }\sum_{s\in\mathcal{S}}\frac{p(h|s)\mu^{m}(s)}{\sum_{s^{\prime}\in\mathcal{S}}p(h|s^{\prime})\mu^{m}(s^{\prime})}g(s|h)=0,$
 that is,
-$\sum_{s\in S}\mu(s|h)g(s|h)=0.$
+$\sum_{s\in\mathcal{S}}\mu(s|h)g(s|h)=0.$
 This means that the agent has no strict incentive to accept $g(\cdot|h)$, therefore acceptance implies $g(s|h)=0$ for all $s\in\mathcal{S}$.
-$\mathbf{3\Rightarrow 1)}$
-The proof that an agent with two different GORs between a state $s$ and a state $s^{\prime}$ can be Dutch-booked resembles the classical Dutch-book arguments of De Finetti and Ramsey.
-Therefore, the proof is deferred to the Appendix.
-$\blacksquare$
-The next example illustrates the theorem.
-Example 5
+$\mathbf{3\Rightarrow 1)}$ It will follow as an implication of the more general Proposition 1. $\blacksquare$
+6.1 Robustness to the Bookmaker Misspecification
+One natural concern about the relevance of the previous result is that a misspecified agent has been shown to be exploitable by a bookmaker who perfectly knows the true data-generating process.
+In principle, if the bookmaker model is also incorrect, then the proposed bet may induce some losses for the bookmaker in some $s\in S$.
+We next show that this is not the case as long as the misspecification is sufficiently small, and we quantify what small means.
+To do so, we start by observing that beliefs that are not completely coherent admit a special form of inconsistency in the generalized discounted odd ratios.
+Lemma 1
+If $\mu$ is not completely coherent, then there exists a state $s$ and a GOR between $s$ and itself with value $r<1$ and where $s^{0},...,s^{n-1}$ are all distinct for each $k\in\left\{i,...,n-1\right\}$, $h^{k}\not=h^{k+1}$.
+The lemma shows that without complete coherence, the agent’s belief can be used to build a GOR that starts and ends at the same state and differs from $1$, the only value it could instead assume under complete coherence.
+This observation motivates the following definition.
+Definition 10
+Let $r<1$ and $n\geq 2$.
+We say that $\mu$ is $(r,n)$-not-coherent if there exists a state $s$ and a GOR between $s$ and itself with value $r<1$ and where $s^{0},...,s^{n-1}$ are all distinct for each $k\in\left\{i,...,n-1\right\}$, $h^{k}\not=h^{k+1}$.
+This definition quantifies the degree of incoherence between the agent’s beliefs $\mu$ and the data generating process $p$, with smaller values of $r$ and $n$ indicating a more significant incoherence.
+Indeed, a smaller $r$ indicates that a larger departure from $1$ of the GOR between $s$ and itself, and a smaller $n$ indicates that such a departure can be obtained by just using a smaller number of comparisons between state-contingency pairs.
+Next, we define what it means for a Dutch book to be robust to misspecification on the bookmaker side.
+Definition 11
+Let $\beta\in\mathbb{R}_{+}$. A Dutch book $g=(g(\cdot|h))_{h\in\mathcal{H}}$ is a $\beta$-robust Dutch book if for every $\hat{p}$ such that
+$\frac{\hat{p}(h|s)}{\hat{p}(h^{\prime}|s)}\frac{p(h^{\prime}|s)}{p\left(h|s\right)}<\beta\qquad\forall h,h^{\prime}\in\mathcal{H},s\in S$
+we have that for all $s\in\mathcal{S}$,
+$\sum_{h\in\mathcal{H}}\hat{p}(h|s)g(s|h)\leq 0,$
+and for some $s\in\mathcal{S}$,
+$\sum_{h\in\mathcal{H}}\hat{p}(h|s)g(s|h)<0.$
+This refinement of the Dutch book concept should be interpreted as follows.
+As before, the agent is endowed with beliefs $\mu$, and the bookmaker considers the data generating process $p$.
+However, she no longer completely trusts it and envisions the possibility that their understanding of how a state translates into different consequences (i.e., $\frac{p(h^{\prime}|s)}{p\left(h|s\right)}$) may be incorrect to some extent.
+Therefore, they require the Dutch book to lead to a payoff that is nonnegative under every state and positive under some state, even if the true data generating process is some different $\hat{p}$ that is sufficiently close to $p$.
+The next result shows that such robustness can be achieved, and it links it to the quantitative notion of incoherence defined above.
+Proposition 1
+Suppose that $\mu$ is $(r,n)$-not-coherent. Then for every $\delta>0$ there exists a $\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$-robust Dutch book.
+Importantly, the result shows that to achieve a Dutch book, what is key is the comparison between the disagreement between the agent and the bookmaker, and the misspecification of the bookmaker.
+If the disagreement between the two, captured by the measure $\left(r,n\right)$ of how far are beliefs $\mu$ from being generated by the bookmaker model $p$ and some prior, is sufficiently high compared to the extent of misspecification of $p$ with respect to $\hat{p}$.
+7 Applications to Correlation Neglect
+We next apply our main theorem to one of the most pervasive and studied forms of misspecification, correlation neglect.
+We start by studying how the use of different financial instruments in venture capital can be viewed as a Dutch book due to neglect of correlation on the startup side.
+We then move to see how some offers in a variant of the classical Monty Hall problem can also be interpreted in the same way.
+Financial Instruments
+Consider a binary-state environment, where states $\mathcal{S}=\left\{G,B\right\}$ indicate whether a startup is good or bad.
+Evidence about the startup’s quality accumulates over time through a sequence of binary signals indicating whether key milestones are achieved.
+That is, the set of contingencies is $\mathcal{H=}\left\{\left\{\gamma,\beta\right\}^{n}\right\}_{n\in\left\{1,...,N\right\}}$, $N\geq 4$, where $\gamma$ was a good performance and $\beta$ was a bad performance in a key milestone.
+The objective data generating process over contingencies is Markov, with two key properties.
+First, good firms are more likely to achieve the milestones, and second, there is persistence in the milestone process.
+Regardless of whether the startup is good or bad, the probability of achieving the second milestone is higher if the first one has been achieved.
+For concreteness, the objective DGP is recursively given by
+$p\left(\gamma|G\right)$ $=$ $\frac{2}{3}=p\left(\beta|B\right)$
+$p\left(\left(h,\gamma,\gamma\right)|\left(h,\gamma\right),G\right)$ $=$ $\frac{3}{4}=p\left(\left(h,\beta,\beta\right)|\left(h,\beta\right),B\right)\quad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-2$
+$p\left(\left(h,\beta,\gamma\right)|\left(h,\beta\right),G\right)$ $=$ $\frac{1}{2}=p\left(\left(h,\gamma,\beta\right)|\left(h,\gamma\right),B\right)\quad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-2\text{.}$
+The startup updates its beliefs using Bayes’ rule and understands that milestones are more likely to be reached when the state is good than when it is bad.
+However, the startup suffers from correlation neglect and treats milestone achievements as i.i.d.
+conditional on the state.
+Concretely, the belief system of the startup is given by
+$\mu\left(G|\varnothing\right)=\frac{1}{2}\text{,}$
+and for all $h\in\left\{\gamma,\beta\right\}^{n},n\leq N$
+$\mu\left(G|h\right)=\frac{\left(\frac{2}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\gamma\right\}|}\left(\frac{1}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\beta\right\}|}}{\left(\frac{2}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\gamma\right\}|}\left(\frac{1}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\beta\right\}|}+\left(\frac{2}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\beta\right\}|}\left(\frac{1}{3}\right)^{|\left\{i\in\left\{1,...,n\right\}:h_{i}=\lambda\right\}|}}\text{.}$ (10)
+By Theorem 1, the startup cannot be deterministically Dutch booked, as the beliefs are derived using Bayesian updating from the conditional probabilities
+$\widetilde{p}\left(\left(h,\gamma\right)|h,G\right)$ $=$ $\frac{2}{3}\qquad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-1$
+$\widetilde{p}\left(\left(h,\gamma\right)|h,B\right)$ $=$ $\frac{1}{3}\qquad\forall h\in\left\{\gamma,\beta\right\}^{n},n\leq N-1\text{.}$
+However, a venture capital offering a combination of “purchase the firm” and “convertible debt” can Dutch-book the startup by exploiting their correlation neglect.
+In particular, and without loss of generality, given the risk neutrality of the agent, suppose that ownership of a good firm is worth $\frac{2}{3}$ in the good state and $\frac{1}{3}$ in the bad state (that is, each state is worth its limit distribution over positive performances).
+The startup initially holds ownership.
+The venture capitalist offers two different contracts in contingencies of length $4$.
+After the sequence $\left(\gamma,\beta,\beta,\gamma\right)$, the venture capitalist offer to buy the startup at a price of $0.51$.
+This corresponds to the bet
+$g\left(G|\left(\gamma,\beta,\beta,\gamma\right)\right)=0.51-\frac{2}{3}\text{ and }g\left(B|\left(\gamma,\beta,\beta,\gamma\right)\right)=0.51-\frac{1}{3}\text{.}$
+By equation (10), $\mu\left(G|\left(\gamma,\beta,\beta,\gamma\right)\right)=\frac{1}{2}$, and thus the startup accepts and sells ownership, overweighting the informativeness of the two consecutive failures.
+Instead, after sequence $\left(\gamma,\gamma,\gamma,\gamma\right)$, convertible debt at a rate below the market rate is offered to the up-and-coming startup.
+This below-market-rate debt amounts to a state-independent financing advantage of $\frac{1}{100}$ to the startup, but in the unlikely case the startup turns out to be of the bad type, the venture capital gains control of half of the ownership of the firm.
+This corresponds to the bet [^16] This contract can also be approximately implemented as a transfer of property if the fraction of milestones over a long horizon is smaller than $1/2$, as this is an event that has probability approaching $1$ under $B$ and 0 under $G$ as the horizon grows.
+$g\left(G|\left(\gamma,\gamma,\gamma,\gamma\right)\right)=\frac{1}{100}\text{ and }g\left(B|\left(\gamma,\gamma,\gamma,\gamma\right)\right)=\frac{1}{100}-\frac{1}{2}\frac{1}{3}\text{.}$
+By equation (10), $\mu\left(G|\left(\gamma,\gamma,\gamma,\gamma\right)\right)=\frac{16}{17}$ and thus the startup accepts the convertible debt, overweighting the informativeness of these successes.
+Easy computations verify that this is indeed a Dutch book.
+[^17] In state $G$, contingency $\left(\gamma,\beta,\beta,\gamma\right)$ is reached with probability $0.0416$ with a startup payoff of 0.1566-0.1566, and $\left(\gamma,\gamma,\gamma,\gamma\right)$ is reached with probability $0.28125$ and a startup payoff of $0.01$.
+In state $B$, contingency $\left(\gamma,\beta,\beta,\gamma\right)$ is reached with probability $0.03125$ with a startup payoff of $0.1766$, and $\left(\gamma,\gamma,\gamma,\gamma\right)$ is reached with probability $0.0416$ and a startup payoff of $0.01-\frac{1}{6}$, resulting in strictly negative expected losses in both states.
+No matter the baseline frequency of successful startups, the venture capitalist can, on average, extract positive returns from both types, exploiting the acquisition of good startups after mixed performance and convertible debt issued to bad startups after an initial sequence of successes.
+We also observe that, consistently with Theorem 2, the generalized odds ratios of the startup do not agree
+$o\left(G,B|\left(\gamma,\gamma,\gamma,\gamma\right)\right)$ $=$ $\frac{\mu\left(g|\left(\gamma,\gamma,\gamma,\gamma\right)\right)}{p\left(\left(\gamma,\gamma,\gamma,\gamma\right)|G\right)}\frac{p\left(\left(\gamma,\gamma,\gamma,\gamma\right)|B\right)}{\mu\left(b|\left(\gamma,\gamma,\gamma,\gamma\right)\right)}=\frac{16/17}{0.28125}\frac{0.0416}{1/17}=2.366$
+$>$ $\frac{1/2}{0.0416}\frac{0.03125}{1/2}=\frac{\mu\left(G|\left(\gamma,\beta,\beta,\gamma\right)\right)}{p\left(\left(\gamma,\beta,\beta,\gamma\right)|G\right)}\frac{p\left(\left(\gamma,\beta,\beta,\gamma\right)|B\right)}{\mu\left(B|\left(\gamma,\beta,\beta,\gamma\right)\right)}=o\left(G,B|\left(\gamma,\beta,\beta,\gamma\right)\right)\text{.}$
+In words, the agent interprets the good type sequence $\left(\gamma,\gamma,\gamma,\gamma\right)$ too positively, as it fails to account for the persistence of performance.
+This application, among other things, illustrates the difference between our Dutch book argument and preexisting arguments against misspecification.
+Typically, it is argued that some misspecifications cannot persist because, in the long run, key statistics of the data force the agent to either reject or incompletely trust their model.
+However, the form of misspecification above would be immune to these arguments.
+Indeed, as the horizon grows, i.e., $N\rightarrow\infty$, the startup beliefs concentrate on the actual state, i.e., $\mu\left(G|h\right)$ converges to $1$ almost surely conditional on being in state $G$, and the same holds for $\mu\left(B|h\right)$ under $B$.
+Moreover, the startup will observe a fraction of success equal to the one predicted by their incorrect model (i.e., the stationary distribution of $\gamma$ under $g$ is $2/3$, which is the probability of $\gamma$ under their i.i.d.
+model).
+Under essentially all the work that studied the persistence of misspecification, these two conditions are sufficient to guarantee that their misspecified model will not be relaxed or changed (Gagnon-Bartsch, Rabin, and Schwartzstein 2023, Fudenberg and Lanzani 2023, Ba 2025, He and Libgober 2025, Lanzani 2025).
+Our argument against misspecification, instead, does not work at the level of beliefs and how they evolve over time, but at the level of material losses deriving from the cross-sectional exploitation of a population of misspecified agents.
+Monty Hall offers
 This is a variation of the famous Monty Hall problem, in which the initially selected door is determined at random and not revealed to the decision maker.
@@
 The state represents the door with the car: $\mathcal{S}=\left\{A,B,C\right\}$.
-Then, one door, call it $d$, is drawn from a uniform distribution (in the Monty Hall problem, this is the door assigned to the agent).
-Finally, another door behind which there is a goat, call it $h\not=d$, is opened; it is the first non-car door preceding $d$ in an $a$-$b$-$c$ circle.
-So, for instance, conditional on $s=B$, door $a$ is opened with probability $2/3$ (the probability that $d\in\left\{b,c\right\}$), and door $c$ is opened with probability $1/3$ (the probability that $d=a$).
-The agent only observes which door is opened, so the possible contingencies are: =$\mathcal{H}=\left\{a,b,c\right\}$.
+Then, one door, call it $d$, is drawn from a uniform distribution (under the Monty Hall interpretation, this is the door initially assigned to the agent).
+Finally, another door behind which there is a goat, call it $h\not=d$, is opened; it is the first non-car door preceding $d$ in an $a$-$b$-$c$ cycle.
+So, for instance, conditional on the state being $B$, door $a$ is opened with probability $2/3$ (the probability that $d\in\left\{b,c\right\}$), and door $c$ is opened with probability $1/3$ (the probability that $d=a$).
+The agent only observes which door is opened, so the possible contingencies are: $\mathcal{H}=\left\{a,b,c\right\}$.
 Overall, conditional on each $s\in\mathcal{S}$, the probabilities of each $h\in\mathcal{H}$ are:
 $p(h|s)$ $h=a$ $h=b$ $h=c$
-$s=A$ 00 $\frac{1}{3}$ $\frac{2}{3}$
-$s=B$ $\frac{2}{3}$ 00 $\frac{1}{3}$
-$s=C$ $\frac{1}{3}$ $\frac{2}{3}$ 00
+$s=A$ 0 $\frac{1}{3}$ $\frac{2}{3}$
+$s=B$ $\frac{2}{3}$ 0 $\frac{1}{3}$
+$s=C$ $\frac{1}{3}$ $\frac{2}{3}$ 0
 We next consider two possible belief systems for the agent.
 First, suppose that, at every $h\in\mathcal{H}$, the agent assigns equal probability to each of the two states in $S(h)$.
+That is, they understand that the car cannot be behind the open door, but they neglect the correlation between the open door and the one containing the car, induced by the open-in-cycle mechanism.
 Then, the agent can be Dutch-booked (in objective expected terms).
 To see this, compute the three discounted odds ratios:
-$o(B,C|a)$ $=$ $\frac{\Pr(a|C)}{\Pr(a|B)}=\frac{1/3}{2/3}=\frac{1}{2};$
-$o(C,A|b)$ $=$ $\frac{\Pr(b|A)}{\Pr(b|C)}=\frac{1/3}{2/3}=\frac{1}{2};$
-$o(B,A|c)$ $=$ $\frac{\Pr(c|A)}{\Pr(c|B)}=\frac{2/3}{1/3}=2.$
+$o(B,C|a)$ $=$ $\frac{p(a|C)}{p(a|B)}=\frac{1/3}{2/3}=\frac{1}{2};$
+$o(C,A|b)$ $=$ $\frac{p(b|A)}{p(b|C)}=\frac{1/3}{2/3}=\frac{1}{2};$
+$o(B,A|c)$ $=$ $\frac{p(c|A)}{p(c|B)}=\frac{2/3}{1/3}=2.$
 Since
@@
 For example, suppose that, at each $h\in\mathcal{H}$, a bookmaker offers to the agent the following bets:
-$g(C|a)$ $=$ 3, $3,\text{ \ }g(B|a)=-2;$
-$g(A|b)$ $=$ 3, $3,\text{ \ }g(C|b)=-2;$
-$g(B|c)$ $=$ 3, $2\displaystyle 3,\text{ \ }g(A|c)=-$ 2 .
+$g(C|a)$ $=$ $3,\text{ \ }g(B|a)=-2;$
+$g(A|b)$ $=$ $3,\text{ \ }g(C|b)=-2;$
+$g(B|c)$ $=$ $3,\text{ \ }g(A|c)=-2.$
 Clearly, the agent accepts all bets, and the objective expected payoff of the bookmaker conditional on each $s\in\mathcal{S}$ is:
 $-g(A|b)p(b|A)-g(A|c)p(c|A)$ $=$ $-g(B|c)p(c|B)-g(B|a)p(a|B)=-g(C|a)p(a|C)-g(C|b)p(b|C)$
-[…]
+$=$ $-3\frac{1}{3}+2\frac{2}{3}>0\text{.}$
 So, the agent is harmed by the “correlation neglect” he experiences.
@@
 He has a uniform prior, and, thinking of the correct updating in the original Monty Hall problem, believes that the car is more likely to be behind the unopened door that has not been assigned to him (i.e., not door $d$).
-For instance, suppose that $h=a$ and the agent correctly computes
-$\Pr\left(d=b|h=a\right)$
-$=$ $\frac{\Pr(h=a|d=b)\Pr(d=b)}{\Pr(h=a|d=b)\Pr(d=b)+\Pr(h=a|d=c)\Pr(d=c)}$
-$=$ $\frac{\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)}{\left(\begin{array}[]{c}\Pr(h=a|d=b,s=B)+\Pr(h=a|d=b,s=C)\\ +\Pr(h=a|d=c,s=B)+\Pr(h=a|d=c,s=C)\end{array}\right)}$
-$=$ $\frac{2}{2+1}=\frac{2}{3}.$
-So, he has $\mu(s=B|h=a)<1/2$. However, given $h=a$, the car is more likely to be behind door $b$:
-$\Pr\left(s=B|h=a\right)$
-$=$ $\frac{\Pr(h=a|s=B)\Pr(s=B)}{\Pr(h=a|s=B)\Pr(s=B)+\Pr(h=a|s=C)\Pr(s=C)}$
-$=$ $\frac{\Pr(s=B)}{\Pr(s=B)+\Pr(d=b)\Pr(s=C)}$
-$=$ $\frac{1}{1+\frac{1}{3}}=\frac{3}{4}.$
+For instance, suppose that we are in contingency $a$ and the agent correctly computes
+$\Pr\left(d=b|h=a\right)$ $=$ $\frac{\Pr(h=a|d=b)\Pr(d=b)}{\Pr(h=a|d=b)\Pr(d=b)+\Pr(h=a|d=c)\Pr(d=c)}$
+$=$ $\frac{\Pr(h=a|\left(d,s\right)=\left(b,B\right))+\Pr(h=a|\left(d,s\right)=\left(b,C\right))}{\left(\begin{array}[]{c}\Pr(h=a|\left(d,s\right)=\left(b,B\right))+\Pr(h=a|\left(d,s\right)=\left(b,C\right))\\ +\Pr(h=a|\left(d,s\right)=\left(c,B\right))+\Pr(h=a|\left(d,s\right)=\left(c,C\right))\end{array}\right)}=\frac{2}{2+1}=\frac{2}{3}.$
+So, he has $\mu(s=B|h=a)<1/2$, as prescribed by correct reasoning in the (original) Monty Hall problem: the car is less likely to be behind the door that is more likely to have been assigned to him.
+However, given contingency $a$, if the distribution over states is uniform, the car is more likely to be behind door $b$:
+$\Pr\left(s=B|h=a\right)$ $=$ $\frac{p\left(a|B\right)\Pr(s=B)}{p\left(a|B\right)\Pr(s=B)+p\left(a|C\right)\Pr(s=C)}$
+$=$ $\frac{\frac{2}{3}\Pr(s=B)}{\frac{2}{3}\Pr(s=B)+\Pr(d=b)\Pr(s=C)}=\frac{\frac{2}{3}}{\frac{2}{3}+\frac{1}{3}}=\frac{2}{3}.$
 In other words, the agent perceives the opposite correlation between the fundamental state and the observed contingency.
 This increases the discrepancy between generalized odds ratios, making it easier for the bookmaker to Dutch-book the agent.
-[^13] To be precise, without limited liability any discrepancy allows the bookmaker to extract an infinite amount of money (in expectation) from the agent.
-But under limited liability, the higher discrepancy allows to strictly increase the bookmaker’s expected profit.
+[^18] To be precise, without limited liability, any discrepancy allows the bookmaker to extract an infinite amount of money (in expectation) from the agent.
+However, under limited liability, the higher discrepancy allows for a strict increase in the bookmaker’s expected profit.
 8 Conclusion
-Dutch-book arguments have already been used to motivate probabilistic sophistication and conditioning.
+Dutch-book arguments have been used to motivate probabilistic sophistication and conditioning.
 We introduce novel Dutch-book arguments that motivate two forms of belief consistency in a learning framework a la Anscombe-Aumann, with separation between subjective uncertainty (the state) and objective uncertainty (lotteries over learning paths conditional on the state).
@@
 A stronger form of coherence, guaranteed by deriving all beliefs from a subjective lexicographic prior with the correct models, is necessary and sufficient to avoid being Dutch-booked “in objective expected terms”, i.e., under each realized state, given the corresponding lottery on the learning paths.
-The latter form of Dutch-booking entails sure aggregate losses across a population of agents, thus showing the perils of heterogenous priors across different individuals, and of model misspecification even in absence of an objective prior.
+The latter form of Dutch-booking entails sure aggregate losses across a population of agents, thereby showing the perils of heterogeneous priors across individuals and of model misspecification even in the absence of an objective prior.
 Appendix A Appendix
@@
 Therefore, the assumption that every contingency is reached with the same probability under every consistent state is satisfied in that context.
-For concreteness, we stick to the sequential games application, but the analysis extends seamlessly to situations where this assumption is satisfied, for instance in situations where the an agent does not choose what to observe but the factors that determine the observations are independent of the state itself, as in the running example of the paper.
-Let $S_{-i}$ be the set of stategy profiles of the opponents of player $i$.
+[^19] For concreteness, we stick to the sequential games application, but the analysis extends seamlessly to situations where this assumption is satisfied, for instance in situations where the an agent does not choose what to observe but the factors that determine the observations are independent of the state itself, as in the running example of the paper.
+Let $S_{-i}$ be the set of strategy profiles of the opponents of player $i$.
 Let $H_{i}$ be the arborescence of information sets of player $i$, endowed with the precedence relation $\prec$.
-Each information set $h_{i}$ corresponds to the observation of some past moves of the opponents, summarized by the event $S_{-i}(h_{i})\subseteq S_{-i}$.
+Each information set $h_{i}$ is associated with the event $S_{-i}(h_{i})\subseteq S_{-i}$ of the opponents strategies that reach $h_{i}$, under some strategy of player $i$.
 A belief system is an array $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ of probability measures over $S_{-i}$ such that $\mu_{i}(S_{-i}(h_{i})|h_{i})=1$ for every $h_{i}\in H_{i}$.
-Taking as primitive a collection of conditioning events $\mathcal{C\subseteq}2^{S_{-i}}\backslash\left\{\emptyset\right\}$, a conditional probability system over $S_{-i}$ is an array of probability measures $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in\mathcal{C}}$ such that $\bar{\mu}_{i}(C|C)=1$ for every $C\in\mathcal{C}$, and for every $D\in\mathcal{C}$ with $D\subset C$, for each $E\subseteq D$,
+[…]
 $\bar{\mu}_{i}(E|C)=\bar{\mu}_{i}(E|D)\bar{\mu}_{i}(D|C).$
-A conditional probability system is complete when it is defined with respect to the collection $\mathcal{C=}2^{S_{-i}}\backslash\left\{\emptyset\right\}$.
+[…]
 Definition 12 (Battigalli et al., 2023)
 A belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ over the opponents’ strategies is forward consistent if, for all $h_{i},h_{i}^{\prime}\in H_{i}$, if $h_{i}\prec h_{i}^{\prime}$, then
-$\mu_{i}(E|h_{i})=\mu_{i}(S_{-i}(h_{i}^{\prime})|h_{i})\mu(E|h_{i}^{\prime})$
+$\mu_{i}(E|h_{i})=\mu_{i}(S_{-i}(h_{i}^{\prime})|h_{i})\mu_{i}(E|h_{i}^{\prime})$
 for every $E\subseteq S_{-i}(h_{i}^{\prime})$.
@@
 Definition 13
-[…]
-$\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i})).$
+A belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ over the opponents strategies is completely consistent if there exists a CCPS $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}}$ such that
+$\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}))\qquad\forall h_{i}\in H_{i}.$
 The definition of complete consistency of Battigalli et al.
-(2023) uses CCPS’s to discipline beliefs at counterfactual contingencies.
+(2023) uses CCPS’s to discipline beliefs in counterfactual contingencies.
 An information set $h$ does not carry any additional information about the strategies of the opponents besides ruling out all the strategies that do not belong to $S_{-i}(h)$; for this reason, it is appropriate to derive each belief by mere conditioning in place of Bayes rule, from the most general theory that explains the contingency.
 Proposition 2
-For every information set $h_{i}$, for every pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h_{i})$, and for every probability distribution over player $i$’s strategies, the probability of reaching the information set is identical under $s_{-i}$ and $s_{-i}^{\prime}$: $p(h_{i}|s_{-i})=p(h_{i}|s_{-i}^{\prime})$.
+For every $h_{i}\in H_{i}$, for every pair $s_{-i},s_{-i}^{\prime}\in S_{-i}(h_{i})$, and for every probability distribution over player $i$’s strategies, the probability of reaching the information set is identical under $s_{-i}$ and $s_{-i}^{\prime}$: $p(h_{i}|s_{-i})=p(h_{i}|s_{-i}^{\prime})$.
 Proof.
@@
 $\blacksquare$
-Although the use of CCPSs is appropriate in this context, one might as well model the theories from which a player derives just as a LCPS, and derive each belief by conditioning the first theory in the list that is consistent with the contingency; the two definitions of complete consistency are equivalent.
+Although the use of CCPSs is appropriate in this context, one might as well model the theories from which a player derives beliefs as a LCPS, and derive each belief by conditioning the first theory in the list that is consistent with the contingency; the two definitions of complete consistency are equivalent.
 [^20] The equivalence between LCPSs and CCPSs is well-known.
@@
   1. 1.
-[…]
-    $\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}));$
+    there exists a CCPS $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}}$ such that
+    $\mu_{i}(\cdot|h_{i})=\bar{\mu}_{i}(\cdot|S_{-i}(h_{i}))\qquad\forall h_{i}\in H_{i};$
   2. 2.
-    there exists a LCPS $\hat{\mu}_{i}=(\mu_{i}^{1},...,\mu_{i}^{n})$ such that for each $h_{i}\in H_{i}$ and $s_{-i}\in S_{-i}(h_{i})$,
-    $\mu_{i}(s_{-i}|h_{i})=\frac{\mu_{i}^{k}(s_{-i})}{\mu_{i}^{k}(S_{-i}(h_{i}))},$
+    there exists a LCPS $\hat{\mu}_{i}=(\mu_{i}^{1},...,\mu_{i}^{n})$ such that
+    $\mu_{i}(s_{-i}|h_{i})=\frac{\mu_{i}^{k}(s_{-i})}{\mu_{i}^{k}(S_{-i}(h_{i}))}\qquad\forall h_{i}\in H_{i},\forall s_{-i}\in S_{-i}(h_{i}),$
     where $k$ is the smallest $m$ such that $\mu_{i}^{k}(S_{-i}(h_{i}))>0$.
 Proof.
+$\mathbf{2\Rightarrow 1)}$ For each $C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}$, derive $\bar{\mu}_{i}(\cdot|C)$ by conditioning the first measure in the LCPS that assigns positive probability to $C$, so that if $C=S_{-i}(h_{i})$ for some $h_{i}\in H_{i}$, then $\bar{\mu}_{i}(\cdot|C)=\mu_{i}(\cdot|h_{i})$.
+There remains to show that $\bar{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C))_{C\in 2^{S_{-i}}\backslash\left\{\emptyset\right\}}$ is a CCPS.
+Now fix $D\subset C$ such that $\bar{\mu}_{i}(D|C)>0$.
+Thus, if $\bar{\mu}_{i}(\cdot|C)$ is derived from $\mu_{i}^{k}$, $\mu_{i}^{k}(D)>0$, and $\mu_{i}^{j}(D)\leq\mu_{i}^{j}(C)=0$ for every $j<k$.
+Therefore, also $\bar{\mu}_{i}(\cdot|D)$ is derived from $\mu_{i}^{k}$ and the chain rule is satisfied.
 […]
-1$\Rightarrow$2) Create a decreasing chain of conditioning events $(C^{1},...,C^{n})$ as follows.
-Let $C^{1}=S_{-i}$.
-For each $k>1$, let $C^{k}$ be the union of all the $s_{-i}\in S_{-i}$ such that $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$.
-Let $n$ be the smallest $k$ such that $C^{k+1}=\emptyset$.
-Consider the list of measures $\hat{\mu}_{i}=(\bar{\mu}_{i}(\cdot|C^{1}),...,\bar{\mu}_{i}(\cdot|C^{n}))$.
-For each $h_{i}\in H_{i}$, let $k$ be the smallest $j\leq n$ such that $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})>0$ — it exists by $C^{n+1}=\emptyset$.
-By $\bar{\mu}_{i}(S_{-i}(h_{i})|C^{j})=0$ for every $j<k$, it follows that $S_{-i}(h_{i})\subseteq C^{j}$.
-Hence, $\mu_{i}(\cdot|h_{i})$ can be derived from $\bar{\mu}_{i}(\cdot|C^{k})$ by updating, and $\bar{\mu}_{i}(\cdot|C^{k})$ is also the first measure in $\hat{\mu}_{i}$ that gives positive probability to $S_{-i}(h_{i})$.
-There remains to show that $\hat{\mu}_{i}$ is a LCPS.
-For each $s_{-i}\in S_{-i}$, by $C^{n+1}=\emptyset$, $\bar{\mu}_{i}(s_{-i}|C^{k})>0$ for some $k\leq n$, so $\hat{\mu}_{i}$ has full joint support.
-Moreover, by $s_{-i}\in C^{k}$, $\bar{\mu}_{i}(s_{-i}|C^{j})=0$ for every $j<k$, and by $\bar{\mu}_{i}(s_{-i}|C^{k})>0$, $s_{-i}\not\in C^{j}$ for every $j>k$, which implies $\bar{\mu}_{i}(s_{-i}|C^{j})=0$.
-Hence, $\hat{\mu}_{i}$ has disjoint supports and is therefore a LCPS.
-$\blacksquare$
-Proposition 2 implies that the main result of this paper also applies to belief consistency as it was defined for sequential games.
+Proposition 3 implies that the Dutch-book theorem also applies to belief consistency as it was defined for sequential games.
 With this, the characterization of complete consistency with the coherence rule among discounted odds ratios can be rewritten using traditional odds ratios in place of discounted odds ratios, because in this context the two coincide by Proposition 2.
@@
 Fix a belief system $\mu_{i}=(\mu_{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$. The following are equivalent:
+  1. 1.
     the belief system is completely consistent;
+  2. 2.
     for every $s_{-i},s_{-i}^{\prime}\in S_{-i}$, all the simple generalized odds ratios between $s_{-i}$ and $s_{-i}^{\prime}$ are identical.
@@
 Therefore, Siniscalchi (2020) generalizes the chain rule to apply also between overlapping, but non-nested observable events.
-We conclude this section by showing the convergence between Siniscalchi’s approach and mine.
-Call “generalized self-odds ratio” a generalized odds ratio between $s_{-i}$ and itself.
+We conclude this section by showing the convergence between Siniscalchi’s approach and ours.
+Call “simple generalized self-odds ratio” a simple generalized odds ratio between $s_{-i}$ and itself.
 Remark 1
-All the generalized odds ratios of each pair are identical if and only if all generalized self-odds ratios are $1$.
+All the simple generalized odds ratios of each pair are identical if and only if all simple generalized self-odds ratios are $1$.
 Proof.
-Only if.
-Fix a generalized self-odds ratios.
+Only if Fix a simple generalized self-odds ratio.
 Break it into two and invert one of the two halves.
 We have two concatenations of odds ratios that start and end with the same element.
-Thus, the two corresponding generalized odds ratios are identical by assumption.
-Therefore, the generalized self-odds ratio is $1$.
-If.
-Fix two concatenations of odds ratios where either no odds ratio is infinite or there is no zero.
-If both contain a zero, the corresponding generalized odds ratios are both zero.
+Thus, the two corresponding simple generalized odds ratios are identical by assumption.
+Therefore, the simple generalized self-odds ratio is $1$.
+If Fix a pair and two concatenations of odds ratios of this pair where either no odds ratio is infinite or there is no zero.
+If both contain a zero, the corresponding simple generalized odds ratios are both zero.
 If they both contain an infinite, they are both infinite.
-If one contains an infinite and the other contains a zero, the generalized self-odds ratio obtained by inverting one of the two is zero or infinite, which violates the assumption.
-Otherwise, the generalized self-odds ratio (in one direction) is finite.
+If one contains an infinite and the other contains a zero, the simple generalized self-odds ratio obtained by inverting one of the two is zero or infinite, which violates the assumption.
+Otherwise, the simple generalized self-odds ratio (in one direction) is finite.
 By assumption, it is $1$.
-Therefore the two generalized odds ratios are identical.
+Therefore the two simple generalized odds ratios are identical.
 $\blacksquare$
 Definition 15 (Siniscalchi 2020)
-A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if for every $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$,
+A belief system $(\mu^{i}(\cdot|h_{i}))_{h_{i}\in H_{i}}$ is consistent if for every $(h_{i}^{1},...,h_{i}^{n})\in H_{i}^{n}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$,
 $\mu^{i}(E|h_{i}^{1})\prod_{m=1}^{n-1}\mu^{i}(S_{-i}(h_{i}^{m})\cap S_{-i}(h_{i}^{m+1})|h_{i}^{m+1})=\mu^{i}(E|h_{i}^{n})\prod_{m=1}^{n-1}\mu^{i}(S_{-i}(h_{i}^{m})\cap S_{-i}(h_{i}^{m+1})|h_{i}^{m}).$ (12)
@@
 Proof. We will use the characterization of complete consistency of Corollary 1.
-If part.
-Fix $(h_{i}^{1},...,h_{i}^{n})\in H_{i}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$.
-If both sides of equation 11 are zero, consistency holds.
-Otherwise, suppose without loss of generality that the right-hand side is not zero, so that equation 11 can be rewritten as
+If Fix $(h_{i}^{1},...,h_{i}^{n})\in H_{i}^{n}$ and $E\subseteq S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{n})$.
+If both sides of equation (12) are zero, consistency holds.
+Otherwise, suppose without loss of generality that the right-hand side is not zero, so that equation (12) can be rewritten as
 $\frac{\mu^{i}(E|h_{i}^{1})}{\mu^{i}(S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{2})|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(S_{-i}(h_{i}^{n-1})\cap S_{-i}(h_{i}^{n})|h_{i}^{n})}{\mu^{i}(E|h_{i}^{n})}=1.$ (13)
-Note that equation 12 holds if for every $s_{-i}\in E$ such that the ratio $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ is not indeterminate,
+Note that equation (13) holds if for every $s_{-i}\in E$ such that the ratio $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ is not indeterminate,
 $\frac{\mu^{i}(s_{-i}|h_{i}^{1})}{\mu^{i}(S_{-i}(h_{i}^{1})\cap S_{-i}(h_{i}^{2})|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(S_{-i}(h_{i}^{n-1})\cap S_{-i}(h_{i}^{n})|h_{i}^{n})}{\mu^{i}(s_{-i}|h_{i}^{n})}=1,$ (14)
 because this means that all the non-indeterminate ratios $\mu^{i}(s_{-i}|h_{i}^{1})/\mu^{i}(s_{-i}|h_{i}^{n})$ are identical, which implies that the ratio $\mu^{i}(E|h_{i}^{1})/\mu^{i}(E|h_{i}^{n})$ takes the same value too.
-We are going to show that for every $m=2,...,n$, there is $s_{-i}^{m}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that
+We are going to show that for every $m\in\left\{2,...,n\right\}$, there is $s_{-i}^{m}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that
 $\frac{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m-1})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})};$
-then, we can rewrite equation 13 as
+then, we can rewrite equation (14) as
 $\frac{\mu^{i}(s_{-i}|h_{i}^{1})}{\mu^{i}(s_{-i}^{2}|h_{i}^{1})}\cdot....\cdot\frac{\mu^{i}(s_{-i}^{n}|h_{i}^{n})}{\mu^{i}(s_{-i}|h_{i}^{n})}=1,$ (15)
@@
 Remark 1).
-Let $s_{-i}^{m}$ be any $s_{-i}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that $\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})>0$; one exists by the fact that the right-hand side of 11 is not zero.
-By complete consistency, for each $s_{-i}^{\prime}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$, we have
+Let $s_{-i}^{m}$ be any $s_{-i}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$ such that $\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})>0$; one exists by the fact that the right-hand side of (12) is not zero.
+By complete consistency and Corollary 1, for each $s_{-i}^{\prime}\in S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})$, we have
 $\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{\prime}|h_{i}^{m})}=\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}{\mu^{i}(s_{-i}^{\prime}|h_{i}^{m-1})}.$
@@
 as we wanted to show.
-Only if part.
-By Remark 1 we can show that for each $s_{-i},\in S_{-i}$, for every generalized self-odds ratio, equation 14 holds.
-Without loss of generality, we can assume that the generalized self-odds ratio is not infinite (because an infinite one can be inverted and becomes zero), so that no denominator in equation 14 is zero.
-For each $m=2,...,n$, by consistency, equation 11 for $E=\left\{s_{-i}^{m}\right\}$ and $\left(h_{i}^{1},...,h_{i}^{n}\right)=(h_{i}^{m-1},h_{i}^{m})$ yields
+Only if By Remark 1 it is enough to show that for each $s_{-i},\in S_{-i}$, for every simple generalized self-odds ratio, equation (15) holds.
+Without loss of generality, we can assume that the simple generalized self-odds ratio is not infinite (because an infinite one can be inverted and becomes zero), so that no denominator in equation (15) is zero.
+For each $m\in\left\{2,...,n\right\}$, by consistency, equation (12) for $E=\left\{s_{-i}^{m}\right\}$ and $\left(h_{i}^{1},...,h_{i}^{n}\right)=(h_{i}^{m-1},h_{i}^{m})$ yields
 $\frac{\mu^{i}(s_{-i}^{m}|h_{i}^{m})}{\mu^{i}(s_{-i}^{m}|h_{i}^{m-1})}=\frac{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m})|h_{i}^{m})}{\mu^{i}(S_{-i}(h_{i}^{m-1})\cap S_{-i}(h_{i}^{m-1})|h_{i}^{m-1})}.$
-Then, substituing in 14, we obtain equation 13, which coincides with equation 12 for $E=\left\{s_{-i}\right\}$, and thus holds by consistency.
+Then, substituting in (15), we obtain equation (14), which coincides with equation (13) for $E=\left\{s_{-i}\right\}$, and thus holds by consistency.
 A.2 Completion of the proofs of the theorems
-Proof of Theorem 1: $1\Rightarrow 2$
-Fix a non-terminal contingency $h$ and let $h^{1},...,h^{n}$ be the next contingencies.
-Let $\alpha^{1},...,\alpha^{n}$ denote the weights that verify the definition of Bayes-plausibility.
-We construct probability distributions $(p(h^{m}|h,s))_{m=1}^{n}$ of reaching each $h^{m}$ given $h$ and each $s\in S(h)$.
+Proof of Theorem 1:
+$\mathbf{1\Rightarrow 2)}$ Fix a non-terminal contingency $h$ and let $h^{1},...,h^{n}$ be the next contingencies.
+Let $\alpha_{1},...,\alpha_{n}$ denote the weights that verify the definition of Bayes-plausibility.
+We construct probability distributions $(\widetilde{p}(h^{m}|h,s))_{m=1}^{n}$ of reaching each $h^{m}$ given $h$ and each $s\in S(h)$.
 Fix $m\in\left\{1,...,n\right\}$.
-If $\alpha^{m}=0$, let $p(h^{m}|h,s)=0$ for all $s\in\mathcal{S}$.
-Suppose now that $\alpha^{m}>0$.
-For each $s\in\mathcal{S}$, let
-$p(h^{m}|h,s)$ $=$ $\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\text{ \ \ if }s\in\mathrm{supp}\mu(\cdot|h),$
-$p(h^{m}|h,s)$ $=$ $\alpha^{m}\text{ \ \ \ \ \ \ \ \ \ if }s\not\in\mathrm{supp}\mu(\cdot|h).$
+If $\alpha_{m}=0$, let $\widetilde{p}(h^{m}|h,s)=0$ for all $s\in\mathcal{S}$.
+Suppose now that $\alpha_{m}>0$.
+For each $s\in S(h)$, let
+$\widetilde{p}(h^{m}|h,s)=\left\{\begin{tabular}[]{ll}$\frac{\mu(s|h^{m})\alpha_{m}}{\mu(s|h)}$&$\text{if }s\in\mathrm{supp}\mu(\cdot|h)$\\ $\alpha^{m}$&otherwise.\end{tabular}\right$.
 For each $s\not\in\mathrm{supp}\mu(\cdot|h)$, we have
-$\sum_{m=1}^{n}p(h^{m}|h,s)=\sum_{m=1}^{n}\alpha^{m}=1.$
+$\sum_{m=1}^{n}\widetilde{p}(h^{m}|h,s)=\sum_{m=1}^{n}\alpha_{m}=1.$
 For each $s\in\mathrm{supp}\mu(\cdot|h)$, we have
-$\sum_{m=1}^{n}p(h^{m}|h,s)=\frac{\sum_{m=1}^{n}\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}=1,$
+$\sum_{m=1}^{n}\widetilde{p}(h^{m}|h,s)=\frac{\sum_{m=1}^{n}\mu(s|h^{m})\alpha_{m}}{\mu(s|h)}=1,$
 where the last equality follows from Bayes-plausibility.
-Now fix $m\in\left\{1,...,n\right\}$. If $\alpha^{m}=0$, then
-$\sum_{s\in S(h)}p(h^{m}|h,s)=0.$
-If $\alpha^{m}>0$, by Bayes-plausibility
+Now fix $m\in\left\{1,...,n\right\}$. If $\alpha_{m}=0$, then
+$\sum_{s\in S(h)}\widetilde{p}(h^{m}|h,s)=0.$
+If $\alpha_{m}>0$, by Bayes-plausibility
 $\mathrm{supp}\mu(\cdot|h^{m})\subseteq\mathrm{supp}\mu(\cdot|h).$ (16)
 So we have
-$\frac{p(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}p(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=\frac{\frac{\mu(s|h^{m})\alpha^{m}}{\mu(s|h)}\mu(s|h)}{\alpha^{m}\sum_{s^{\prime}\in\mathrm{supp}\mu(\cdot|h)}\mu(s|h^{m})}=\mu(s|h^{m}),\qquad\forall s\in\mathcal{S}$
-where the last equality follows from (15). Thus, $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(p(h^{m}|h,s))_{s\in S(h)}$ whenever possible.
-Proof of Theorem 2: $3\Rightarrow 1$
-We prove the counterpositive.
+$\frac{\widetilde{p}(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}\widetilde{p}(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=\frac{\frac{\mu(s|h^{m})\alpha_{m}}{\mu(s|h)}\mu(s|h)}{\alpha_{m}\sum_{s^{\prime}\in\mathrm{supp}\mu(\cdot|h)}\mu(s^{\prime}|h^{m})}=\mu(s|h^{m}),\qquad\forall s\in S\left(h\right)$
+where the last equality follows from (16) and
+$\frac{\widetilde{p}(h^{m}|h,s)\mu(s|h)}{\sum_{s^{\prime}\in S(h)}\widetilde{p}(h^{m}|h,s^{\prime})\mu(s^{\prime}|h)}=0=\mu(s|h^{m})\qquad\forall s\notin S\left(h\right)\text{.}$
+Thus, $\mu(\cdot|h^{m})$ can be derived with Bayes rule from $\mu(\cdot|h)$ using $(\widetilde{p}(h^{m}|h,s))_{s\in S(h)}$ whenever possible.
+$\mathbf{2\Rightarrow 3)}$ Fix a system of bets that the agent accepts.
+We show that it is not a deterministic Dutch book.
+Fix an initial contingency $h$.
+Along the paths $\bar{L}$ that start from $h$ and are consistent with some $s\insupp$$\mu(\cdot|h)$, forward-consistency is equivalent to updating by conditioning a joint belief over states and paths with marginal $\mu(\cdot|h)$ over the states.
+Thus, the Dutch-book argument for conditioning (Teller, 1973; Lewis, 1999) guarantees that the bookmaker either makes zero profit along every $l\in\bar{L}$, or negative profit along some $\bar{l}\in\bar{L}$, under some state $\bar{s}$ that is consistent with $\bar{l}$.
+[^21] This applies to betting on the state-path pair, and hence, as a subcase, also to our case of betting on the state only.
+In the second case, the system is not a deterministic Dutch book, as it violates condition of equation (3) with $s=\bar{s}$ and $l=\bar{l}$.
+The first case is impossible: since acceptance (given the tie-breaking rule) guarantees that the agent expects a positive payoff under every bet along the paths in $\bar{L}$, one could lower the payouts to the agent, preserve acceptance, and give positive profits to the bookmaker along all the paths, contradicting the Dutch-book argument for conditioning.
+Proof of Lemma 1.
 Suppose that there exist two different GORs between some $s$ and $s^{\prime}$.
@@
 we obtain a new GOR between $s$ and itself, which is not $1$; write it as
-$r:=o(s^{0},s^{1}|h^{1})\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-1},s^{n}|h^{n}),$
-where $n\geq 2$, $s^{0}=s^{n}=s$, and for each $k=1,...,n$,
-$o(s^{k-1},s^{k}|h^{k})=\frac{p(h^{k}|s^{k-1})}{p(h^{k}|s^{k})}\frac{\mu(s^{k-1}|h^{k})}{\mu(s^{k}|h^{k})}.$ (16)
-Wwithout loss of generality, we can suppose that:
-$s^{0},...,s^{n-1}$ are all distinct, because if $s^{j}=s^{k}$, either $r^{\prime}=o(s^{j},s^{j+1}|h^{j})\cdot...\cdot o(s^{k-1},s|h^{k})$, or $r^{\prime\prime}=r/r^{\prime}$ is a new GOR between $s$ and itself that is not $1$.
-$r<1$, because if a GOR is larger than $1$, inverting all the DORs we obtain a GOR smaller than $1$ (still between $s$ and itself).
-With this, we construct a Dutch Book that the agent is willing to accept.
-For every $h\in\mathcal{H}\backslash\left\{h^{1},...,h^{n}\right\}$, let $g(\widetilde{s}|h)=0$ for every $\widetilde{s}\in\mathcal{S}$, so that $i$ is willing to accept $g(\cdot|h)$.
-Fix $\varepsilon>0$.
-Let $g(\cdot|h^{1})$ be the following gamble on $S(h^{1})$:
+$r:=\prod_{i=1}^{n}o(s^{i-1},s^{i}|h^{i}),$
+where $n\geq 2$, $s^{0}=s^{n}=s$, and for each $k\in\left\{1,...,n\right\}$,
+$o(s^{k-1},s^{k}|h^{k})=\frac{p(h^{k}|s^{k})}{p(h^{k}|s^{k-1})}\frac{\mu(s^{k-1}|h^{k})}{\mu(s^{k}|h^{k})}.$ (17)
+Without loss of generality, we can suppose that:
+$s^{0},...,s^{n-1}$ are all distinct, because if $s^{j}=s^{k}$, either $r^{\prime}=o(s^{j},s^{j+1}|h^{j})\cdot...\cdot o(s^{k-1},s^{k}|h^{k})$ is a new GOR between $s^{j}$ and itself that is not $1$, or $r^{\prime\prime}=r/r^{\prime}$ is a new GOR between $s$ and itself that is not $1$;
+for each k=i,,n1k=i,...,n-1, $h^{k}\not=h^{k+1}$, because if $h^{k}=h^{k+1}=h$, we have $o(s^{k-1},s^{k}|h^{k})\cdot o(s^{k},s^{k+1}|h^{k+1})=o(s^{k-1},s^{k+1}|h)$ and we can substitute the left-hand side with the right-hand side in the GOR (eliminating state $s^{k}$) — likewise, we can suppose $h^{n}\not=h^{1}$;
+$r<1$, because if a GOR is larger than $1$, inverting all the DORs we obtain a GOR smaller than $1$ (still between $s$ and itself). $\blacksquare$
+Proof of Proposition 1.
+Consider the GOR given by the definition of $(r,n)$-not-coherent.
+We fix a $\delta>0$ and use this GOR to construct a $\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$-robust Dutch book.
+Fix $\varepsilon>0$ and let $\alpha=\sqrt[n]{\frac{1}{r}}$.
+Define $g(s^{0}|h^{1})$ and $g(s^{1}|h^{1})$ as:
 $g(s^{0}|h^{1})$ $=$ $-1,$
-$g(s^{1}|h^{1})$ $=$ $-g(s^{0}|h^{1})\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon=\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon,$ (17)
-and $g(s^{\prime}|h^{1})=0$ for every $s^{\prime}\not=s^{0},s^{1}$. Recursively, for each $m=2,...,n$, let $g(\cdot|h^{m})$ be the following gamble on $S(h^{m})$:
-$g(s^{m-1}|h^{m})$ $=$ $-g(s^{m-1}|h^{m-1})\frac{p(h^{m-1}|s^{m-1})}{p(h^{m}|s^{m-1})}-\varepsilon,$ (18)
-$g(s^{m}|h^{m})$ $=$ $-g(s^{m-1}|h^{m})\frac{\mu(s^{m-1}|h^{m})}{\mu(s^{m}|h^{m})}+\varepsilon,$ (19)
-and $g(s^{\prime}|h^{m})=0$ for every $s^{\prime}\not=s^{m},s^{m-1}$. For each $m=1,...,n$, the agent is willing to accept $g(\cdot|h^{m})$ because
-$\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})+\mu(s^{m}|h^{m})\cdot g(s^{m}|h^{m})$
-$=$ $\mu(s^{m-1}|h^{m})\cdot g(s^{m-1}|h^{m})-g(s^{m-1}|h^{m})\cdot\mu(s^{m-1}|h^{m})+\varepsilon\cdot\mu(s^{m}|h^{m})$
-$=$ $\varepsilon\cdot\mu(s^{m}|h^{m})>0\text{,}$
-where $\mu(s^{m}|h^{m})>0$ because $r\not=\infty$.
-Thus, $i$ is willing to accept $g=(g(\cdot|h))_{h\in\mathcal{H}}$.
-If $m\not=n$, since $s^{m}\not=s^{m^{\prime}}$ for every $m^{\prime}\not=m$, we also have
-$\sum_{h\in\mathcal{H}}p(h|s^{m})g(s^{m}|h)$ $=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})+p(h^{m+1}|s^{m})g(s^{m}|h^{m+1})$
-$=$ $p(h^{m}|s^{m})g(s^{m}|h^{m})-p(h^{m+1}|s^{m})g(s^{m}|h^{m})\frac{p(h^{m}|s^{m})}{p(h^{m+1}|s^{m})}-p(h^{m+1}|s^{m})\varepsilon$
-$=$ $-p(h^{m+1}|s^{m})\varepsilon<0,$
-where $p(h^{m+1}|s^{m})>0$ because $s^{m}\in S(h^{m+1})$ (DORs at a contingency are only defined between consistent states).
-Thus, $s^{m}$ verifies the strict inequality condition of a Dutch book.
-Every $\widetilde{s}\in\mathcal{S}\backslash\left\{s^{1},...,s^{n}\right\}$ trivially satisfies the weak inequality condition.
-Finally, we need to show that, for some $\varepsilon>0$, $s$ satisfies the weak inequality condition as well, so that $g$ is a Dutch book.
+$g(s^{1}|h^{1})$ $=$ $-g(s^{0}|h^{1})\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon=\frac{\mu(s^{0}|h^{1})}{\mu(s^{1}|h^{1})}+\varepsilon.$
+Recursively, for each $m\in\left\{2,...,n\right\}$ define $g(s^{m-1}|h^{m})$ and $g(s^{m}|h^{m})$ as: [^23] Since $h^{m}\not=h^{m-1}$, we are not redefining $g(s^{m-1}|h^{m-1})$.
+$g(s^{m-1}|h^{m})$ $=$ $-g(s^{m-1}|h^{m-1})\frac{p(h^{m-1}|s^{m-1})}{p(h^{m}|s^{m-1})}\sqrt[n]{\frac{1}{r}},$ (18)
+$g(s^{m}|h^{m})$ $=$ $-g(s^{m-1}|h^{m})\frac{\mu(s^{m-1}|h^{m})}{\mu(s^{m}|h^{m})}+\varepsilon.$ (19)
+Complete the system of bets with $g(\widetilde{s}|h)=0$ wherever $g(\widetilde{s}|h)$ is not already defined.
+For each $m\in\left\{1,...,n\right\}$, the agent is willing to accept $g(\cdot|h^{m})$ because
+$\mu(s^{m-1}|h^{m})g(s^{m-1}|h^{m})+\mu(s^{m}|h^{m})g(s^{m}|h^{m})$
+$=$ $\mu(s^{m-1}|h^{m})g(s^{m-1}|h^{m})-g(s^{m-1}|h^{m})\mu(s^{m-1}|h^{m})+\varepsilon\mu(s^{m}|h^{m})=\varepsilon\mu(s^{m}|h^{m})>0\text{,}$
+where $\mu(s^{m}|h^{m})>0$ because $r\not$$\infty$.
+[^24] If $h^{m}=h^{k}=h$ for some $k\not=m$, the same inequality holds for the other involved states, so $g(\cdot|h)$ is accepted.
+Thus, the agent is willing to accept $g=(g(\cdot|h))_{h\in\mathcal{H}}$.
+Let $\hat{p}$ be such that $\max_{h,h^{\prime}\in\mathcal{H},s\in S}\frac{\hat{p}(h|s)}{\hat{p}(h^{\prime}|s)}\frac{p(h^{\prime}|s)}{p\left(h|s\right)}<\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$.
+For each $m\in\left\{1,...,n-1\right\}$ (recall that $s^{n}=s^{1}$), since $s^{m}\not=s^{m^{\prime}}$ for every $m^{\prime}\not=m$, we have
+$\sum_{h\in\mathcal{H}}\hat{p}(h|s^{m})g(s^{m}|h)$ $=$ $\hat{p}(h^{m}|s^{m})g(s^{m}|h^{m})+\hat{p}(h^{m+1}|s^{m})g(s^{m}|h^{m+1})$
+$=$ $\hat{p}(h^{m}|s^{m})g(s^{m}|h^{m})-\hat{p}(h^{m+1}|s^{m})g(s^{m}|h^{m})\frac{p(h^{m}|s^{m})}{p(h^{m+1}|s^{m})}\sqrt[n]{\frac{1}{r}}<0,$
+[…]
+$p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$ $\leq$ $0\iff g(s|h^{1})+\frac{p(h^{n}|s)}{p(h^{1}|s)}g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)\leq 0$
+$\iff$ $g(s|h^{1})+\frac{p(h^{n}|s)/\hat{p}(h^{n}|s)}{p(h^{1}|s)/\hat{p}(h^{1}|s)}\frac{\hat{p}(h^{n}|s)}{\hat{p}(h^{1}|s)}g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)\leq 0$
+$\implies$ $g(s|h^{1})+\frac{\hat{p}(h^{n}|s)}{\hat{p}(h^{1}|s)}g(s|h^{n})\leq 0$
+$\iff$ $\hat{p}(h^{1}|s)g(s|h^{1})+\hat{p}(h^{n}|s)g(s|h^{n})\leq 0$
+$\iff$ $\sum_{h\in\mathcal{H}}\hat{p}(h|s)g(s|h)\leq 0\text{.}$
 Note that, for $\varepsilon=0$,
-$\sum_{h\in\mathcal{H}}p(h|s)g(s|h)$
-$=$ $p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})$
-$=$ $-p(h^{1}|s)-p(h^{n}|s)g(s^{n-1}|h^{n})\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
-$=$ $-p(h^{1}|s)+p(h^{n}|s)g(s^{n-1}|h^{n-1})\frac{p(h^{n-1}|s^{n-1})}{p(h^{n}|s^{n-1})}\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}$
-$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-1}|h^{n-1})o(s^{n-1},s|h^{n})$
-$=$ $-p(h^{1}|s)-p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-1})\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
-$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-2})\frac{p(h^{n-2}|s^{n-2})}{p(h^{n-1}|s^{n-2})}\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\cdot o(s^{n-1},s|h^{n})$
-$=$ $-p(h^{1}|s)+p(h^{n-2}|s^{n-2})g(s^{n-2}|h^{n-2})o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h^{n})$
-$=$ $-p(h^{1}|s)-p(h^{1}|s^{1})g(s^{0}|h^{1})\frac{\mu(s|h^{1})}{\mu(s^{1}|h^{1})}\cdot o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-2},s^{n-1}|h^{n-1})\cdot o(s^{n-1},s|h)$
-$=$ $-p(h^{1}|s)+p(h^{1}|s)r<0\text{,}$
+$p(h^{1}|s)g(s|h^{1})+p(h^{n}|s)g(s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)-p(h^{n}|s)g(s^{n-1}|h^{n})\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)+p(h^{n}|s)g(s^{n-1}|h^{n-1})\frac{p(h^{n-1}|s^{n-1})}{p(h^{n}|s^{n-1})}\sqrt[n]{\frac{1}{r}}\frac{\mu(s^{n-1}|h^{n})}{\mu(s|h^{n})}\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-1}|h^{n-1})\sqrt[n]{\frac{1}{r}}o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)-\beta p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-1})\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}\sqrt[n]{\frac{1}{r}}o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)+p(h^{n-1}|s^{n-1})g(s^{n-2}|h^{n-2})\frac{p(h^{n-2}|s^{n-2})}{p(h^{n-1}|s^{n-2})}\left(\sqrt[n]{\frac{1}{r}}\right)^{2}\frac{\mu(s^{n-2}|h^{n-1})}{\mu(s^{n-1}|h^{n-1})}o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)+p(h^{n-2}|s^{n-2})g(s^{n-2}|h^{n-2})\left(\sqrt[n]{\frac{1}{r}}\right)^{2}o(s^{n-2},s^{n-1}|h^{n-1})o(s^{n-1},s|h^{n})\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)-p(h^{1}|s^{1})g(s^{0}|h^{1})\frac{\mu(s|h^{1})}{\mu(s^{1}|h^{1})}\left(\sqrt[n]{\frac{1}{r}}\right)^{n-1}o(s^{1},s^{2}|h^{2})\cdot...\cdot o(s^{n-2},s^{n-1}|h^{n-1})o(s^{n-1},s|h)\left(\sqrt[n]{\frac{1}{r}}-\delta\right)$
+$=$ $-p(h^{1}|s)+p(h^{1}|s)r\left(\sqrt[n]{\frac{1}{r}}\right)^{n-1}\left(\sqrt[n]{\frac{1}{r}}-\delta\right)=-p(h^{1}|s)+p(h^{1}|s)r\left(\sqrt[n]{\frac{1}{r}}\right)^{n}\frac{\left(\sqrt[n]{\frac{1}{r}}-\delta\right)}{\sqrt[n]{\frac{1}{r}}}$
+$\leq$ $-p(h^{1}|s)+p(h^{1}|s)\frac{\left(\sqrt[n]{\frac{1}{r}}-\delta\right)}{\sqrt[n]{\frac{1}{r}}}<0\text{,}$
 where the second equality uses (19), the third (18), the fourth (17), and so on. By continuity, the strict inequality is preserved for sufficiently small $\varepsilon>0$. $\blacksquare$